### v1

# Adjoint log canonical foliated singularities on surfaces

Abstract.
Let $(X,\mathcal{F})$ be a foliated surface over $\mathbb{C}$. We study the singularities of the adjoint foliated divisor $K_{\mathcal{F}}+\epsilon K_{X}$. We provide a complete classification of $\epsilon$-adjoint log canonical singularities of foliated surfaces for $\epsilon\in(0,1/3)$. Moreover, we prove that for any $\epsilon\in(0,1/5)$, every $\epsilon$-adjoint log canonical singularity is log canonical for $\mathcal{F}$, and that for any $\epsilon\in(0,1/4)$, every $\epsilon$-adjoint canonical singularity is log canonical for $\mathcal{F}$. Finally, we present examples showing that both bounds are sharp.
1. Introduction
Throughout this paper, we work over the complex numbers $\mathbb{C}$.
The study of singularities lies at the core of algebraic geometry. Singular varieties arise naturally, for instance, in the minimal model program and in the construction of moduli spaces. Among the most fundamental classes are (log) canonical singularities, which have been extensively studied, particularly in dimension two. In fact, log canonical surface singularities admit a complete classification (see [Alex92], [KM98, Sec. 4.1], or Theorem A.8).
In recent years, the theory of foliations in algebraic geometry has attracted increasing attention. It is therefore natural to seek foliated analogues of the classical theory of singularities. In this direction, McQuillan [McQ08] introduced the notion of (log) canonical singularities for foliations (cf. Definition 2.12), and obtained a classification of canonical foliated singularities on surfaces (see [McQ08, Corollary I.2.2, Fact I.2.4, Theorem III.3.2]). Subsequently, Y. Chen classified log canonical foliated singularities on surfaces (see [Chen23, Theorem 1.1] or Theorem B.9).
More recently, Pereira and Svaldi introduced adjoint foliated structures in [PS19]. This framework was further developed by Spicer and Svaldi for rank-one foliations on surfaces [SS23], and by Cascini–Han–Liu–Meng–Spicer–Svaldi–Xie for algebraically integrable foliations on higher-dimensional varieties [CHL+24]. Adjoint foliated structures have proved effective in addressing several difficulties in the birational geometry of foliations. For instance, effective birationality fails even for rank-one foliations on surfaces [Lu25]. The underlying idea is that, instead of studying the canonical divisor $K_{\mathcal{F}}$ alone, one considers divisors of the form $K_{\mathcal{F}}+\epsilon K_{X}$ for $0<\epsilon\ll 1$, which allows one to exploit classical results for $K_{X}$.
In this paper, we study singularities of adjoint foliated structures $K_{\mathcal{F}}+\epsilon K_{X}$ on surfaces. Our goal is to understand how $\epsilon$-adjoint singularities relate to the usual singularities of foliations. More precisely, we consider the following questions.
  • Q1.
    Does there exist $\epsilon_{0}>0$ such that $\epsilon$-adjoint canonical singularities are canonical for all $\epsilon\in(0,\epsilon_{0})$?
  • Q2.
    Does there exist $\epsilon_{0}>0$ such that $\epsilon$-adjoint log canonical singularities are log canonical for all $\epsilon\in(0,\epsilon_{0})$?
We first observe that Question (Q1) admits a negative answer (see Example 2.15). This shows that $\epsilon$-adjoint canonical singularities need not be canonical even for $\epsilon>0$ sufficiently small.
Motivated by this failure of Question (Q1), we are led to consider the following weaker variant.
  • Q3.
    Does there exist $\epsilon_{0}>0$ such that $\epsilon$-adjoint canonical singularities are log canonical for all $\epsilon\in(0,\epsilon_{0})$?
The main purpose of this paper is to address these questions. We give a complete classification of $\epsilon$-adjoint log canonical singularities of foliations for $\epsilon\in(0,1/3)$, and prove that Questions (Q2) and (Q3) admit affirmative answers.
We now state our main results.
Theorem 1.1 (cf. Theorem 4.14).
Fix $\epsilon\in(0,1/5)$. Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface, and assume that $p$ is an $\epsilon$-adjoint log canonical singularity of $(Y,\mathcal{G})$. Let
$\pi\colon(X,\mathcal{F})\to(Y,\mathcal{G})$
be the minimal resolution (cf. Definition 2.5) of $(Y,\mathcal{G},p)$, with exceptional divisor $E=\bigcup_{i=1}^{r}E_{i}$. Then $E$ belongs to one of the following families:
  1. (1)
    Chains of $\mathcal{F}$-invariant curves: either an $\mathcal{F}$-chain (Figure 3), two (1)(-1)-$\mathcal{F}$-curves of self-intersection 2-2 joined by a bad tail (Figure 4), or a chain of (2)(-2)-$\mathcal{F}$-curves (Definition 2.18).
  2. (2)
    $\mathcal{F}$-dihedral graphs: two (1)(-1)-$\mathcal{F}$-curves of self-intersection 2-2 joined by a bad tail, which in turn is attached to a chain of (2)(-2)-$\mathcal{F}$-curves (see Figure 5).
  3. (3)
    Elliptic Gorenstein leaves: either a rational $\mathcal{F}$-invariant curve with a single node, or a cycle of (2)(-2)-$\mathcal{F}$-curves (see Figure 6).
  4. (4)
    Smooth elliptic non-$\mathcal{F}$-invariant curves: a smooth elliptic non-$\mathcal{F}$-invariant curve $E_{1}$ with ${\rm tang}(\mathcal{F},E_{1})=0$.
  5. (5)
    $\mathcal{F}$-star chains: a chain containing a unique smooth rational non-$\mathcal{F}$-invariant curve $E_{k}$ with ${\rm tang}(\mathcal{F},E_{k})=0$, while the remaining components form $\mathcal{F}$-chains (see Figure 8).
  6. (6)
    $\mathcal{F}$-star graphs: a smooth rational non-$\mathcal{F}$-invariant curve $C$ with ${\rm tang}(\mathcal{F},C)=0$ at the center, together with attached $\mathcal{F}$-chains of type $(l;m_{1},\dots,m_{l})$, where $(l;m_{1},\dots,m_{l})$ is one of the following:
    1. (a)
      $l=3$, $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, $(2,3,4)$, $(2,3,5)$;
    2. (b)
      $l=3$, $(m_{1},m_{2},m_{3})=(2,3,6)$, $(2,4,4)$, or $(3,3,3)$;
    3. (c)
      $l=4$, $(m_{1},\dots,m_{4})=(2,2,2,2)$.
    (See Definition 2.22 and Figure 7.)
Moreover, configurations of type (3), (4), (6-b), and (6-c) are not $\epsilon$-adjoint klt.
Remark 1.2.
For $\epsilon\in[1/5,1/3)$, there are additional configurations: (cf. Theorem 4.14)
  1. (7)
    A rational smooth $\mathcal{F}$-invariant curve $C$ with ${\rm Z}(\mathcal{F},C)=3$ connecting to two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$ of type $(2,3)$.
  2. (8)
    A rational smooth $\mathcal{F}$-invariant curve $C$ with ${\rm Z}(\mathcal{F},C)=3$ connecting to two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$ of type $(2,3)$ and a chain of (2)(-2)-$\mathcal{F}$-curves.
Here, by an $\mathcal{F}$-chain $\Theta$ of type $m$ we mean that $m$ denotes the absolute value of the determinant of the intersection matrix of $\Theta$.
Remark 1.3.
For $\epsilon\in(0,1/4)$, cases (3), (4), (6), (7) and (8) fail to be $\epsilon$-adjoint canonical (see Theorem 5.6).
Therefore, we have the following proposition.
Proposition 1.4.
Let $(X,\mathcal{F})$ be a foliated surface.
  1. (1)
    For any $\epsilon\in(0,1/5)$, if $p$ is an $\epsilon$-adjoint log canonical singularity of $(X,\mathcal{F})$, then $p$ is a log canonical singularity of $\mathcal{F}$. Moreover, $p$ is a log canonical singularity of $X$. If furthermore $p$ is $\epsilon$-adjoint klt, then $p$ is a klt singularity of $X$.
  2. (2)
    For any $\epsilon\in(0,1/4)$, if $p$ is an $\epsilon$-adjoint canonical singularity of $(X,\mathcal{F})$, then $p$ is a log canonical singularity of $\mathcal{F}$ and a klt singularity of $X$.
Proof.
This is a direct consequence of Theorem 4.14, Theorem 5.6 and Theorem A.8. ∎
Remark 1.5.
  1. (1)
    Proposition 1.4(1) answers a question posed to the author by Professor Jihao Liu. In [SS23], the corresponding result was proved under the assumption that $p$ is smooth. In [Vas25], it was proved under the assumption that $p$ is a klt singularity.
  2. (2)
    The bounds in Proposition 1.4 are both sharp, see Example 2.16.
As a direct corollary of Proposition 1.4(2), we obtain the following result.
Corollary 1.6.
Let $X$ be a smooth projective surface and $\mathcal{F}$ a rank one foliation with canonical singularities. Then for any $\epsilon\in(0,1/4)$, there exists a birational morphism $\varphi:X\to Y$ (which is a $(K_{\mathcal{F}}+\epsilon K_{X})$-non-positive contraction) such that either
  1. (1)
    $K_{\mathcal{G}}+\epsilon K_{Y}$ is nef, where $\mathcal{G}=\varphi_{*}\mathcal{F}$; or
  2. (2)
    there exists a morphism $f:Y\to Z$ such that $\rho(Y/Z)=1$ and $(K_{\mathcal{G}}+\epsilon K_{Y})$ is $f$-ample.
Moreover, $Y$ has klt singularities and $\mathcal{G}$ has log canonical singularities.
Proof.
This directly follows from Proposition 1.4(2) and [KM98, Corollary 3.17]. ∎
Remark 1.7.
Corollary 1.6 improves [SS23, Theorem 1.1], where the authors treat the case $\epsilon\in(0,1/5)$. It also provides a useful complement to [LWX25, Theorem 1.1].
2. Preliminaries
2.1. Foliations and singularities
A foliation $\mathcal{F}$ on a normal surface $X$ is a rank 1 saturated subsheaf $T_{\mathcal{F}}$ of the tangent sheaf $T_{X}$. Note that $T_{X}$ is reflexive as it is isomorphic to ${\rm Hom}_{\mathcal{O}_{X}}(\Omega_{X},\mathcal{O}_{X})$. Thus $T_{\mathcal{F}}$ is also reflexive and we can define the canonical divisor $K_{\mathcal{F}}$ of the foliation as a Weil divisor on $X$ satisfying $\mathcal{O}_{X}(-K_{\mathcal{F}})\cong T_{\mathcal{F}}$. A foliated surface $(X,\mathcal{F})$ consists of a normal surface $X$ together with a foliation $\mathcal{F}$ on $X$.
A point $p\in X$ is called a singularity of the foliation $\mathcal{F}$ if either a singular point of $X$ or a point at which the quotient $T_{X}/T_{\mathcal{F}}$ is not locally free. We say $p$ is a regular point of $\mathcal{F}$ if $p$ is a smooth point of $X$ and $T_{X}/T_{\mathcal{F}}$ is locally free at $p$.
Let $p$ be a singularity of $\mathcal{F}$, at which $X$ is smooth. Locally, $\mathcal{F}$ is defined by a vector field
(2.1) $\nu=A(x,y)\frac{\partial}{\partial x}+B(x,y)\frac{\partial}{\partial y},$
where $p=(0,0)$. The two eigenvalues $\lambda_{1},\lambda_{2}$ of the linear part $(D\nu)(p)$ of $\nu$ at $p$ are well defined.
Definition 2.1.
Let $p$ be a singularity of $\mathcal{F}$ (at which $X$ is smooth) and let $\lambda_{1},\lambda_{2}$ be as above.
  1. (1)
    The singularity $p$ is called non-degenerated if the two eigenvalues $\lambda_{1},\lambda_{2}$ are both nonzero.
  2. (2)
    The singularity $p$ is called reduced if one of the two eigenvalues, say, $\lambda_{2}$, is nonzero and and the quotient $\lambda=\lambda_{1}/\lambda_{2}$ is not a positive rational number. In particular, if $\lambda=0$, we call $p$ a saddle-node.
A foliation $\mathcal{F}$ is said to be reduced if any singularity of $\mathcal{F}$ is reduced.
Remark 2.2.
If $p$ is a saddle-node, by a suitable transition of coordinates, the foliation $\mathcal{F}$ at $p$ can be represented by
$\nu$ $=\left(x+axy^{k}+yF(x,y)\right)\frac{\partial}{\partial x}+y^{k+1}\frac{\partial}{\partial y},$
$\text{or}\quad\omega$ $=\left(x+axy^{k}+yF(x,y)\right){\rm d}y-y^{k+1}{\rm d}x,$
where $a\in\mathbb{C}$, $k\in\mathbb{Z}^{+}$, and $F$ is a holomorphic function which vanishes at $p=(0,0)$ up to order $k$. The curve $(y=0)$ is a separatrix, called strong separatrix. If $F=0$, then the separatrix $(x=0)$ is called the weak separatrix.
Theorem 2.3 (Seidenberg).
Given any foliated surface $(X,\mathcal{F})$, there exists a sequence of blowing-ups $\sigma:X^{\prime}\to X$, such that the induced foliation $\mathcal{F}^{\prime}$ on $X^{\prime}$ is reduced.
Proof.
See [Sei68] or [Bru15, Theorem 1.1]. ∎
2.2. Minimal resolution of foliated surfaces
Definition 2.4.
Let $(X,\mathcal{F})$ be a foliated surface. A curve $C\subset X$ is said to be $\mathcal{F}$-exceptional if the following conditions are satisfied:
  1. (1)
    $C$ is a smooth rational curve with self-intersection 1-1;
  2. (2)
    the contraction of $C$ to a point $p$ yields a foliated surface $(X^{\prime},\mathcal{F}^{\prime})$, where $p$ is either a regular point or a reduced singularity of $\mathcal{F}^{\prime}$.
Definition 2.5.
Let $(X,\mathcal{F})$ be a foliated surface. Let $p$ is a singularity of $(X,\mathcal{F})$. A resolution of $(X,\mathcal{F},p)$ is a birational morphism
$\pi:(Y,\mathcal{G})\longrightarrow(X,\mathcal{F}),$
such that
  • (i)
    $Y-\pi^{-1}(p)\cong X-p$,
  • (ii)
    for any $q\in\pi^{-1}(p)$, $p$ is a smooth point of $Y$ and $p$ is a reduced (or regular) point of the induced foliation $\mathcal{G}:=\pi^{*}\mathcal{F}$.
We say the resolution $\pi$ is minimal if moreover
  • (iii)
    there exists no $\mathcal{G}$-exceptional curve contained in $\pi^{-1}(p)$.
2.3. Index Theorems
Let $(X,\mathcal{F})$ be a foliated surface with $X$ smooth.
A curve $C\subseteq X$ is said to be $\mathcal{F}$-invariant if the inclusion $T_{\mathcal{F}}|_{C}\rightarrow T_{X}|_{C}$ factors through $T_{C}$, where $T_{C}$ is the tangent bundle of $C$.
2.3.1. Non-invariant curves
For a non-$\mathcal{F}$-invariant curve $C$ and $p\in C$, we can define the tangency order of $\mathcal{F}$ along $C$ at $p$ to be
${\rm tang}(\mathcal{F},C,p):=\dim_{\mathbb{C}}\frac{\mathcal{O}_{p}}{\langle f,\nu(f)\rangle},$
where $\nu$ is the local generator of $\mathcal{F}$ at $p$ and $f=0$ is the local equation of $C$. We define
${\rm tang}(\mathcal{F},C):=\sum_{p\in C}{\rm tang}(\mathcal{F},C,p).$
Proposition 2.6.
We have
(2.2) ${\rm tang}(\mathcal{F},C)=K_{\mathcal{F}}\cdot C+C^{2}\,(\geq 0).$
Proof.
See [Bru15, Proposition 2.2]. ∎
Remark 2.7.
${\rm tang}(\mathcal{F},C)=0$ implies that $C$ is smooth.
2.3.2. Invariant curves
Let $C$ be an $\mathcal{F}$-invariant curve on $X$, and let $p\in C$ be a point. In a neighborhood of $p$, the foliation $\mathcal{F}$ is defined by a 1-form $\omega$, and $C$ is locally given by the equation $f=0$. Since $C$ is $\mathcal{F}$-invariant, we may write
$g\omega=hdf+f\eta,$
where $\eta$ is a holomorphic 1-form and $g,h$ are holomorphic functions defined around $p$, with $h$ and $f$ are coprime. We define
(2.3) ${\rm Z}(\mathcal{F},C,p)$ $:=\text{vanishing order of $\left.\frac{h}{g}\right|_{C}$ at $p$}$
(2.4) ${\rm CS}(\mathcal{F},C,p)$ $:={\rm Res}_{p}\left\{\left.-\frac{\eta}{h}\right|_{C}\right\}.$
By definition, ${\rm Z}(\mathcal{F},C,p)={\rm CS}(\mathcal{F},C,p)=0$ if $p$ is not a singularity of $\mathcal{F}$. If $p$ is a reduced singularity of $\mathcal{F}$, then ${\rm Z}(\mathcal{F},C,p)\geq 0$. Let
${\rm Z}(\mathcal{F},C):=\sum_{p\in C}{\rm Z}(\mathcal{F},C,p),\quad{\rm CS}(\mathcal{F},C):=\sum_{p\in C}{\rm CS}(\mathcal{F},C,p).$
Proposition 2.8.
Let $C$ be an $\mathcal{F}$-invariant curve on $X$. Then
(2.5) ${\rm Z}(\mathcal{F},C)$ $=K_{\mathcal{F}}\cdot C+2-2p_{a}(C),$
(2.6) ${\rm CS}(\mathcal{F},C)$ $=C^{2},$
where $p_{a}(C)$ denotes the arithmetic genus of $C$. The second equality above is called the Camacho-Sad formula.
Proof.
See [Bru15, Proposition 2.3 and Theorem 3.2]. ∎
Lemma 2.9.
Let $p$ is a reduced singularity of the foliated surface $(X,\mathcal{F})$.
  1. (1)
    If $p$ is nondegerate and $\mathcal{F}$ locally around $p$ defined by $\omega=\lambda y(1+o(1)){\rm d}x-x(1+o(1)){\rm d}y$, then
    $\begin{cases}{\rm CS}(\mathcal{F},x=0,p)=1/\lambda,\qquad{\rm CS}(\mathcal{F},y=0,p)=\lambda,\\ {\rm Z}(\mathcal{F},x=0,p)={\rm Z}(\mathcal{F},y=0,p)=1\end{cases}$
  2. (2)
    If $p$ is a saddle-node and $\mathcal{F}$ locally around $p$ defined by $\omega=y^{k+1}{\rm d}x-(x(1+ay^{k})+yo(k)){\rm d}y$, where $k\in\mathbb{Z}^{>0}$ and $a\in\mathbb{C}$, then
    ${\rm CS}(\mathcal{F},y=0,p)=0,\quad\text{and}\quad{\rm Z}(\mathcal{F},y=0,p)=1.$
    If furthermore there exists a weak separatrix $(x=0)$, then
    ${\rm CS}(\mathcal{F},x=0,p)=a,\quad\text{and}\quad{\rm Z}(\mathcal{F},x=0,p)=k+1.$
Proof.
See [Bru15, pp. 30-31]. ∎
Next, we recall the separatrix theorem.
Theorem 2.10 (Separatrix theorem).
Let $\mathcal{F}$ be a foliation on a smooth projective surface $X$ and let $C\subset X$ be a connected compact $\mathcal{F}$-invariant curve such that:
  • (i)
    All the singularities of $\mathcal{F}$ on $C$ are reduced (in particular, $C$ has only normal crossing singularities);
  • (ii)
    If $C_{1},\cdots,C_{n}$ are the irreducible components of $C$, then the intersection matrix $(C_{i}C_{j})_{1\leq i,j\leq n}$ is negative definite and the dual graph $\Gamma$ is a tree.
Then there exists at least one point $p\in C\cap{\rm Sing}(\mathcal{F})$ and a separatrix through $p$ not contained in $C$.
Proof.
See [Bru15, Theorem 3.4]. ∎
2.4. Pairs and Triples
Let $X$ be a normal surface and let $\mathcal{F}$ be a foliation on $X$. Let $D$ be an $\mathbb{R}$-divisor on $X$. Then we can uniquely decompose $D$ as
$D=D^{\mathrm{inv}}_{\mathcal{F}}+D^{\mathrm{n\!-\!inv}}_{\mathcal{F}},$
where $D^{\mathrm{inv}}_{\mathcal{F}}$ (resp. $D^{\mathrm{n\!-\!inv}}_{\mathcal{F}}$) denotes the $\mathcal{F}$-invariant part (resp. the non-$\mathcal{F}$-invariant part) of $D$, namely the sum of the components of $D$ whose support is $\mathcal{F}$-invariant (resp. non-$\mathcal{F}$-invariant). If there is no risk of confusion, we will omit the subscript $\mathcal{F}$ from the notation.
We say that $(X,\Delta)$ is a pair if $\Delta$ is an effective $\mathbb{R}$-divisor and $K_{X}+\Delta$ is $\mathbb{R}$-Cartier. A foliated pair $(\mathcal{F},\Delta)$ on $X$ consists of a foliation $\mathcal{F}$ on $X$ and an effective $\mathbb{R}$-divisor $\Delta$ such that $K_{\mathcal{F}}+\Delta$ is $\mathbb{R}$-Cartier.
Definition 2.11.
A foliated triple $(X,\mathcal{F},\Delta)$ consists of a foliation $\mathcal{F}$ on a normal surface $X$ and an effective $\mathbb{R}$-divisor $\Delta\geq 0$ such that both $K_{\mathcal{F}}+\Delta^{\mathrm{n\!-\!inv}}$ and $K_{X}+\Delta$ are $\mathbb{R}$-Cartier.
2.5. Log canonical foliated singularities
Let $\pi\colon X^{\prime}\to X$ be a birational morphism and let $(\mathcal{F},\Delta)$ be a foliated pair on $X$. Denote by $\mathcal{F}^{\prime}$ the induced foliation on $X^{\prime}$. Then we may write
$K_{\mathcal{F}^{\prime}}=\pi^{*}(K_{\mathcal{F}}+\Delta)+\sum_{E}a(E,\mathcal{F},\Delta)\,E,$
where the sum runs over all prime divisors $E$ on $X^{\prime}$ and
$\pi_{*}\!\left(\sum_{E}a(E,\mathcal{F},\Delta)\,E\right)=-\Delta.$
The rational number $a(E,\mathcal{F},\Delta)$ is called the discrepancy of the foliated pair $(\mathcal{F},\Delta)$ with respect to $E$.
Definition 2.12.
Let $X$ be a normal surface and let $(\mathcal{F},\Delta)$ be a foliated pair on $X$. We say that $(\mathcal{F},\Delta)$ is terminal (resp. canonical, klt, log canonical) if
$a(E,\mathcal{F},\Delta)>0\quad(\text{resp.\ }\geq 0,\ >-\iota(E)\ \text{and }\lfloor\Delta\rfloor=0,\ \geq-\iota(E))$
for any birational morphism $\pi\colon X^{\prime}\to X$ and for any prime divisor $E$ on $X^{\prime}$, where
$\iota(E):=\begin{cases}1,&\text{if $E$ is not $\mathcal{F}$-invariant},\\ 0,&\text{if $E$ is $\mathcal{F}$-invariant}.\end{cases}$
Moreover, we say that $(\mathcal{F},\Delta)$ is log terminal if
$a(E,\mathcal{F},\Delta)>-\iota(E)$
for any birational morphism $\pi\colon X^{\prime}\to X$ and for any $\pi$-exceptional prime divisor $E$ on $X^{\prime}$.
2.6. $(\epsilon,\delta)$-adjoint log canonical foliated singularities
Definition 2.13.
Let $(X,\mathcal{F},\Delta)$ be a foliated triple on a normal surface $X$, and fix $\epsilon\geq 0$. The $\epsilon$-adjoint divisor of $(X,\mathcal{F},\Delta)$ is defined by
$K_{(X,\mathcal{F},\Delta),\epsilon}:=(K_{\mathcal{F}}+\Delta^{\mathrm{n\!-\!inv}})+\epsilon(K_{X}+\Delta).$
Definition 2.14.
Let $(X,\mathcal{F},\Delta)$ be a foliated triple. Fix $\epsilon>0$ and $\delta\geq 0$. We say that $(X,\mathcal{F},\Delta)$ is $(\epsilon,\delta)$-adjoint log canonical (resp. $(\epsilon,\delta)$-adjoint klt) if for any birational morphism $\pi\colon X^{\prime}\to X$, writing
$K_{(X^{\prime},\mathcal{F}^{\prime},\Delta^{\prime}),\epsilon}=\pi^{*}K_{(X,\mathcal{F},\Delta),\epsilon}+\sum_{i}a_{i}E_{i},$
where the $E_{i}$ are $\pi$-exceptional divisors and $\Delta^{\prime}:=\pi_{*}^{-1}\Delta$ is the strict transform of $\Delta$, we have, for all $i$,
$a_{i}\geq(\iota(E_{i})+\epsilon)(-1+\delta)$
$(\text{resp.}\quad a_{i}>(\iota(E_{i})+\epsilon)(-1+\delta)\quad\text{and}\quad\lfloor\Delta\rfloor=0).$
When $\delta=1$, we call $(\epsilon,\delta)$-adjoint log canonical $\epsilon$-adjoint canonical. When $\delta=0$, we call $(\epsilon,\delta)$-adjoint log canonical simply $\epsilon$-adjoint log canonical.
Moreover, we say that $(\mathcal{F},\Delta)$ is $\epsilon$-adjoint terminal if $a_{i}>0$ for all $i$.
Next we introduce two important examples.
Example 2.15.
Let $p$ be a singularity of a foliated surface $(X,\mathcal{F})$. Assume that $X$ is smooth at $p$ and that $\mathcal{F}$ is locally defined near $p=(0,0)$ by the $1$-form
$\omega=ny\,{\rm d}x-x\,{\rm d}y.$
Consider the minimal resolution $\sigma\colon(X^{\prime},\mathcal{F}^{\prime})\to(X,\mathcal{F})$ over the point $p$ (see Figure 1).
$\cdots\cdots$
$\bar{E}_{2}(-2)$
$\bar{E}_{1}(-2)$
$\bar{E}_{n-1}(-2)$
$E_{n}(-1)$
Figure 1.
We denote by $q_{i}$ (resp. $E_{i}$), $i=1,\cdots,n$, the successive blow-up points (resp. the exceptional curves). In this case, we have
$l(q_{i})=a(q_{i})=1,\quad i\leq n-1,\qquad\text{and}\qquad l(q_{n})=2>a(q_{n})=1.$
In particular, each $E_{i}$ is $\mathcal{F}^{\prime}$-invariant for $i\leq n-1$ and $E_{n}$ is not $\mathcal{F}^{\prime}$-invariant. A direct computation gives
$K_{\mathcal{F}^{\prime}}$ $=\sigma^{*}(K_{\mathcal{F}})-E_{n},$
$K_{X^{\prime}}$ $=\sigma^{*}(K_{X})+\bar{E}_{1}+2\bar{E}_{2}+\cdots+(n-1)\bar{E}+nE_{n}.$
Therefore,
$K_{\mathcal{F}^{\prime}}+\epsilon K_{X^{\prime}}=\sigma^{*}(K_{\mathcal{F}}+K_{X})+\epsilon\sum_{i=1}^{n-1}i\bar{E}_{i}+(n\epsilon-1)E_{n}.$
For any $\epsilon\in(0,1)$, there exists some $n$ such that $n\epsilon-1\geq 0$. In this case, $p$ is $\epsilon$-adjoint canonical but not canonical.
Example 2.16.
Let $p$ be a singularity of a foliated surface $(X,\mathcal{F})$. Assume that $X$ is smooth at $p$ and that $\mathcal{F}$ is locally defined near $p=(0,0)$ by the $1$-form
$\omega=x\,\mathrm{d}x+y^{2}\,\mathrm{d}y.$
Consider the minimal resolution $\sigma\colon(X^{\prime},\mathcal{F}^{\prime})\to(X,\mathcal{F})$ over the point $p$ (see Figure 2).
$\bar{E}_{1}(-3)$
$\bar{E}_{2}(-2)$
$E_{3}(-1)$
Figure 2.
We denote by $q_{i}$ (resp. $E_{i}$) the successive blow-up points (resp. the exceptional curves). In this case, we have
$l(q_{1})=l(q_{2})=1\qquad\text{and}\qquad l(q_{3})=a(q_{3})=2.$
In particular, each $E_{i}$ is $\mathcal{F}^{\prime}$-invariant. A direct computation gives
$K_{\mathcal{F}^{\prime}}$ $=\sigma^{*}(K_{\mathcal{F}})-E_{3},$
$K_{X^{\prime}}$ $=\sigma^{*}(K_{X})+\bar{E}_{1}+2\bar{E}_{2}+4E_{3}.$
Therefore,
$K_{\mathcal{F}^{\prime}}+\frac{1}{4}K_{X^{\prime}}$ $=\sigma^{*}\!\left(K_{\mathcal{F}}+\frac{1}{4}K_{X}\right)+\frac{1}{4}\bar{E}_{1}+\frac{1}{2}\bar{E}_{2},$
$K_{\mathcal{F}^{\prime}}+\frac{1}{5}K_{X^{\prime}}$ $=\sigma^{*}\!\left(K_{\mathcal{F}}+\frac{1}{5}K_{X}\right)+\frac{1}{5}\bar{E}_{1}+\frac{2}{5}\bar{E}_{2}-\frac{1}{5}E_{3}.$
It follows that $p$ is $\tfrac{1}{4}$-adjoint canonical and $\tfrac{1}{5}$-adjoint log canonical, but it is not a log canonical singularity of $\mathcal{F}$.
2.7. Notation and basic configurations
In this subsection, let $(X,\mathcal{F})$ be a foliated surface with $X$ smooth. We introduce notation and basic configurations that will be used throughout the paper.
Definition 2.17 ($\mathcal{F}$-chain).
A divisor $\Theta=\Gamma_{1}+\cdots+\Gamma_{r}$ is called an $\mathcal{F}$-chain if
  1. (1)
    $\Theta$ is a Hirzebruch–Jung string (cf. [BHPV04, Ch.III; Sec.5]);
  2. (2)
    each irreducible component $\Gamma_{j}$ is $\mathcal{F}$-invariant;
  3. (3)
    ${\rm Sing}(\mathcal{F})\cap\Theta$ consists of reduced, non-degenerate singularities;
  4. (4)
    ${\rm Z}(\mathcal{F},\Gamma_{1})=1$ and ${\rm Z}(\mathcal{F},\Gamma_{i})=0$ for all $i\geq 2$.
The curve $\Gamma_{1}$ is called the first curve of the $\mathcal{F}$-chain (see Figure 3).
$\cdots\cdots$
$\Gamma_{2}$
$\Gamma_{1}$
$\Gamma_{r}$
Figure 3. $\mathcal{F}$-chain with the first curve $\Gamma_{1}$
Definition 2.18.
An irreducible curve $C$ is called a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve) if
  1. (1)
    $C$ is a smooth rational $\mathcal{F}$-invariant curve, and
  2. (2)
    ${\rm Z}(\mathcal{F},C)=1$ (resp. ${\rm Z}(\mathcal{F},C)=2$).
Definition 2.19.
A curve $C$ is called a bad tail if
  1. (1)
    $C$ is a smooth rational irreducible $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$ and $C^{2}\leq-2$;
  2. (2)
    $C$ intersects two (1)(-1)-$\mathcal{F}$-curves whose self-intersections are both 2-2.
(See Figure 4.)
$C$
2-2
2-2
Figure 4. $C$ is a bad tail
Definition 2.20.
An $\mathcal{F}$-dihedral graph is a divisor
$F=\Gamma_{1}+\Gamma_{2}+\cdots+\Gamma_{r},$
where
  1. (1)
    $\Gamma_{1}$ and $\Gamma_{2}$ are (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2;
  2. (2)
    $\Gamma_{3}$ is a bad tail, which is attached to a chain of (2)(-2)-$\mathcal{F}$-curves $\Gamma_{4}+\cdots+\Gamma_{r}$.
(See Figure 5.)
$\Gamma_{1}(-2)$
$\cdots\cdots$
$\Gamma_{2}(-2)$
$\Gamma_{4}$
$\Gamma_{3}$
$\Gamma_{r}$
Figure 5. $\mathcal{F}$-dihedral graph
Definition 2.21 (egl).
An elliptic Gorenstein leaf is either a rational $\mathcal{F}$-invariant curve with a single node, or a cycle of (2)(-2)-$\mathcal{F}$-curves. (See Figure 6.)
or
$\vdots$
Figure 6. An elliptic Gorenstein leaf
Definition 2.22 ($\mathcal{F}$-star graph).
We say that
$F=C+\sum_{i=1}^{s}\Theta_{i}$
is an $\mathcal{F}$-star graph (centered at $C$) if
  1. (1)
    $C$ is a smooth irreducible non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$;
  2. (2)
    for each $i=1,\dots,s$, $\Theta_{i}=\sum_{j=1}^{r(i)}\Gamma_{ij}$ is an $\mathcal{F}$-chain whose first curve $\Gamma_{i1}$ satisfies
    $C\cdot\Theta_{i}=C\cdot\Gamma_{i1}=1.$
We say that an $\mathcal{F}$-star graph is of type $(s;m_{1},\dots,m_{s})$, where
$m_{i}=\det\!\bigl(-\Gamma_{ij}\cdot\Gamma_{ik}\bigr)_{1\leq j,k\leq r(i)}.$
The curve $C$ is called the center of the $\mathcal{F}$-star graph (see Figure 7).
$C$
$\Theta_{1}$
$\cdots\cdots$
$\Theta_{2}$
$\cdots\cdots$
$\Theta_{s}$
$\cdots\cdots$
$\cdots\cdots\cdots\cdots$
Figure 7. $\mathcal{F}$-star graph centered at $C$
Definition 2.23 ($\mathcal{F}$-star chain).
An $\mathcal{F}$-star chain is an $\mathcal{F}$-star graph whose underlying divisor is a chain.
$C$
$\Theta_{1}$
$\Theta_{2}$
$\cdots\cdots$
$\cdots\cdots$
Figure 8. $\mathcal{F}$-star chain centered at $C$, where $\Theta_{1}$ and $\Theta_{2}$ may be empty.
Remark 2.24.
The configurations introduced above arise naturally in the birational geometry of foliated surfaces. They will serve as the basic configurations throughout this paper.
2.8. A negativity criterion on surfaces
Lemma 2.25.
Let $D=\sum_{i=1}^{n}a_{i}C_{i}$ be a $\mathbb{Q}$-divisor such that the intersection matrix
$(C_{i}\cdot C_{j})_{1\leq i,j\leq n}$
is negative definite.
  1. (1)
    If $D\cdot C_{i}\leq 0$ for all $i=1,\dots,n$, then $D\geq 0$.
  2. (2)
    If $E$ is an effective $\mathbb{Q}$-divisor and $(E-D)\cdot C_{j}\leq 0$ for all $j=1,\dots,n$, then $E-D\geq 0$.
Proof.
Both statements follow from [Luc01, Lemma 14.9 and Lemma 14.15]. ∎
3. Negative definite configurations of curves
In this section, we collect several standard facts on negative definite configurations of curves on smooth surfaces, which will be used repeatedly in the study of exceptional divisors arising in resolutions of singularities.
Throughout this section, let $\mathfrak{E}=\bigcup_{i=1}^{n}E_{i}$ be a finite collection of proper irreducible curves on a smooth surface $X$ whose intersection matrix $(E_{i}\cdot E_{j})$ is negative definite. Set $E:=E_{1}+\cdots+E_{n}$.
3.1. Dual graphs and intersection equations
Associated to $\mathfrak{E}$ is its weighted dual graph $G_{E}$, whose vertices correspond to the curves $E_{i}$, with an edge of weight $E_{i}\cdot E_{j}$ between distinct vertices and with vertex weight $E_{i}^{2}$. A subgraph of $G_{E}$ is the dual graph of a divisor $F\subset E$.
Let $D$ be a divisor on $X$, and let $F=C_{1}+\cdots+C_{r}\subset E$ be a divisor whose intersection matrix is negative definite. Then the system of linear equations
$\sum_{i=1}^{r}a_{i}\,C_{i}\cdot C_{j}=D\cdot C_{j},\qquad j=1,\dots,r,$
admits a unique solution $(a_{1},\dots,a_{r})$.
Definition 3.1.
We denote by $M(D,F)$ the unique divisor supported on $F$ satisfying
$M(D,F)\cdot C_{i}=D\cdot C_{i}\quad\text{for all }i=1,\dots,r.$
We say that $F$ (or $G_{F}$) is a $D^{\geq 0}$-graph (resp. $D^{>0}$-graph) if the coefficients of $M(D,F)$ are all non-negative (resp. positive). If $F=C$ is irreducible, then $C$ is a $D^{>0}$-curve if and only if $D\cdot C<0$.
3.2. $D^{>0}$-chains
A divisor $F=C_{1}+\cdots+C_{r}\subset E$ is called a chain if $C_{i}\cdot C_{i+1}=1$ and $C_{i}\cdot C_{j}=0$ for $|i-j|\geq 2$. If moreover $F$ is a $D^{\geq 0}$- (resp. $D^{>0}$-)graph, we call it a $D^{\geq 0}$- (resp. $D^{>0}$-)chain.
Lemma 3.2.
Let $F=C_{1}+\cdots+C_{r}$ be a $D^{>0}$-chain and let $C\subset E$ be a curve not contained in $F$ such that $F\cdot C=C_{r}\cdot C=1$. Then $F+C$ is a $D^{>0}$-chain if and only if
$D\cdot C<M(D,F)\cdot C.$
Proof.
Write
$M(D,F)=\sum_{i=1}^{r}\gamma_{i}C_{i},\qquad M(D,F+C)=\sum_{i=1}^{r+1}\gamma^{\prime}_{i}C_{i},$
where $C_{r+1}:=C$. Then
$\bigl(M(D,F+C)-M(D,F)\bigr)\cdot C_{i}=\begin{cases}0,&i=1,\dots,r,\\ D\cdot C-M(D,F)\cdot C,&i=r+1.\end{cases}$
If $D\cdot C<M(D,F)\cdot C$, then
$\bigl(M(D,F+C)-M(D,F)\bigr)\cdot C_{i}\leq 0\quad\text{for }i=1,\cdots,r+1,$
and hence Lemma 2.25 implies $M(D,F+C)\geq M(D,F)$. Thus $\gamma^{\prime}_{i}>0$ for all $i=1,\dots,r+1$, and $F+C$ is a $D^{>0}$-chain.
Conversely, if $D\cdot C\geq M(D,F)\cdot C$, then $\gamma^{\prime}_{r+1}\leq 0$, so $F+C$ is not a $D^{>0}$-chain. ∎
Remark 3.3.
An analogous statement holds for $D^{\geq 0}$-chains with the inequality $D\cdot C\leq M(D,F)\cdot C$.
For later estimates along chains, we introduce the following determinants associated to $F$.
(3.1) $\begin{cases}\mu_{0}:=\lambda_{r+1}=1,\qquad\mu_{1}:=\lambda_{r}=1,\\ \mu_{k}:=\det(-C_{i}\cdot C_{j})_{1\leq i,j\leq k-1},&k=2,\dots,r+1,\\ \lambda_{l}:=\det(-C_{i}\cdot C_{j})_{l+1\leq i,j\leq r},&l=0,\dots,r-1,\\ n:=\lambda_{0}=\mu_{r+1}.\end{cases}$
Lemma 3.4.
  1. (1)
    For $i=1,\dots,r$, we have
    $\lambda_{i-1}-e_{i}\lambda_{i}+\lambda_{i+1}=0,\qquad\mu_{i-1}-e_{i}\mu_{i}+\mu_{i+1}=0,$
    where $e_{i}:=-C_{i}^{2}$.
  2. (2)
    For $i<j$, the quantity $\lambda_{i}\mu_{j}-\lambda_{j}\mu_{i}$ is a positive multiple of $n$. In particular,
    $\lambda_{i}\mu_{i+1}-\lambda_{i+1}\mu_{i}=n.$
Proof.
This follows from a straightforward computation. ∎
Since the intersection matrix $(C_{i}\cdot C_{j})_{1\leq i,j\leq r}$ is negative definite, for each $i$ there exists a unique effective $\mathbb{Q}$-divisor $M_{i}(F)$ supported on $F$ such that
$M_{i}(F)\cdot C_{i}=-1,\qquad M_{i}(F)\cdot C_{j}=0\quad(j\neq i).$
A direct computation yields
$M_{i}(F)=\frac{\lambda_{i}}{n}\sum_{k=1}^{i}\mu_{k}C_{k}+\frac{\mu_{i}}{n}\sum_{k=i+1}^{r}\lambda_{k}C_{k}.$
Consequently,
$M(D,F)=\sum_{i=1}^{r}(-D\cdot C_{i})\,M_{i}(F).$
Moreover, we have the following lemma:
Lemma 3.5.
Writing $M(D,F)=\sum_{i=1}^{r}\gamma_{i}C_{i}$, we have
(3.2) $\gamma_{i}=\frac{\lambda_{i}}{n}\sum_{k=1}^{i}\mu_{k}\,(-D\cdot C_{k})+\frac{\mu_{i}}{n}\sum_{k=i+1}^{r}\lambda_{k}\,(-D\cdot C_{k}),\qquad i=1,\dots,r.$
Lemma 3.6.
Assume that $D\cdot C_{i}\geq 0$ for $i\geq 2$. Then $F$ is a $D^{>0}$-chain if and only if
$\sum_{k=1}^{r}\mu_{k}\,D\cdot C_{k}<0.$
In this case, for any $k\geq 2$,
$D\cdot C_{k}<\frac{-D\cdot C_{1}}{\mu_{k}}.$
Proof.
Using (3.2), we may write
$\gamma_{i}=\frac{\lambda_{i}}{n}\sum_{k=1}^{r}\mu_{k}\,(-D\cdot C_{k})-\sum_{k=i+1}^{r}\frac{\lambda_{i}\mu_{k}-\lambda_{k}\mu_{i}}{n}\,(-D\cdot C_{k}).$
By Lemma 3.4, we have $\lambda_{i}\mu_{k}-\lambda_{k}\mu_{i}\geq n$ for $k>i$. Since $D\cdot C_{k}\geq 0$ for $k\geq 2$, it follows that
$\gamma_{i}\geq\frac{\lambda_{i}}{n}\sum_{k=1}^{r}\mu_{k}\,(-D\cdot C_{k})=\lambda_{i}\gamma_{r}.$
Hence $F$ is a $D^{>0}$-chain if and only if $\gamma_{r}>0$, which is equivalent to $\sum_{k=1}^{r}\mu_{k}\,D\cdot C_{k}<0$.
In this case, for $i\geq 2$,
$\mu_{i}\,D\cdot C_{i}\leq\sum_{k=2}^{r}\mu_{k}\,D\cdot C_{k}<\mu_{i}\,(-D\cdot C_{1}),$
and the desired inequality follows. ∎
3.3. Special $D^{>0}$-chains
Definition 3.7.
A $D^{>0}$- (resp. $D^{\geq 0}$-)chain $F=C_{1}+\cdots+C_{r}$ is called special if
$D\cdot C_{1}<0\quad\text{and}\quad D\cdot C_{i}\geq 0\ \text{for }i\geq 2.$
The next proposition provides a numerical criterion for positivity along special $D^{>0}$-chains, which will be crucial in the classification.
Proposition 3.8.
Let $F=C_{1}+\cdots+C_{r}$ be a special $D^{>0}$-chain. Write $M(D,F)=\sum_{i=1}^{r}\gamma_{i}C_{i}$. Then:
  • (1)
    $\gamma_{k}\leq\frac{(-D\cdot C_{1})\lambda_{k}}{n},\qquad k=1,\dots,r.$
  • (2)
    For each $k=1,\dots,r$, the sub-chain $F_{k}=C_{1}+\cdots+C_{k}$ is a special $D^{>0}$-chain, and
    $0<M(D,F_{1})<M(D,F_{2})<\cdots<M(D,F_{r})=M(D,F).$
    In particular,
    $D\cdot C_{k}<M(D,F_{k-1})\cdot C_{k}\leq\frac{-D\cdot C_{1}}{\mu_{k}},\qquad k\geq 2,$
    where $\mu_{k}=\det(-C_{i}\cdot C_{j})_{1\leq i,j\leq k-1}$.
Proof.
(1) follows directly from (3.2). (2) follows from Lemma 3.2 and Lemma 3.6. ∎
4. Classification of $\epsilon$-adjoint lc singularities with $\epsilon\in(0,1/3)$
Let $(Y,\mathcal{G})$ be a foliated pair on a normal surface $Y$ and let $p\in Y$ be a singular point of $(Y,\mathcal{G})$. Let
$\pi:(X,\mathcal{F})\longrightarrow(Y,\mathcal{G})$
be the minimal resolution of $(Y,\mathcal{G})$ at $p$. Denote by $\mathfrak{E}=\bigcup_{i\in I}E_{i}$ the set of $\pi$-exceptional curves and set
$E:=\sum_{i\in I}E_{i}.$
We can write
$K_{(X,\mathcal{F},E),\epsilon}=\pi^{*}K_{(Y,\mathcal{G}),\epsilon}+\sum_{i}a_{i}E_{i}.$
Intersecting both sides with $E_{j}$, we obtain the system of linear equations
(4.1) $\sum_{i\in I}a_{i}\,E_{i}\cdot E_{j}=K_{(X,\mathcal{F},E),\epsilon}\cdot E_{j},\qquad\text{for all $j\in I$}.$
It follows that $p$ is an $\epsilon$-adjoint log canonical (resp. $\epsilon$-adjoint klt) singularity if and only if
$a_{i}\geq 0\quad(\text{resp.\ }a_{i}>0)\quad\text{for all }i.$
Our goal is to classify all possible configurations of the exceptional divisor $E$ arising from $\epsilon$-adjoint lc (resp. klt) singularities. Equivalently, this amounts to determining all possible dual graphs of $E$ which are $K^{\geq 0}$-graphs (resp. $K^{>0}$-graphs), where
$K:=K_{(X,\mathcal{F},E),\epsilon},\qquad\epsilon\in(0,\tfrac{1}{3}).$
4.1. Outline of the classification
The classification proceeds by a systematic analysis of the intersection numbers $K\cdot C$ for irreducible components $C\subset E$.
We first study components satisfying $K\cdot C\leq 0$. Such curves are shown to be extremely restricted: they are either invariant or non-invariant curves of specific numerical types, which are completely classified in Lemmas 4.2 and 4.3. These curves form the basic building blocks of the exceptional divisor.
Starting from components with $K\cdot C<0$, we then introduce certain linear configurations, called special $K^{>0}$-chains. These chains play a central role in the analysis. A key feature is their rigidity: only finitely many numerical types can occur, and these are explicitly classified in Proposition 4.5.
Once all special chains are identified, we subtract their contribution from $K$ and consider the modified divisor
$K_{1}:=K-\sum_{F}M(K,F),$
where the sum runs over all special $K^{>0}$-chains $F\subset E$. To simplify the analysis and to avoid interactions between different chains, we impose a technical assumption ()(*), under which all special $K^{>0}$-chains are pairwise disjoint. (The case where two special chains overlap is treated separately in Proposition 4.10.) Under this assumption, the remaining components of $E$ can be analyzed according to the sign of $K_{1}\cdot C$.
We then classify the irreducible components that are not contained in any special $K^{>0}$-chain according to whether they are $\mathcal{F}$-invariant and to the sign of $K_{1}\cdot C$. This leads to strong numerical restrictions on their self-intersection numbers and valencies in the dual graph. Combining these results yields a complete list of possible configurations for $E$ under assumption ()(*), stated in Proposition 4.13.
Finally, we show that this list is exhaustive for $\epsilon$-adjoint log canonical singularities. This leads to a finite classification of exceptional divisors, summarized in Theorem 4.14.
4.2. $K$-non-positive curves
Definition 4.1.
Let $C$ be an irreducible curve contained in $E$. We define the contributions coming from non-invariant and invariant components by
(4.2) $d_{1}(C):=\left(E^{\rm n\!-\!inv}-\iota(C)C\right)\cdot C,\qquad d_{2}(C):=\left(E^{\rm inv}-(1-\iota(C))C\right)\cdot C,$
and set
$d(C):=d_{1}(C)+d_{2}(C)=(E-C)\cdot C.$
Lemma 4.2.
Let $C\subset E$ be an irreducible $\mathcal{F}$-invariant curve. Then
(4.3) $KC=(1+\epsilon)\bigl(2p_{a}(C)-2+d(C)\bigr)+{\rm Z}(\mathcal{F},C)-d_{2}(C).$
Moreover, $KC\leq 0$ if and only if $C$ belongs to one of the following cases:
  1. (1)
    $d(C)=0$.
    1. (a)
      $C$ is an elliptic Gorenstein leaf. In this case, $KC=0$.
    2. (b)
      $C$ is a (2)(-2)-$\mathcal{F}$-curve. In this case, $KC=-2\epsilon$.
    3. (c)
      $C$ is a (1)(-1)-$\mathcal{F}$-curve. In this case, $KC=-(1+2\epsilon)$.
  2. (2)
    $d(C)=1$.
    1. (a)
      $C$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$. In this case, $KC=-(1+\epsilon)$.
    2. (b)
      $C$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=1$ and $d_{2}=0$. In this case, $KC=-\epsilon$.
    3. (c)
      $C$ is a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$. In this case, $KC=-\epsilon$.
  3. (3)
    $d(C)=2$.
    1. (a)
      $C$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$. In this case, $KC=0$.
    2. (b)
      $C$ is a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=2$. In this case, $KC=0$.
Proof.
We compute
$KC$ $=(K_{\mathcal{F}}+E^{\rm n\!-\!inv})\cdot C+\epsilon(K_{X}+E)\cdot C$
$=K_{\mathcal{F}}\cdot C+\epsilon(K_{X}+C)\cdot C+E^{\rm n\!-\!inv}\cdot C+\epsilon(E-C)\cdot C$
$=K_{\mathcal{F}}\cdot C+\epsilon(2p_{a}(C)-2)+d_{1}(C)+\epsilon d(C)$
$=K_{\mathcal{F}}\cdot C+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d(C)-d_{2}(C).$
By Proposition 2.8,
$K_{\mathcal{F}}\cdot C=2p_{a}(C)-2+{\rm Z}(\mathcal{F},C).$
Thus,
$KC=(1+\epsilon)\bigl(2p_{a}(C)-2+d(C)\bigr)+{\rm Z}(\mathcal{F},C)-d_{2}(C),$
where ${\rm Z}(\mathcal{F},C)-d_{2}(C)\geq 0$.
If $p_{a}(C)\geq 1$, then $KC\geq 0$. Moreover, $KC=0$ if and only if $p_{a}(C)=1$ and $K_{\mathcal{F}}\cdot C=d(C)=0$. In this case, $C$ is an elliptic Gorenstein leaf. This is case (1-a).
Next we assume $p_{a}(C)=0$. Then
$KC=(1+\epsilon)(d(C)-2)+{\rm Z}(\mathcal{F},C)-d_{2}(C),$
and $KC\leq 0$ implies $d(C)\leq 2$.
  • If $d(C)=0$, then $d_{2}(C)=0$ and
    $KC=-2(1+\epsilon)+{\rm Z}(\mathcal{F},C).$
    Since $C^{2}<0$ and $p_{a}(C)=0$, we have ${\rm Z}(\mathcal{F},C)\geq 1$, by Proposition 2.8 and Lemma 2.9.
    • If ${\rm Z}(\mathcal{F},C)\geq 3$, then $KC\geq 1-2\epsilon>0$.
    • If ${\rm Z}(\mathcal{F},C)=2$, then $C$ is a (2)(-2)-$\mathcal{F}$-curve and $KC=-2\epsilon$. This is case (1-b).
    • If ${\rm Z}(\mathcal{F},C)=1$, then $C$ is a (1)(-1)-$\mathcal{F}$-curve and $KC=-(1+2\epsilon)$. This is case (1-c).
  • If $d(C)=1$, then
    $KC=-(1+\epsilon)+{\rm Z}(\mathcal{F},C)-d_{2}(C).$
    • If ${\rm Z}(\mathcal{F},C)-d_{2}(C)\geq 2$, then $KC\geq 1-\epsilon>0$.
    • If $d_{2}(C)=0$ and ${\rm Z}(\mathcal{F},C)=1$, then $C$ is a (1)(-1)-$\mathcal{F}$-curve with $KC=-\epsilon$. This is case (2-b).
    • If $d_{2}(C)=1$ and ${\rm Z}(\mathcal{F},C)=1$, then $C$ is a (1)(-1)-$\mathcal{F}$-curve with $KC=-(1+\epsilon)$. This is case (2-a).
    • If $d_{2}(C)=1$ and ${\rm Z}(\mathcal{F},C)=2$, then $C$ is a (2)(-2)-$\mathcal{F}$-curve with $KC=-\epsilon$. This is case (2-c).
  • If $d(C)=2$, then $KC={\rm Z}(\mathcal{F},C)-d_{2}(C)\geq 0$. Moreover, $KC=0$ if and only if ${\rm Z}(\mathcal{F},C)=d_{2}(C)\in\{1,2\}$.
    • If ${\rm Z}(\mathcal{F},C)=d_{2}(C)=1$, then $C$ is a (1)(-1)-$\mathcal{F}$-curve. This is case (3-a).
    • If ${\rm Z}(\mathcal{F},C)=d_{2}(C)=2$, then $C$ is a (2)(-2)-$\mathcal{F}$-curve. This is case (3-b).
We now turn to the non-$\mathcal{F}$-invariant case.
Lemma 4.3.
Let $C\subset E$ be an irreducible non-$\mathcal{F}$-invariant curve. Then
(4.4) $KC={\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+d_{1}(C)+\epsilon d(C).$
Moreover, $KC\leq 0$ if and only if ${\rm tang}(\mathcal{F},C)=0$, and $C$ is one of the following explicitly listed curves:
  1. (1)
    $d(C)=0$.
    1. (a)
      $C$ is a smooth elliptic curve. In this case, $KC=0$.
    2. (b)
      $C$ is a smooth rational curve. In this case, $KC=-2\epsilon$.
  2. (2)
    $d(C)=1$. $C$ is a smooth rational curve with $d_{1}(C)=0$ and $d_{2}(C)=1$. In this case, $KC=-\epsilon$.
  3. (3)
    $d(C)=2$. $C$ is a smooth rational curve with $d_{1}(C)=0$ and $d_{2}(C)=2$. In this case, $KC=0$.
Proof.
We compute
$KC$ $=(K_{\mathcal{F}}+E^{\rm n\!-\!inv})\cdot C+\epsilon(K_{X}+E)\cdot C$
$=(K_{\mathcal{F}}+C)\cdot C+\epsilon(K_{X}+C)\cdot C+(E^{\rm n\!-\!inv}-C)\cdot C+\epsilon(E-C)\cdot C.$
By Proposition 2.6, the genus formula, and the definitions of $d_{1}(C)$ and $d(C)$, this can be rewritten as
$KC={\rm tang}(\mathcal{F},C)+\epsilon\bigl(2p_{a}(C)-2\bigr)+d_{1}(C)+\epsilon d(C).$
If $p_{a}(C)\geq 1$, then $KC\geq 0$. Moreover, $KC=0$ if and only if $p_{a}(C)=1$ and ${\rm tang}(\mathcal{F},C)=d(C)=0$. In this case, $C$ is a smooth elliptic curve. This is case (1-a).
Next we assume $p_{a}(C)=0$. If ${\rm tang}(\mathcal{F},C)+d_{1}(C)\geq 1$, then $KC\geq 1-2\epsilon>0$. Thus, we may assume ${\rm tang}(\mathcal{F},C)=d_{1}(C)=0$, in which case
$KC=(d(C)-2)\epsilon.$
Then $KC\leq 0$ implies case (1-b), (2) or (3). ∎
4.3. Special $K^{>0}$-chains in $E$
Let $C$ be an irreducible component of $E$. By Lemmas 4.2 and 4.3, any irreducible component $C\subset E$ with $KC<0$ satisfies $d(C)\leq 1$. If $d(C)=0$, then necessarily $E=C$. Otherwise, when $d(C)=1$, such a curve intersects ECE-C in a unique point, and hence naturally gives rise to a linear configuration inside $E$. This motivates the introduction of the following notion.
Definition 4.4.
Let $F=C_{1}+\cdots+C_{r}$ be a chain of irreducible curves in $E$. We say that $F$ is a special $K^{>0}$-chain (resp. a special $K^{\geq 0}$-chain) if $F$ is a $K^{>0}$-graph (resp. $K^{\geq 0}$-graph) satisfying
$KC_{1}<0,\qquad KC_{i}\geq 0,\quad i\geq 2.$
We call $C_{1}$ the first curve of $F$.
Proposition 4.5.
Suppose $E$ has at least two components. Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{>0}$-chain with first curve $C_{1}$. In all cases below, we have
$d(C_{1})=1,\qquad d(C_{i})=2\ \text{and}\ KC_{i}=0\ \text{for all }i\geq 2.$
Moreover, $F$ belongs to one of the following configurations:
  • (A)
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve, and $C_{i}$ is a (2)(-2)-$\mathcal{F}$-curve for all $i\geq 2$.
  • (B)
    Each $C_{i}$ is a (2)(-2)-$\mathcal{F}$-curve.
  • (C)
    $C_{i}$ is a (2)(-2)-$\mathcal{F}$-curve for all $i<r$, and $C_{r}$ is a (1)(-1)-$\mathcal{F}$-curve.
  • (D)
    There exists some $k\in[1,r]$ such that $C_{k}$ is a non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{k})=0$. Moreover, $C_{k-1}$ (if $k\geq 2$) and $C_{k+1}$ (if $k\leq r-1$) are (1)(-1)-$\mathcal{F}$-curves, and all remaining $C_{i}$ are (2)(-2)-$\mathcal{F}$-curves.
In addition,
[…]
Proof.
Since $KC_{1}<0$ and $d(C_{1})\geq 1$, by Lemmas 4.2 and 4.3, the curve $C_{1}$ must fall into one of the following cases:
  • (i)
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$. In this case, $KC_{1}=-(1+\epsilon)$.
  • (ii)
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=1$ and $d_{2}=0$. In this case, $KC_{1}=-\epsilon$.
  • (iii)
    $C_{1}$ is a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$. In this case, $KC_{1}=-\epsilon$.
  • (iv)
    $C_{1}$ is a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{1})=0$, $d_{1}=0$ and $d_{2}=1$. In this case, $KC_{1}=-\epsilon$.
If $r=1$, then $F=C_{1}$ and the statement follows directly from the above list. Hence we assume $r\geq 2$.
Since $KC_{i}\geq 0$ for all $i\geq 2$, Proposition 3.8 implies
$KC_{k}<\frac{-KC_{1}}{\mu_{k}},\qquad k\geq 2,$
where $\mu_{k}=\det\big(-C_{i}C_{j}\big)_{1\leq i,j\leq k-1}$.
Cases (ii)–(iv). In these cases $KC_{1}=-\epsilon$, hence $0\leq KC_{i}<\epsilon$ for all $i\geq 2$. By (4.3) and (4.4), for each $i\geq 2$ we have $KC_{i}=0$, and $C_{i}$ must be one of the following:
  • a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$;
  • a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=2$;
  • a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{i})=0$, $d_{1}=0$ and $d_{2}=2$.
A simple combinatorial analysis yields the types (B), (C), and (D), and any additional case is excluded by the separatrix theorem.
Case (i). Here $KC_{1}=-(1+\epsilon)$ and $d(C_{1})=d_{2}(C_{1})=1$. In particular, we have $C_{1}^{2}\leq-2$, and hence
$0\leq KC_{k}<\frac{1+\epsilon}{2}<1-\epsilon,\qquad k\geq 2.$
Since $d_{2}(C_{1})=1$, the curve $C_{2}$ is $\mathcal{F}$-invariant. Moreover, as $0\leq KC_{2}<1-\epsilon$ and $d_{2}(C_{2})\geq 1$, it follows from (4.3) that $d(C_{2})=d_{2}(C_{2})=2$ and $KC_{2}=0$. Thus $C_{2}$ is either a (2)(-2)-$\mathcal{F}$-curve or a (1)(-1)-$\mathcal{F}$-curve. The latter is excluded by the separatrix theorem.
By the same argument, each curve $C_{i}$ with $i\geq 3$ must be a (2)(-2)-$\mathcal{F}$-curve satisfying $d(C_{i})=d_{2}(C_{i})=2$ and $KC_{i}=0$. Therefore, the chain $F$ is of type (A). ∎
Remark 4.6.
Types (A) and (C) are both $\mathcal{F}$-chains. However, when viewed as special $K^{>0}$-chains, they are distinct.
The following result will be used repeatedly to control how a special $K^{\geq 0}$-chain can meet the rest of the exceptional divisor.
Corollary 4.7.
Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{\geq 0}$-chain in $E$ with $KC_{1}<0$. If $C$ is an irreducible curve in $E\setminus F$ such that $F\cdot C=C_{r}\cdot C=1$, then
[…]
where $n=\det(-C_{i}C_{j})_{1\leq i,j\leq r}$. Moreover, such a curve $C$ is non-$\mathcal{F}$-invariant if and only if $F$ is of type (C).
4.4. Classification of the exceptional divisor
Let $G_{E}$ denote the dual graph of $E$, which is connected. We classify all possible configurations of $E$ according to the structure of $G_{E}$ and the distribution of the intersection numbers $KC$.
4.4.1. Case I: $E$ irreducible
Proposition 4.8.
If $G_{E}=\{[C]\}$, then $C$ belongs to one of the following cases:
  • (1)
    $C$ is an irreducible elliptic Gorenstein leaf. In this case, $KC=0$.
  • (2)
    $C$ is a smooth elliptic non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$. In this case, $KC=0$.
  • (3)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve. In this case, $KC=-2\epsilon$.
  • (4)
    $C$ is a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$. In this case, $KC=-2\epsilon$.
  • (5)
    $C$ is a (1)(-1)-$\mathcal{F}$-curve. In this case, $KC=-(1+2\epsilon)$.
Proof.
It follows directly from Lemma 4.2 and Lemma 4.3. ∎
4.4.2. Case II: $E$ is a $K^{=0}$-graph
We next consider the case where all components of $E$ have zero intersection with the adjoint divisor.
Proposition 4.9.
If $E$ is a $K^{=0}$-graph, that is, $KC=0$ for any component $C$ of $E$, then $E$ belongs to one of the following cases:
  1. (1)
    $E=C$ is an irreducible curve, where $C$ is an irreducible elliptic Gorenstein leaf, or a smooth elliptic non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$.
  2. (2)
    $E$ is a cycle of (2)(-2)-$\mathcal{F}$-curves.
Proof.
If $E$ consists of a single component $C$, then $KC=0$ and $d(C)=0$. Thus $C$ falls into case (1-a) of Lemma 4.2 or Lemma 4.3. This is case (1).
Assume now that $E$ has at least two components. Then $d(C)\geq 1$ for any component $C$ of $E$. Since $KC=0$, Lemma 4.2 and Lemma 4.3 imply that each $C$ is a smooth rational curve with $d(C)=2$. Hence $E=C_{1}+\cdots+C_{r}$ is a cycle of smooth rational curves.
By Lemma 4.2 and Lemma 4.3, each $C_{i}$ must belong to one of the following types:
  • (i)
    a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$;
  • (ii)
    a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=2$;
  • (iii)
    a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$, $d_{1}=0$, and $d_{2}=2$.
A direct analysis using the separatrix theorem shows that cases (i) and (iii) cannot occur. Therefore, each $C_{i}$ is a (2)(-2)-$\mathcal{F}$-curve, and $E$ is a cycle of such curves. This is case (2). ∎
4.4.3. The presence of special $K^{>0}$-chains
We now assume that $E$ does not belong to Case I or Case II. Then $E$ necessarily contains at least one special $K^{>0}$-chain.
Proposition 4.10.
Suppose that there exist two special $K^{>0}$-chains $F_{1}$ and $F_{2}$ with distinct first curves, such that they share a common component or satisfy $F_{1}F_{2}>0$. Then $E=C_{1}+\cdots+C_{r}$ is a chain belonging to one of the following cases:
  • (1)
    $E$ is an $\mathcal{F}$-chain.
  • (2)
    $E$ is a chain of (2)(-2)-$\mathcal{F}$-curves.
  • (3)
    There exists some $k\in[1,r]$ such that $C_{k}$ is a non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{k})=0$, and both $C_{1}+\cdots+C_{k-1}$ and $C_{k+1}+\cdots+C_{r}$ are $\mathcal{F}$-chains whose first curves are $C_{k-1}$ and $C_{k+1}$, respectively. (See Figure 8.)
Proof.
By Proposition 4.5, a simple combinatorial analysis yields the desired cases, and the separatrix theorem excludes all others. ∎
To separate the remaining cases, we impose the following technical assumption.
    $E$ does not belong to Case I or Case II, and any two special $K^{>0}$-chains with distinct first curves are pairwise disjoint.
Under assumption ()(*), maximal special $K^{>0}$-chains can be defined. We introduce the following notation.
  • Let $\{\Theta_{1},\dots,\Theta_{s}\}$ denote the set of all maximal special $K^{>0}$-chains of type (A) contained in $E$.
  • Let $\{\Xi_{1},\dots,\Xi_{t}\}$ denote the set of all maximal special $K^{>0}$-chains of type (B), (C), or (D) contained in $E$.
Define
(4.5) $K_{1}:=K-\sum_{i=1}^{s}M(K,\Theta_{i})-\sum_{j=1}^{t}M(K,\Xi_{j}).$
It is clear that:
  • $E$ is a $(K_{1})^{\geq 0}$-graph;
  • for any component $C$ of $\Theta_{i}$ (resp. $\Xi_{j}$), one has $K_{1}C=0$.
Lemma 4.11.
Under assumption ()(*), let $C$ be an irreducible non-$\mathcal{F}$-invariant curve contained in $E$, and assume that $C$ is not contained in any special $K^{>0}$-chain. If $K_{1}C\leq 0$, then one of the following holds:
  1. (1)
    $C$ is a smooth rational curve with ${\rm tang}(\mathcal{F},C)=0$, and $E$ is an $\mathcal{F}$-star graph (cf. Definition 2.22) of type $(l;m_{1},\dots,m_{l})$ centered at $C$, where $(l;m_{1},\dots,m_{l})$ is one of the following:
    1. (a)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, $(2,3,4)$, or $(2,3,5)$;
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, or $(2,4,4)$, or $(3,3,3)$;
    3. (c)
      $l=4$ and $(m_{1},\dots,m_{4})=(2,2,2,2)$.
    (In cases (b) and (c), we have $K_{1}C=0$.)
  2. (2)
    $C$ is a smooth rational curve with ${\rm tang}(\mathcal{F},C)=0$ meeting transversely two (1)(-1)-$\mathcal{F}$-curves $\Gamma_{1}$ and $\Gamma_{2}$ whose self-intersection are 2-2, where
    $d(\Gamma_{1})=d(\Gamma_{2})=1,\qquad d(C)=d_{2}(C)=3.$
    In this case, $K_{1}C=0$.
  3. (3)
    $C$ is a smooth rational curve with
    ${\rm tang}(\mathcal{F},C)=0,\qquad d_{1}(C)=0,\qquad d_{2}(C)=2,$
    and is disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
Proof.
In this case we have $d(C)\geq 1$. If $FC>0$ for some special $K^{>0}$-chain $F$, then $F$ must be of type (C) by Proposition 4.5 and Corollary 4.7.
Without loss of generality, we assume that $C$ meets transversely the chains $\Xi_{1},\dots,\Xi_{l}$. Let $m_{j}$ denote the absolute value of the determinant of the intersection matrix of $\Xi_{j}$. Since each $\Xi_{j}$ is of type (C), for $j\leq l$, we have $m_{j}\geq 2$. By (4.4), we obtain
$KC={\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d_{1}(C)+\epsilon d_{2}(C).$
Thus,
[…]
$={\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d_{1}+\epsilon(d_{2}-l)+\sum_{j=1}^{l}\epsilon\!\left(1-\frac{1}{m_{j}}\right)$
$\geq{\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d_{1}+\epsilon(d_{2}-l)+\frac{\epsilon l}{2}$
$={\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d_{1}+\frac{\epsilon(d_{2}-l)}{2}+\frac{\epsilon d_{2}}{2},$
where $d_{2}\geq l$ and $m_{j}\geq 2$.
Note that if ${\rm tang}(\mathcal{F},C)\geq 1$ or $p_{a}(C)\geq 1$ or $d_{1}\geq 1$, then $K_{1}C>0$, where we use $d=d_{1}+d_{2}\geq 1$ and $\epsilon\in(0,1/3)$. Hence we may assume
${\rm tang}(\mathcal{F},C)=p_{a}(C)=d_{1}(C)=0.$
In this case,
$K_{1}C=\epsilon(d_{2}-l-2)+\sum_{j=1}^{l}\epsilon\!\left(1-\frac{1}{m_{j}}\right)\geq\epsilon(d_{2}-l-2)+\frac{\epsilon l}{2}.$
  • If $d_{2}-l\geq 3$, or $d_{2}-l=2$ with $l\geq 1$, or $d_{2}-l=1$ with $l\geq 3$, or $d_{2}-l=0$ with $l\geq 5$, then $K_{1}C>0$.
  • If $d_{2}-l=2$ with $l=0$, then $K_{1}C=0$. Here $C$ is disjoint from all special $K^{>0}$-chains and $d=d_{2}=2$. This is case (3).
  • If $d_{2}-l=1$ with $l=2$, then
    $K_{1}C=\epsilon\!\left(1-\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\geq 0,$
    and equality holds iff $(m_{1},m_{2})=(2,2)$. This is case (2).
  • If $d_{2}-l=1$ with $l=1$, then $d=d_{2}=2$ and $KC=0$. In this situation $\Xi_{1}+C$ forms a special $K^{>0}$-chain, contradicting the maximality of $\Xi_{1}$.
  • If $d_{2}-l=1$ with $l=0$, or $d_{2}-l=0$ with $l=1$, then $d=1$ and $KC<0$. Thus $C$ would be the first curve of a special $K^{>0}$-chain, contradicting our assumption.
  • If $d_{2}-l=0$ with $l=0$, then $d=d_{2}=0$, contradicting our assumption.
  • If $d_{2}-l=0$ with $l=2$, then $d=d_{2}=2$ and $KC=0$. In this situation $\Xi_{1}+C$ forms a special $K^{>0}$-chain, contradicting the maximality of $\Xi_{1}$.
  • If $d_{2}-l=0$ with $l\in\{3,4\}$, then
    $K_{1}C=-2\epsilon+\sum_{i=1}^{l}\epsilon(1-\frac{1}{m_{j}})=\epsilon\cdot\left(-2+l-\sum_{i=1}^{l}\frac{1}{m_{j}}\right)$
    Hence, the condition $K_{1}C\leq 0$ is equivalent to
    $\frac{1}{m_{1}}+\cdots+\frac{1}{m_{l}}\geq l-2.$
    In this case $E=C+\Xi_{1}+\cdots+\Xi_{l}$ is an $\mathcal{F}$-star graph, giving case (1).
Lemma 4.12.
Under the assumption ()(*), let $C$ be an irreducible $\mathcal{F}$-invariant curve contained in $E$, and assume that $C$ is not contained in any special $K^{>0}$-chain. If $K_{1}C\leq 0$, then one of the following holds:
  1. (1)
    $E=\Theta_{1}+\Theta_{2}+C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, meeting transversely two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of the following:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2;
    2. (b)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
    3. (c)
      $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
  2. (2)
    $E=C+\Theta_{1}+\Theta_{2}+\Xi_{1}$ is a fork centered at $C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, $\Theta_{1}$ and $\Theta_{2}$ are $\mathcal{F}$-chains, and $\Xi_{1}$ is a chain of (2)(-2)-$\mathcal{F}$-curves. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of the following:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2;
    2. (b)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
    3. (c)
      $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
  3. (3)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve that meets transversely two (1)(-1)-$\mathcal{F}$-curves $\Gamma_{1}$ and $\Gamma_{2}$, where
    $d(\Gamma_{1})=d(\Gamma_{2})=1,\qquad d(C)=d_{2}(C)=3.$
    In this case, $K_{1}C=0$.
  4. (4)
    $C$ is a (1)(-1)-$\mathcal{F}$-curve with
    $d_{1}(C)=d_{2}(C)=1,$
    and is disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
  5. (5)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve with
    $d(C)=d_{2}(C)=2,$
    and is disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
In particular, cases (2-b), (2-c), (3-b) and (3-c) cannot occur if $\epsilon<\tfrac{1}{5}$.
Proof.
Without loss of generality, we assume $C$ meets transversely with $\Theta_{1},\cdots,\Theta_{k}$ at $p_{1},\cdots,p_{k}$ respectively, and $C$ meets transversely with $\Xi_{1},\cdots,\Xi_{l}$ at $q_{1},\cdots,q_{l}$ respectively. Let
$V:=\sum_{i=1}^{k}M(K,\Theta_{i})+\sum_{j=1}^{l}M(K,\Xi_{i}).$
By Corollary 4.7, we have
(4.6) $VC=\sum_{i=1}^{k}\frac{1+\epsilon}{n_{i}}+\sum_{j=1}^{l}\frac{\epsilon}{m_{j}}$
where $n_{i}$ (resp. $m_{i}$) denotes the absolute value of the determinant of the intersection matrix of $\Theta_{i}$ (resp. $\Xi_{j}$).
By (4.3),
$KC=(1+\epsilon)(2p_{a}(C)-2+d(C))+{\rm Z}(\mathcal{F},C)-d_{2}(C).$
Then
$K_{1}C$ $=KC-VC$
$=(1+\epsilon)(2p_{a}(C)-2+d-k-l)+l+{\rm Z}(\mathcal{F},C)-d_{2}$
$\qquad+\sum_{i=1}^{k}(1+\epsilon)(1-\frac{1}{n_{i}})+\sum_{j=1}^{l}\epsilon(1-\frac{1}{m_{j}}),$
where $d-k-l\geq 0$ and ${\rm Z}(\mathcal{F},C)-d_{2}\geq 0$.
If $p_{a}(C)\geq 1$, then $K_{1}C\leq 0$ implies $k=l=d=0$, which is a contradiction with the assumption ()(*).
Next we assume $p_{a}(C)=0$. Then the dual graph of $C+\sum_{i=1}^{k}\Theta_{i}+\sum_{j=1}^{l}\Xi_{j}$ is a tree and
$K_{1}C$ $=(1+\epsilon)(d-k-l-2)+l+{\rm Z}(\mathcal{F},C)-d_{2}$
$\qquad+\sum_{i=1}^{k}(1+\epsilon)(1-\frac{1}{n_{i}})+\sum_{j=1}^{l}\epsilon(1-\frac{1}{m_{j}})$
$\geq(1+\epsilon)(d-k-l-2)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\frac{k(1+\epsilon)}{2}$
Suppose $d-k-l\geq 3$. Then $K_{1}C>0$.
Suppose dkl=2d-k-l=2. Then $K_{1}C\leq 0$ implies $k=l=0$, $d=2$, and ${\rm Z}(\mathcal{F},C)=d_{2}\in\{1,2\}$. Thus, $C$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$ or a (2)(-2)-$\mathcal{F}$-curve with $d=d_{2}=2$, and it is disjoint from all special $K^{>0}$-chains. This is case (4) or (5). In this case, $K_{1}C=KC=0$.
Suppose dkl=1d-k-l=1. Then
$K_{1}C$ $=-(1+\epsilon)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\sum_{i=1}^{k}(1+\epsilon)(1-\frac{1}{n_{i}})+\sum_{j=1}^{l}\epsilon(1-\frac{1}{m_{j}})$
$\geq-(1+\epsilon)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\tfrac{k}{2}(1+\epsilon).$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}\geq 2$, then
    $K_{1}C\geq-(1+\epsilon)+2=1-\epsilon>0.$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=1$ and $k\geq 1$, then
    $K_{1}C\geq-(1+\epsilon)+1+\frac{1}{2}(1+\epsilon)=\frac{1}{2}(1-\epsilon)>0.$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=1$ and $k=0$, then $d=l+1\in\{1,2\}$.
    • If $l=0$, then $d=1$ and ${\rm Z}(\mathcal{F},C)-d_{2}=1$. Thus, $KC=-\epsilon$. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
    • If $l=1$, then $d=2$ and ${\rm Z}(\mathcal{F},C)-d_{2}=0$. Thus, $KC=0$. In this case, $\Xi_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Xi_{1}$.
  • If $l={\rm Z}(\mathcal{F},C)-d_{2}=0$, then
    $K_{1}C=-(1+\epsilon)+\sum_{i=1}^{k}(1+\epsilon)(1-\frac{1}{n_{i}})\geq(1+\epsilon)(\frac{k}{2}-1).$
    Thus, $K_{1}C\leq 0$ implies $k\in\{0,1,2\}$.
    • If $k=0$, then $d=1$ and $KC=-(1+\epsilon)$. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
    • If $k=1$, then $d=2$ and $KC=0$. In this case, $\Theta_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Theta_{1}$.
    • If $k=2$, then $d=3$ and
      $K_{1}C=(1+\epsilon)(1-\frac{1}{n_{1}}-\frac{1}{n_{2}}).$
      Hence, $K_{1}C\leq 0$ implies $(n_{1},n_{2})=(2,2)$. Thus, $C$ meets transversely two (1)(-1)-$\mathcal{F}$-curves $\Gamma_{1}$ and $\Gamma_{2}$ with
      $d(\Gamma_{1})=d(\Gamma_{2})=1,\qquad\Gamma_{1}^{2}=\Gamma_{2}^{2}=-2.$
      By the separatrix theorem, we see $d_{2}(C)={\rm Z}(\mathcal{F},C)>2$. So $d_{2}(C)=d(C)=3$. This is case (3). In this case, $K_{1}C=0$.
Suppose dkl=0d-k-l=0. Then
$K_{1}C$ $=-2(1+\epsilon)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\sum_{i=1}^{k}(1+\epsilon)(1-\frac{1}{n_{i}})+\sum_{j=1}^{l}\epsilon(1-\frac{1}{m_{j}})$
$\geq-2(1+\epsilon)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\tfrac{k}{2}(1+\epsilon).$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}\geq 3$, then
    $K_{1}C\geq-2(1+\epsilon)+3=1-2\epsilon>0.$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=2$ and $k\geq 1$, then
    $K_{1}C\geq-2(1+\epsilon)+2+\frac{1}{2}(1+\epsilon)=\frac{1}{2}(1-3\epsilon)>0.$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=2$ and $k=0$, then $d=l\in\{1,2\}$.
    • If $l=1$, then $d=1$ and ${\rm Z}(\mathcal{F},C)-d_{2}=1$ and so $KC=-\epsilon$. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
    • If $l=2$, then $d=2$ and ${\rm Z}(\mathcal{F},C)-d_{2}=0$ and so $KC=0$. In this case, $\Xi_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Xi_{1}$.
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=1$ and $k\geq 3$, then
    $K_{1}C\geq-2(1+\epsilon)+1+\frac{3}{2}(1+\epsilon)=\frac{1}{2}(1-\epsilon)>0.$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=1$ and $k=2$, then $d=l+2\in\{2,3\}$ and
    $K_{1}C\geq-2(1+\epsilon)+1+\sum_{i=1}^{2}(1+\epsilon)(1-\frac{1}{n_{i}})=1-(1+\epsilon)(\frac{1}{n_{1}}+\frac{1}{n_{2}}).$
    Thus, $K_{1}C\leq 0$ implies
    $\frac{1}{n_{1}}+\frac{1}{n_{2}}\geq\frac{1}{1+\epsilon}>\frac{1}{1+1/3}=\frac{3}{4}.$
    So $(n_{1},n_{2})=(2,2)$ or $(2,3)$, where $(n_{1},n_{2})=(2,3)$ occurs only if $\epsilon\geq 1/5$. By the separatrix theorem, we see ${\rm Z}(\mathcal{F},C)\geq 3$.
    • If $l=0$, then $d=2$ and $d_{2}={\rm Z}(\mathcal{F},C)-1\geq 2$. Thus, $d_{2}=d=2$. This is case (1).
    • If $l=1$, then $d=3$ and $d_{2}={\rm Z}(\mathcal{F},C)\geq 3$. Thus, $d=d_{2}=3$. Moreover, by the separatrix theorem, $\Xi_{1}$ must be a special $K^{>0}$-chain of type (B). Hence, we get the case (2).
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=1$ and $k\in\{0,1\}$, then $d=k+l\in\{1,2\}$.
    • If $d=1$, then $KC=-(1+\epsilon)+{\rm Z}(\mathcal{F},C)-d_{2}\leq-\epsilon$. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
    • If $d=2$, then $k=l=1$ and ${\rm Z}(\mathcal{F},C)-d_{2}=0$. In this case, $KC=0$, and so $\Xi_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Xi_{1}$.
  • If $l={\rm Z}(\mathcal{F},C)-d_{2}=0$, then ${\rm Z}(\mathcal{F},C)=d_{2}$ and
    $k\leq d_{2}\leq d=k,$
    which imply ${\rm Z}(\mathcal{F},C)=d_{2}=k$, contradicting the separatrix theorem.
Proposition 4.13.
Under the assumption ()(*), $E$ belongs to one of the following cases:
  1. (1)
    $E$ is an $\mathcal{F}$-star graph of type $(l;m_{1},\dots,m_{l})$ centered at $C$, where $C$ is a smooth rational curve with ${\rm tang}(\mathcal{F},C)=0$ and $(l;m_{1},\dots,m_{l})$ is one of the following:
    1. (a)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, $(2,3,4)$, or $(2,3,5)$;
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, or $(2,4,4)$, or $(3,3,3)$;
    3. (c)
      $l=4$ and $(m_{1},\dots,m_{4})=(2,2,2,2)$.
    (In cases (b) and (c), $p$ is not $\epsilon$-adjoint klt.)
  2. (2)
    $E=\Theta_{1}+\Theta_{2}+C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, meeting transversely two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of the following:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2;
    2. (b)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
    3. (c)
      $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
  3. (3)
    $E=C+\Theta_{1}+\Theta_{2}+\Xi_{1}$ is a fork centered at $C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, $\Theta_{1}$ and $\Theta_{2}$ are $\mathcal{F}$-chains, and $\Xi_{1}$ is a chain of (2)(-2)-$\mathcal{F}$-curves. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of the following:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2;
    2. (b)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
    3. (c)
      $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
In particular, cases (2-b), (2-c), (3-b) and (3-c) cannot occur if $\epsilon<\tfrac{1}{5}$.
Proof.
By Lemma 4.11 and 4.12, a simple combinatorial analysis yields the desired cases, and the separatrix theorem excludes all others. ∎
4.4.4. Summary of the classification
Theorem 4.14.
Fix $\epsilon\in(0,\tfrac{1}{3})$. Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface, and assume that $p$ is an $\epsilon$-adjoint log canonical singularity of $(Y,\mathcal{G})$. Let
$\pi\colon(X,\mathcal{F})\to(Y,\mathcal{G})$
be the minimal resolution (cf. Definition 2.5) of $(Y,\mathcal{G},p)$, with exceptional divisor $E=\bigcup_{i=1}^{r}E_{i}$. Then $E$ is one of the following:
  1. (1)
    An $\mathcal{F}$-chain. (See Figure 3)
  2. (2)
    A chain of (2)(-2)-$\mathcal{F}$-curves. (See Definition 2.18.)
  3. (3)
    $E=\Theta_{1}+\Theta_{2}+C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, meeting transversely two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2;
    2. (b)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
    3. (c)
      $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
  4. (4)
    $E=C+\Theta_{1}+\Theta_{2}+\Xi_{1}$ is a single fork centered at $C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, $\Theta_{1}$ and $\Theta_{2}$ are $\mathcal{F}$-chains, and $\Xi_{1}$ is a chain of (2)(-2)-$\mathcal{F}$-curves. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2; (see Figure 5)
    2. (b)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
    3. (c)
      $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
  5. (5)
    An elliptic Gorenstein leaf, which is either a rational $\mathcal{F}$-invariant curve with a single node, or a cycle of (2)(-2)-$\mathcal{F}$-curves. (See Figure 6)
  6. (6)
    An $\mathcal{F}$-star chain, that is, a chain with exactly one non-$\mathcal{F}$-invariant curve $E_{k}$, for some $k\in\{1,\dots,r\}$. Moreover, ${\rm tang}(\mathcal{F},E_{k})=0$, and $E_{1}+\cdots+E_{k-1}$ (resp. $E_{k+1}+\cdots+E_{r}$) is an $\mathcal{F}$-chain whose first curve is $E_{k-1}$ (resp. $E_{k+1}$). (See Figure 8.)
  7. (7)
    A smooth elliptic non-$\mathcal{F}$-invariant curve $E_{1}$ with ${\rm tang}(\mathcal{F},E_{1})=0$.
  8. (8)
    An $\mathcal{F}$-star graph of type $(l;m_{1},\dots,m_{l})$ centered at $C$ (cf. Definition 2.22), where $C$ is a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$, and $(l;m_{1},\dots,m_{l})$ is one of:
    1. (a)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, $(2,3,4)$, or $(2,3,5)$;
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, or $(2,4,4)$, or $(3,3,3)$;
    3. (c)
      $l=4$ and $(m_{1},\dots,m_{4})=(2,2,2,2)$.
    (See Definition 2.22 and Figure 7.)
Moreover, the following hold:
  • (i)
    Type (5), (7), (8-b), and (8-c) are not $\epsilon$-adjoint klt.
  • (ii)
    Type (3-b), (3-c), (4-b) and (4-c) cannot occur if $\epsilon\in(0,1/5)$.
Proof.
The statement follows by combining Proposition 4.8, Proposition 4.9, Proposition 4.10, and Proposition 4.13. ∎
5. $\epsilon$-adjoint canonical singularities with $\epsilon\in(0,1/4)$
Throughout this section, we fix $\epsilon\in(0,1/4)$. We characterize those configurations in Theorem 4.14 that are $\epsilon$-adjoint canonical.
Lemma 5.1.
Type (5), (7), (8-b), and (8-c) are not $\epsilon$-adjoint canonical.
Proof.
This follows directly from Theorem 4.14. ∎
Lemma 5.2.
Type (3-b), (3-c), (4-b), and (4-c) are not $\epsilon$-adjoint canonical.
Proof.
Let $K^{\prime}:=K_{\mathcal{F}}+\epsilon K_{X}$. For each $\Theta_{i}$, let $M(K^{\prime},\Theta_{i})$ denote the effective divisor supported on $\Theta_{i}$ such that
$M(K^{\prime},\Theta_{i})\cdot\Gamma=K^{\prime}\cdot\Gamma$
for any irreducible component $\Gamma$ of $\Theta_{i}$. Set
$W:=K^{\prime}-M(K^{\prime},\Theta_{1})-M(K^{\prime},\Theta_{2}).$
Suppose that $p$ is an $\epsilon$-adjoint canonical singularity. Then $E$ is a $(K^{\prime})^{\geq 0}$-graph and hence also a $W^{\geq 0}$-graph. Clearly, $W\cdot\Gamma\geq 0$ for any component $\Gamma\subset E$ with $\Gamma\neq C$. We claim that $W\cdot C>0$, which yields a contradiction.
We have
$M(K^{\prime},\Theta_{1})=\frac{1}{2}\Gamma_{11}.$
  • If $\Theta_{2}$ is of type (3-b) or (4-b), then
    $M(K^{\prime},\Theta_{2})=\frac{1-\epsilon}{3}\Gamma_{21}.$
    Note that
    $K^{\prime}\cdot C=K_{\mathcal{F}}\cdot C+\epsilon K_{X}\cdot C\geq 1-\epsilon.$
    Hence,
    $W\cdot C$ $=K^{\prime}\cdot C-M(K^{\prime},\Theta_{1})\cdot C-M(K^{\prime},\Theta_{2})\cdot C$
    $\geq 1-\epsilon-\frac{1}{2}-\frac{1-\epsilon}{3}=\frac{1-4\epsilon}{6}>0.$
  • If $\Theta_{2}$ is of type (3-c) or (4-c), then
    $M(K^{\prime},\Theta_{2})=\frac{2}{3}\Gamma_{21}+\frac{1}{3}\Gamma_{22}.$
    In this case, we must have $C^{2}\leq-2$; otherwise, the intersection matrix of $\Gamma_{11}+C+\Gamma_{21}$ is not negative definite. Hence $K^{\prime}\cdot C\geq K_{\mathcal{F}}C=-C^{2}\geq 2$, and therefore
    $W\cdot C\geq 2-\frac{1}{2}-\frac{1}{3}=\frac{7}{6}>0.$
Lemma 5.3.
Type (8-a) is not $\epsilon$-adjoint canonical.
Proof.
We follow the notation of Lemma 5.2. It suffices to show $W\cdot C>0$, where $C$ is as in type (8) of Theorem 4.14 and
$W:=K^{\prime}-\sum_{i=1}^{3}M(K^{\prime},\Theta_{i}).$
In case (8-a), we have $e_{C}:=-C^{2}\geq 2$, since the intersection matrix of $C+\Theta_{1}+\Theta_{2}+\Theta_{3}$ is negative definite. Thus,
$K^{\prime}\cdot C=K_{\mathcal{F}}C+\epsilon K_{X}C=(1+\epsilon)e_{C}-2\epsilon\geq 2.$
Case $(2,2,n)$, $n\geq 2$. We have
$M(K^{\prime},\Theta_{1})=\frac{1}{2}\Gamma_{11},\qquad M(K^{\prime},\Theta_{2})=\frac{1}{2}\Gamma_{21}.$
Thus,
$M(K^{\prime},\Theta_{1})\cdot C=M(K^{\prime},\Theta_{2})\cdot C=\frac{1}{2}.$
On the other hand, since the coefficients of $M(K^{\prime},\Theta_{3})$ are at most (n1)/n(n-1)/n, we have
$M(K^{\prime},\Theta_{3})\cdot C\leq\frac{n-1}{n}.$
Hence,
$W\cdot C\geq 2-\frac{1}{2}-\frac{1}{2}-\frac{n-1}{n}=\frac{1}{n}>0.$
Case $(2,3,n)$, $n=3,4,5$. Here $\Theta_{1}=\Gamma_{11}$ with $\Gamma_{11}^{2}=-2$, and $\Theta_{2}$ is either
  • (i)
    $\Gamma_{21}$ with $\Gamma_{21}^{2}=-3$, or
  • (ii)
    $\Gamma_{21}+\Gamma_{22}$ with $\Gamma_{21}^{2}=\Gamma_{22}^{2}=-2$.
In case (i),
$M(K^{\prime},\Theta_{1})C=\frac{1}{2},\quad M(K^{\prime},\Theta_{2})C=\frac{1-\epsilon}{3},\quad M(K^{\prime},\Theta_{3})C\leq\frac{n-1}{n}\leq\frac{4}{5},$
which imply
$W\cdot C\geq 2-\frac{1}{2}-\frac{1-\epsilon}{3}-\frac{4}{5}=\frac{11}{30}+\frac{\epsilon}{3}>0.$
In case (ii),
$M(K^{\prime},\Theta_{1})C=\frac{1}{2},\quad M(K^{\prime},\Theta_{2})C=\frac{2}{3},\quad M(K^{\prime},\Theta_{3})C\leq\frac{n-1}{n}\leq\frac{4}{5},$
which imply
$W\cdot C\geq 2-\frac{1}{2}-\frac{2}{3}-\frac{4}{5}=\frac{1}{30}>0.$
Lemma 5.4.
If type (3-a) (resp. type (4-a)) is $\epsilon$-adjoint canonical, then every component of $E$ has self-intersection 2-2.
Proof.
The proof is similar to that of Lemma 5.2. ∎
Lemma 5.5.
If type (6) is $\epsilon$-adjoint canonical, then the non-$\mathcal{F}$-invariant curve $E_{k}$ is a (1)(-1)-curve.
Proof.
We use the notation as in Lemma 5.2. Let $C:=E_{k}$. In this case, $C$ meets $\mathcal{F}$-chains $\Theta_{1},\dots,\Theta_{\ell}$ with $\ell\leq 2$. Set
$W:=K^{\prime}-\sum_{i=1}^{\ell}M(K^{\prime},\Theta_{i}).$
Assume $C^{2}\leq-2$. Then
$K^{\prime}\cdot C=K_{\mathcal{F}}\cdot C+\epsilon K_{X}\cdot C\geq 2.$
Since $M(K^{\prime},\Theta_{i})\cdot C<1$ for all $i$, we obtain
$W\cdot C>2-\ell\geq 0,$
contradicting the fact that $E$ is a $W^{\geq 0}$-graph. ∎
The above lemmas exclude most configurations in Theorem 4.14 under the $\epsilon$-adjoint canonical condition for $\epsilon\in(0,1/4)$. The remaining cases yield the following classification.
Theorem 5.6.
Fix $\epsilon\in(0,\tfrac{1}{4})$. Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface, and assume that $p$ is an $\epsilon$-adjoint canonical singularity of $(Y,\mathcal{G})$. Let
$\pi\colon(X,\mathcal{F})\to(Y,\mathcal{G})$
be the minimal resolution, with exceptional divisor $E=\bigcup_{i=1}^{r}E_{i}$. Then $E$ is one of the following:
  1. (1)
    An $\mathcal{F}$-chain. (See Figure 3.)
  2. (2)
    A chain of (2)(-2)-$\mathcal{F}$-curves. (See Definition 2.18.)
  3. (3)
    A chain $E_{1}+E_{2}+E_{3}$, where $E_{1}$ and $E_{3}$ are (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2 and $E_{2}$ is a bad tail. Moreover, in this case, $E_{2}^{2}=-2$. (See Figure 4.)
  4. (4)
    An $\mathcal{F}$-dihedral graph, all of whose components have self-intersection 2-2. (See Figure 5.)
  5. (5)
    An $\mathcal{F}$-star chain, that is, a chain with exactly one smooth rational non-$\mathcal{F}$-invariant curve $E_{k}$ with ${\rm tang}(\mathcal{F},E_{k})=0$, such that the two sides are $\mathcal{F}$-chains. Moreover, in this case, $E_{k}^{2}=-1$. (See Figure 8.)
Here, types (1)(4) are canonical foliated singularities, and types (3) and (4) are not $\epsilon$-adjoint terminal.
Appendix A Log canonical surface singularities
In this appendix, we give a proof of the classification of log canonical surface singularities, following an approach analogous to the program developed in Section 4.
Let $Y$ be a normal surface and let $p$ be a log canonical singularity of $Y$. Let
$\pi:X\to Y$
be a minimal resolution of $(Y,p)$ with exceptional curves $\mathfrak{E}=\bigcup_{i\in I}E_{i}$, and set
$E:=\sum_{i\in I}E_{i}.$
Then we can write
$K:=K_{X}+\sum_{i\in I}E_{i}=\pi^{*}K_{Y}+\sum_{i\in I}a_{i}E_{i},$
so that $E$ is a $K^{\geq 0}$-graph.
A.1. $K$-non-positive curves in $E$
Lemma A.1.
Let $C\subset E$ be an irreducible $\mathcal{F}$-invariant curve. Then
(A.1) $KC=2p_{a}(C)-2+d(C).$
Moreover, $KC\leq 0$ if and only if $C$ is one of:
  1. (1)
    $d(C)=0$:
    1. (a)
      a smooth rational curve. In this case, KC=2KC=-2.
    2. (b)
      a smooth elliptic curve. In this case, $KC=0$.
    3. (c)
      a rational curve with a single node. In this case, $KC=0$.
  2. (2)
    $C$ is a smooth rational curve with $d(C)=1$. In this case, KC=1KC=-1.
  3. (3)
    $C$ is a smooth rational curve with $d(C)=2$. In this case, $KC=0$.
A.2. Special $K^{>0}$-chain
Proposition A.2.
Suppose $E$ has at least two components. Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{>0}$-chain with the first curve $C_{1}$. Then $F$ is a Hirzebruch-Jung chain with
$d(C_{1})=1,\qquad d(C_{i})=2,\quad i\geq 2,$
and
$KC_{1}=1,\qquad KC_{2}=0,\quad i\geq 2.$
Proof.
The proof follows similarly to Proposition 4.5 and is in fact simpler in this setting. ∎
A.3. Classification
Let $G_{E}$ denote the dual graph of $E$, which is connected.
A.3.1. Case I: $G_{E}$ consists of a single curve
Proposition A.3.
If $G_{E}=\{[C]\}$, then $C$ is one of:
  1. (1)
    a smooth rational curve. In this case, KC=2KC=-2.
  2. (2)
    a smooth elliptic curve. In this case, $KC=0$.
  3. (3)
    a rational curve with a single node. In this case, $KC=0$.
Proof.
This follows from Lemma A.1. ∎
A.3.2. Case II: $E$ is a $K^{=0}$-graph
Proposition A.4.
If $KC=0$ for every component $C\subset E$, then $E$ is one of:
  1. (1)
    an irreducible curve: a smooth elliptic curve or a rational curve with a single node,
  2. (2)
    a cycle of smooth rational curve.
Proof.
This follows from Lemma A.1. ∎
A.3.3. The presence of special $K^{>0}$-chain
We now assume that $E$ does not belong to Case I or Case II. Then $E$ necessarily contains at least one special $K^{>0}$-chain.
Proposition A.5.
Suppose that there exists two special $K^{>0}$-chains $F_{1}$ and $F_{2}$ with distinct first curves, such that they share a common component or satisfy $F_{1}F_{2}>0$. Then $E$ is a Hirzebruch-Jung chain.
Proof.
This follows from Proposition A.2. ∎
To separate the remaining cases, we impose the following technical assumption.
    $E$ does not belong to Case I or Case II, and any two special $K^{>0}$-chains with distinct first curves are pairwise disjoint.
Under the assumption ()(*), we can define the maximal special $K^{>0}$-chain. Let $\{\,\Theta_{1},\cdots,\Theta_{s}\,\}$ denotes the set of all maximal special $K^{>0}$-chains contained in $E$. Define
$K_{1}:=K-\sum_{i=1}^{s}M(K,\Theta_{i}).$
It is clear that $E$ is a $(K_{1})^{\geq 0}$-graph and for any component $C$ of $\Theta_{i}$, one has $K_{1}C=0$.
Lemma A.6.
Under the assumption ()(*), let $C$ be an irreducible curve contained in $E$, and assume $C$ is not contained in any special $K^{>0}$-chain. If $K_{1}C\leq 0$, then one of the following holds:
  1. (1)
    $C$ is smooth rational curve with $d(C)=3$, meeting two special $K^{>0}$-chains $\Theta_{1},\Theta_{2}$ of type $(2,2)$. In this case, $K_{1}C=0$.
  2. (2)
    $C$ is a smooth rational curve with $d(C)=2$, disjoint from all special $K^{>0}$-chain. In this case, $K_{1}C=KC=0$.
  3. (3)
    $E=C+\Theta_{1}+\cdots+\Theta_{l}$ is a single fork centered at $C$, where $C$ is a smooth rational curve and $(\Theta_{1},\cdots,\Theta_{l})$ is of type $(m_{1},\cdots,m_{l})$, which belongs to one of
    1. (a)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, or $(2,3,4)$, or $(2,3,5)$;
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, $(2,4,4)$, $(3,3,3)$.
    3. (c)
      $l=4$ and $(m_{1},m_{2},m_{3},m_{4})=(2,2,2,2)$.
    In the case (b) and (c), $K_{1}C=0$.
Proof.
Without loss of generality, we assume that $C$ meets transversely the chains $\Theta_{1},\cdots,\Theta_{k}$. Let $m_{j}$ denots the absolute value of the determinant of the intersection matrix of $\Theta_{j}$. Note that $m_{j}\geq 2$. Thus, we have
$K_{1}C$ $=KC-\sum_{i=1}^{k}M(K,\Theta_{i})C$
$=2p_{a}(C)-2+d(C)-\sum_{i=1}^{k}\frac{1}{m_{i}}$
$=2p_{a}(C)-2+(d-k)+\sum_{i=1}^{k}(1-\frac{1}{m_{i}}),$
where $d-k\geq 0$.
If $p_{a}(C)\geq 1$, then $K_{1}C\leq 0$ implies $p_{a}(C)=1$ and $d=k=0$, contradicting our assumption. Next we assume $p_{a}(C)=0$. Then
$K_{1}C=d-k-2+\sum_{i=1}^{k}(1-\frac{1}{m_{i}}).$
  • If $d-k\geq 2$, then $K_{1}C\leq 0$ implies $k=0$ and $d=2$. In this case, $K_{1}C=KC=0$ and $C$ is a smooth rational curve with $d(C)=2$, disjoint from all special $K^{>0}$-chain. This is case (2).
  • If dk=1d-k=1, then
    $K_{1}C=-1+\sum_{i=1}^{k}(1-\frac{1}{m_{i}})\geq-1+\frac{k}{2}$
    • If $k\geq 3$, then $K_{1}C>0$.
    • If $k=2$, then $K_{1}C\leq 0$ implies $d=3$, $k=2$ and $(m_{1},m_{2})=(2,2)$. This is case (1). In this case, $K_{1}C=0$.
    • If $k=1$, then $d=2$ and $KC=0$. In this case, $\Theta_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Theta_{1}$.
    • If $k=0$, then $d=1$ and KC=1KC=-1. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
  • If dk=0d-k=0, then
    $K_{1}C=-2+\sum_{i=1}^{k}(1-\frac{1}{m_{i}})\geq-2+\frac{k}{2}.$
    $K_{1}C\leq 0$ implies $k\leq 4$.
    • If $k=0$, then $d=0$, contradicting our assumption.
    • If $k=1$, then $d=1$ and KC=1KC=-1. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
    • If $k=2$, then $d=2$ and $KC=0$. In this case, $\Theta_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Theta_{1}$.
    • If $k=3$, then $K_{1}C\leq 0$ implies
      $\frac{1}{m_{1}}+\frac{1}{m_{2}}+\frac{1}{m_{3}}\geq 1.$
      Then $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,l)$ for $l=3,4,5,6$, or $(2,4,4)$, or $(3,3,3)$. These are case (3-a) and (3-b).
      In particular, $K_{1}C=0$ for $(m_{1},m_{2},m_{3})=(2,3,6)$, or $(2,4,4)$, or $(3,3,3)$.
    • If $k=4$, then $K_{1}C\leq 0$ implies $(m_{1},\cdots,m_{4})=(2,2,2,2)$. In this case, $K_{1}C=0$. This is case (3-c).
Proposition A.7.
Under the assumption ()(*), $E$ belongs to one of the following cases:
  1. (1)
    $E$ has two forks with the dual graph as follows:
[…]
    where each component is a smooth rational curve with self-intersection at most 2-2. In this case, $p$ is not klt.
  2. (2)
    $E=C+\Theta_{1}+\cdots+\Theta_{l}$ is a single fork centered at $C$, where $C$ is a smooth rational curve and $(\Theta_{1},\cdots,\Theta_{l})$ is of type $(m_{1},\cdots,m_{l})$, which belongs to one of
    1. (a)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, or $(2,3,4)$, or $(2,3,5)$;
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, $(2,4,4)$, $(3,3,3)$.
    3. (c)
      $l=4$ and $(m_{1},m_{2},m_{3},m_{4})=(2,2,2,2)$.
    In the case (b) and (c), $K_{1}C=0$.
Proof.
This follows from Lemma A.6. ∎
A.3.4. Summary of the classification
Theorem A.8.
Let $p$ be a log canonical singularity of a normal surface $Y$. Let
$\pi:X\to Y$
be the minimal resolution of $(Y,p)$, with the exceptional divisor $E=\cup_{i}E_{i}$. Then $E$ is one of the following:
  1. (1)
    $E$ is a Hirzebruch-Jung chain.
  2. (2)
    $E=C$ is an irreducible curve, where $C$ is a smooth elliptic curve or a rational curve with a single node.
  3. (3)
    $E$ is a cycle of smooth rational curves.
  4. (4)
    $E$ has two forks with the dual graph as follows:
[…]
    where each component is a smooth rational curve with self-intersection at most 2-2.
  5. (5)
    $E$ is a single fork centered at a smooth rational curve $C$, and $E\setminus C$ are pairwise disjoint Hirzebruch-Jung chains $\Theta_{1},\cdots,\Theta_{l}$ of type $(m_{1},\cdots,m_{l})$, which belongs to one of
    1. (a)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, or $(2,3,4)$, or $(2,3,5)$;
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, $(2,4,4)$, $(3,3,3)$.
    3. (c)
      $l=4$ and $(m_{1},m_{2},m_{3},m_{4})=(2,2,2,2)$.
Here type (2), (3), (4), (5-b) and (5-c) are not klt.
Proof.
The statement follows by combining Proposition A.3, Proposition A.4, Proposition A.5, and Proposition A.7. ∎
Appendix B Log canonical foliated singularities
In this appendix, we give a proof of the classification of log canonical foliated singularities, following an approach analogous to the program developed in Section 4, differing from that of Y. Chen.
Let $(Y,\mathcal{G})$ be a foliated pair on a normal surface $Y$, and let $p\in Y$ be a log canonical singularity of $\mathcal{G}$. Let
$\pi:(X,\mathcal{F})\longrightarrow(Y,\mathcal{G})$
be a minimal resolution of $(Y,\mathcal{G},p)$, with exceptional curves $\mathfrak{E}=\bigcup_{i\in I}E_{i}$, and set
$E:=\sum_{i\in I}E_{i}.$
Then we can write
$K:=K_{\mathcal{F}}+\sum_{i\in I}\iota(E_{i})E_{i}=\pi^{*}K_{\mathcal{G}}+\sum_{i\in I}a_{i}E_{i},$
so that $E$ is a $K^{\geq 0}$-graph.
B.1. $K$-non-positive curves in $E$
Lemma B.1.
Let $C\subset E$ be an irreducible $\mathcal{F}$-invariant curve. Then
(B.1) $KC=2p_{a}(C)-2+d(C)+{\rm Z}(\mathcal{F},C)-d_{2}(C).$
Moreover, $KC\leq 0$ if and only if $C$ is one of:
  1. (1)
    $d(C)=0$:
    1. (a)
      an irreducible elliptic Gorenstein leaf. In this case, $KC=0$,
    2. (b)
      a (2)(-2)-$\mathcal{F}$-curve. In this case, $KC=0$,
    3. (c)
      a (1)(-1)-$\mathcal{F}$-curve. In this case, KC=1KC=-1.
  2. (2)
    $d(C)=1$:
    1. (a)
      a (1)(-1)-$\mathcal{F}$-curve, $d_{1}=1$, $d_{2}=0$. In this case, $KC=0$.
    2. (b)
      a (1)(-1)-$\mathcal{F}$-curve, $d_{1}=0$, $d_{2}=1$. In this case, KC=1KC=-1,
    3. (c)
      a (2)(-2)-$\mathcal{F}$-curve, $d_{1}=0$, $d_{2}=1$. In this case, $KC=0$.
  3. (3)
    $d(C)=2$:
    1. (a)
      a (1)(-1)-$\mathcal{F}$-curve, $d_{1}=d_{2}=1$. In this case, $KC=0$.
    2. (b)
      a (2)(-2)-$\mathcal{F}$-curve, $d_{1}=0$, $d_{2}=2$. In this case, $KC=0$.
Proof.
This follows from a direct computation. ∎
Lemma B.2.
Let $C\subset E$ be a non-$\mathcal{F}$-invariant curve. Then
(B.2) $KC={\rm tang}(\mathcal{F},C)+d_{1}(C)\geq 0,$
and $KC=0$ if and only if $C$ is smooth, with ${\rm tang}(\mathcal{F},C)=d_{1}(C)=0$.
Proof.
This follows from a direct computation. ∎
B.2. Special $K^{>0}$-chains
Proposition B.3.
Suppose $E$ has at least two components. Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{>0}$-chain with first curve $C_{1}$. Then $F$ is an $\mathcal{F}$-chain with
$d(C_{1})=d_{2}(C_{1})=1,\quad d(C_{i})=d_{2}(C_{i})=2\,(2\leq i\leq r-1),\quad d(C_{r})=d_{2}(C_{r})\in\{1,2\},$
and
$KC_{1}=-1,\quad KC_{i}=0\text{ for }i\geq 2.$
Proof.
The proof follows similarly to Proposition 4.5 and is in fact simpler in this setting. ∎
Lemma B.4.
Any two special $K^{>0}$-chains with distinct first curves are disjoint.
Proof.
This directly follows from the separatrix theorem. ∎
B.3. Classification
Let $G_{E}$ denote the dual graph of $E$, which is connected.
B.3.1. Case I: $G_{E}$ consists of a single curve
Proposition B.5.
If $G_{E}=\{[C]\}$, then $C$ is one of:
  1. (1)
    an irreducible elliptic Gorenstein leaf. In this case, $KC=0$.
  2. (2)
    a smooth non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$. In this case, $KC=0$.
  3. (3)
    a (2)(-2)-$\mathcal{F}$-curve. In this case, $KC=0$.
  4. (4)
    a (1)(-1)-$\mathcal{F}$-curve. In this case, KC=1KC=-1.
Proof.
This follows from Lemma B.1 and Lemma B.2. ∎
B.3.2. Case II: $E$ is a $K^{=0}$-graph
Proposition B.6.
If $KC=0$ for every component $C\subset E$, then $E$ is one of:
  1. (1)
    An irreducible curve: irreducible elliptic Gorenstein leaf or smooth non-$\mathcal{F}$-invariant elliptic curve with ${\rm tang}(\mathcal{F},C)=0$.
  2. (2)
    A chain of (2)(-2)-curves.
  3. (3)
    A cycle of (2)(-2)-$\mathcal{F}$-curves.
  4. (4)
    An $\mathcal{F}$-star graph centered at a smooth non-$\mathcal{F}$-invariant curve $C$ with ${\rm tang}(\mathcal{F},C)=0$.
Proof.
This follows from Lemma B.1, Lemma B.2 and the separatrix theorem. ∎
B.3.3. Remaining cases
Let $\{\Theta_{1},\dots,\Theta_{s}\}$ be the set of all maximal special $K^{>0}$-chains in $E$, and set
$K_{1}:=K-\sum_{i=1}^{s}M(K,\Theta_{i}).$
Lemma B.7.
Let $C\subset E$ be irreducible and not in any special $K^{>0}$-chain. If $K_{1}C\leq 0$, then $C$ is one of the following:
  1. (1)
    $E=C_{1}+C+C_{2}$ is a chain where $C_{1},C_{2}$ are (1)(-1)-$\mathcal{F}$-curves, and $C_{2}$ is a bad tail. In this case, $K_{1}C=0$.
  2. (2)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve meeting two (1)(-1)-$\mathcal{F}$-curves $C_{1},C_{2}$, with $d(C_{1})=d(C_{2})=1$, $C_{1}^{2}=C_{2}^{2}=-2$ and $d(C)=d_{2}(C)=3$. In this case, $K_{1}C=0$.
  3. (3)
    $C$ is smooth non-$\mathcal{F}$-invariant, ${\rm tang}(\mathcal{F},C)=d_{1}(C)=0$, disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
  4. (4)
    $C$ is a (1)(-1)-$\mathcal{F}$-curve, $d_{1}=d_{2}=1$, disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
  5. (5)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve, $d_{1}=0$, $d_{2}=2$, disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
  6. (6)
    $C$ is a (1)(-1)-$\mathcal{F}$-curve, $d=d_{1}=1$, disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
  7. (7)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve, $d=d_{2}=1$, disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
Proof.
If $C$ is not $\mathcal{F}$-invariant, then it is disjoint from all special $K^{>0}$-chains. Hence $K_{1}\cdot C=K\cdot C\geq 0$, with equality if and only if $C$ is smooth, ${\rm tang}(\mathcal{F},C)=d_{1}(C)=0$, and $C$ is disjoint from all special $K^{>0}$-chains. This corresponds to case (3).
We now assume that $C$ is $\mathcal{F}$-invariant. Without loss of generality, suppose that $C$ meets $\Theta_{1},\dots,\Theta_{k}$ transversely at distinct points $p_{1},\dots,p_{k}$. Let $n_{i}$ denote the absolute value of the determinant of the intersection matrix of $\Theta_{i}$. Note that $n_{i}\geq 2$ for all $i$. Then
$K_{1}\cdot C$ $=K\cdot C-\sum_{i=1}^{k}M(K,\Theta_{i})\cdot C$
$=2p_{a}(C)-2+d(C)+{\rm Z}(\mathcal{F},C)-d_{2}(C)-\sum_{i=1}^{k}\frac{1}{n_{i}}$
$=2p_{a}(C)-2+(d-k)+({\rm Z}(\mathcal{F},C)-d_{2})+\sum_{i=1}^{k}\Bigl(1-\frac{1}{n_{i}}\Bigr),$
where $d\geq k$ and ${\rm Z}(\mathcal{F},C)\geq d_{2}$.
If $p_{a}(C)\geq 1$, then $K_{1}\cdot C\leq 0$ forces $p_{a}(C)=1$ and $d=k=0$, contradicting our assumption.
We therefore assume $p_{a}(C)=0$. In this case,
$K_{1}\cdot C$ $=-2+(d-k)+({\rm Z}(\mathcal{F},C)-d_{2})+\sum_{i=1}^{k}\Bigl(1-\frac{1}{n_{i}}\Bigr)$
$\geq-2+(d-k)+({\rm Z}(\mathcal{F},C)-d_{2})+\frac{k}{2}$
By the separatrix theorem, we have ${\rm Z}(\mathcal{F},C)\geq k+1$.
  • If $d-k\geq 3$, then $K_{1}\cdot C\geq 1>0$.
  • If dk=2d-k=2, then $K_{1}\cdot C\leq 0$ implies $k=0$, $d=2$ and ${\rm Z}(\mathcal{F},C)=d_{2}\in\{1,2\}$. Hence $C$ is either a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$, or a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=2$. In both cases, $C$ is disjoint from all special $K^{>0}$-chains and $K_{1}C=KC=0$. This yields case (4) or (5).
  • Suppose dk=1d-k=1.
    • If ${\rm Z}(\mathcal{F},C)-d_{2}\geq 2$, then $K_{1}\cdot C>0$.
    • If ${\rm Z}(\mathcal{F},C)-d_{2}=1$, then $K_{1}\cdot C\leq 0$ implies $k=0$, $d=1$ and ${\rm Z}(\mathcal{F},C)=d_{2}+1\in\{1,2\}$. Hence $C$ is either a (1)(-1)-$\mathcal{F}$-curve with $d=d_{1}=1$, or a (2)(-2)-$\mathcal{F}$-curve with $d=d_{2}=1$. Both cases satisfy $K_{1}C=KC=0$ and are disjoint from all special $K^{>0}$-chains. This corresponds to case (6) or (7).
    • If ${\rm Z}(\mathcal{F},C)-d_{2}=0$, then
      $k+1\leq{\rm Z}(\mathcal{F},C)=d_{2}\leq d=k+1,$
      so $d=d_{2}={\rm Z}(\mathcal{F},C)=k+1$.
        If $k\geq 3$, then $K_{1}C>0$.
        If $k=2$, then $K_{1}C\leq 0$ forces $(n_{1},n_{2})=(2,2)$ and $K_{1}\cdot C=0$, giving case (2).
        If $k=1$, then $K_{1}C\leq 0$ implies that $C$ is a (2)(-2)-$\mathcal{F}$-curve with $d=d_{2}=2$. In this case $KC=0$, so $\Theta_{1}+C$ forms a special $K^{>0}$-chain, contradicting the maximality of $\Theta_{1}$.
        If $k=0$, then $K_{1}C=-1$, so $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
  • If dk=0d-k=0, then $1\leq d_{2}=d=k<{\rm Z}(\mathcal{F},C)$.
    • If ${\rm Z}(\mathcal{F},C)-d_{2}\geq 2$, then $K_{1}C>0$.
    • If ${\rm Z}(\mathcal{F},C)-d_{2}=1$, then
      $K_{1}C=-1+\sum_{i=1}^{k}\Bigl(1-\frac{1}{n_{i}}\Bigr)\geq-1+\frac{k}{2}.$
      Hence $K_{1}C\leq 0$ implies $k\in\{1,2\}$. If $k=1$, then $KC=0$ and $\Theta_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Theta_{1}$. If $k=2$, this gives case (1).
Proposition B.8.
If $E$ does not belong to Case I or II, then $E$ is one of:
  1. (1)
    An $\mathcal{F}$-chain $C_{1}+\cdots+C_{r}$.
  2. (2)
    A chain $C_{1}+C+C_{2}$ where $C_{1},C_{2}$ are (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2 and $C_{2}$ is a bad tail. In this case, $p$ is not log terminal.
  3. (3)
    An $\mathcal{F}$-dihedral graph. (See Figure 5.)
Proof.
This follows directly from Lemma B.7 together with the separatrix theorem. ∎
Conclusion
We now summarize the classification.
Theorem B.9 (Y. Chen).
Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface with log canonical singularity at $p$, and let $\pi:(X,\mathcal{F})\to(Y,\mathcal{G})$ be the minimal resolution with exceptional divisors $E=\cup_{i}E_{i}$. Then $E$ is one of the following:
  1. (1)
    An $\mathcal{F}$-chain. (See Figure 3.)
  2. (2)
    A chain of (2)(-2)-$\mathcal{F}$-curves. (See Definition 2.18.)
  3. (3)
    A chain of three $\mathcal{F}$-invariant curves $E_{1}+E_{2}+E_{3}$, where $E_{1},E_{3}$ are (1)(-1)-$\mathcal{F}$-curves and $E_{2}$ is a bad tail. (See Figure 4.)
  4. (4)
    a $\mathcal{F}$-dihedral graph: two (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2 joined by a bad tail connecting to a chain of (2)(-2)-$\mathcal{F}$-curves. (See Figure 5.)
  5. (5)
    An elliptic Gorenstein leaf: either a rational $\mathcal{F}$-invariant curve with a single node or a cycle of (2)(-2)-$\mathcal{F}$-curves. (See Figure 6.)
  6. (6)
    An $\mathcal{F}$-star chain: a chain with exactly one non-$\mathcal{F}$-invariant curve $E_{k}$, ${\rm tang}(\mathcal{F},E_{k})=0$, with the remaining curves forming $\mathcal{F}$-chains. (See Figure 8.)
  7. (7)
    An $\mathcal{F}$-star graph centered at a smooth non-$\mathcal{F}$-invariant curve $C$ with ${\rm tang}(\mathcal{F},C)=0$. (See Definition 2.22 and Figure 7.)
Type (1) is terminal; types (1)–(5) are canonical.
Proof.
Combine Propositions B.5, B.6, and B.8. ∎
Acknowledgements
The author sincerely thanks Professors Shengli Tan, Jun Lu, and Xin Lü for their continuous encouragement and support, and Professor Xiaohang Wu for many helpful discussions. He is also grateful to Professor Jihao Liu for posing question Q2, which inspired the initial ideas of this work. He further thanks the referees for carefully reading the manuscript, pointing out errors, and providing detailed comments that greatly improved the paper.