### v2

# Sharp Thresholds for $\epsilon$-Adjoint Singularities of Foliated Surfaces

Abstract.
Let $(X,\mathcal{F})$ be a foliated surface over the complex numbers. We study the variation of $\epsilon$-adjoint singularities, defined by the adjoint divisor $K_{\mathcal{F}}+\epsilon K_{X}$ ($\epsilon>0$), and analyze their stability as $\epsilon$ varies.
We prove that a sharp first stability threshold occurs at $\epsilon=1/5$: for $\epsilon\in(0,1/5)$, every $\epsilon$-adjoint log canonical singularity is foliated log canonical, while at $\epsilon=1/5$ a boundary configuration enters the admissible region. In the adjoint canonical setting, the maximal stability interval is $\epsilon\in(0,1/4)$. Both thresholds are optimal and arise from explicit extremal configurations.
These results are obtained via a complete classification of $\epsilon$-adjoint log canonical singularities for $\epsilon\in(0,1/3)$ in terms of negative definite exceptional configurations.
1. Introduction
Throughout this paper, we work over the complex numbers $\mathbb{C}$.
Singularities play a central role in birational geometry, arising naturally in the minimal model program and in the construction of moduli spaces. Among them, (log) canonical singularities form one of the most fundamental classes. In dimension two, log canonical surface singularities admit a complete classification (see [Alex92], [KM98, Sec. 4.1], or Theorem 2.24).
For foliations, McQuillan [McQ08] introduced (log) canonical singularities of foliations and classified canonical singularities of foliations on surfaces. The log canonical case was later established by Y. Chen [Chen23, Theorem 1.1] (see also Theorem 2.25). These results provide a detailed structural description of log canonical singularities of foliations on surfaces.
Adjoint foliated structures, introduced by Pereira and Svaldi [PS19], interpolate between the intrinsic birational geometry of a foliation and the ambient geometry of the surface. This framework was further developed by Spicer and Svaldi for rank-one foliations on surfaces [SS23], and by Cascini–Han–Liu–Meng–Spicer–Svaldi–Xie for algebraically integrable foliations in higher dimensions [CHL+24]. Adjoint divisors
$K_{\mathcal{F}}+\epsilon K_{X},\qquad 0<\epsilon\ll 1,$
provide a natural deformation of the foliated canonical class and address structural pathologies in the birational theory of foliations. For instance, the canonical ring of a foliation need not be finitely generated even if $K_{\mathcal{F}}$ is big and nef (cf. [McQ08, Theorem IV.2.2]), and effective birationality may fail even for rank-one foliations [Lu25].
A natural question is whether sufficiently small adjoint perturbations preserve canonical singularities. However, this is not always the case: there exist adjoint canonical singularities for which arbitrarily small adjoint deformations may destroy canonicity (see Example 6.1). Thus the canonical condition is not stable under small variations of $\epsilon$.
Since canonical singularities form a subclass of log canonical ones, we first study adjoint log canonical singularities, with adjoint canonical singularities imposing additional constraints.
Passing to a minimal resolution, the $\epsilon$-adjoint log canonical singularities correspond to rational polyhedral regions defined by a finite system of linear inequalities, whose coefficients depend on $\epsilon$ and on the configuration of exceptional curves. In this way, the variation of $\epsilon$ can be interpreted as a linear deformation of discrepancy inequalities. As $\epsilon$ increases, the polyhedral region shrinks linearly, and critical values
$0<\epsilon_{1}<\epsilon_{2}<\cdots$
occur when one inequality becomes sharp, giving rise to combinatorial wall-crossing phenomena at which new configurations may enter the admissible region. Our primary focus is on the first critical value $\epsilon_{1}$, which governs the initial stability behavior of adjoint singularities.
In dimension two, the geometry of negative definite configurations allows these extremal situations to be tracked explicitly. To study $\epsilon_{1}$, it is convenient to work with a slightly larger interval
$\epsilon\in(0,1/3),$
where the upper bound $1/3$ is dictated by the structure of the discrepancy inequalities, marking the next potential threshold beyond which new configurations may appear. Within this interval, the $\epsilon$-adjoint log canonical singularities are controlled by a finite set of structural types, which we classify completely below.
Theorem 1.1 (Main classification theorem).
Fix $\epsilon\in(0,1/3)$. Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface such that $p$ is $\epsilon$-adjoint log canonical. Let
$\pi\colon(X,\mathcal{F})\to(Y,\mathcal{G})$
be the minimal resolution of $(Y,\mathcal{G},p)$, and write $E=\bigcup_{i=1}^{r}E_{i}$ for the exceptional divisor.
Then $E$ belongs to one of the following structural types:
  1. (1)
    Foliated canonical configurations. All components are $\mathcal{F}$-invariant, including $\mathcal{F}$-chains, two (1)(-1)-$\mathcal{F}$-curves of self-intersection 2-2 joined by a bad tail, chains of (2)(-2)-$\mathcal{F}$-curves, $\mathcal{F}$-dihedral forks, and elliptic Gorenstein leaves.
  2. (2)
    Foliated lc (non-canonical) configurations. There exists a smooth non-$\mathcal{F}$-invariant curve $E_{0}$ with ${\rm tang}(\mathcal{F},E_{0})=0$, and all other components (if any) form $\mathcal{F}$-chains attached to $E_{0}$. The curve $E_{0}$ is either rational (with attached chains) or elliptic (in which case $E=E_{0}$).
  3. (3)
    Boundary configurations. A smooth rational $\mathcal{F}$-invariant curve $C$ with ${\rm Z}(\mathcal{F},C)=3$ connecting two $\mathcal{F}$-chains of type $(2,3)$, possibly together with an additional chain of (2)(-2)-$\mathcal{F}$-curves. (Here, by an $\mathcal{F}$-chain $\Theta$ of type $m$ we mean that $m$ denotes the absolute value of the determinant of the intersection matrix of $\Theta$.)
Configurations of type (1) and (2) are precisely the foliated log canonical singularities. Type (3) is not foliated log canonical. It becomes $\epsilon$-adjoint log canonical precisely at $\epsilon=1/5$, and does not occur for $\epsilon<1/5$.
A complete classification, with all subtypes and details, is given in Theorem 4.15.
1.1. Log canonical stability
The boundary configuration enters the admissible region precisely at $\epsilon=1/5$, marking the first wall in the $\epsilon$-direction and the first stability transition in the adjoint log canonical setting. Below this value, adjoint log canonicity forces foliated log canonicity, whereas at $\epsilon=1/5$ the boundary configuration becomes admissible.
Proposition 1.2.
[Corollary 4.16+Example 6.3] Fix $\epsilon\in(0,1/5)$. Suppose $p$ is an $\epsilon$-adjoint log canonical singularity of $(X,\mathcal{F})$. Then:
  • $p$ is log canonical both as a foliated singularity and as a surface singularity;
  • moreover, if $p$ is $\epsilon$-adjoint klt, then $p$ is klt as a singularity of $X$.
The threshold $\epsilon=1/5$ is sharp: at this value, the $\epsilon$-adjoint log canonical foliated singularity may fail to be log canonical as a foliated singularity.
Remark 1.3.
Proposition 1.2 extends previous results: [SS23, Lemma 2.19] treats the case where $p$ is smooth, and [Vas25, Lemma 2.13] considers the case where $p$ is klt as a surface singularity.
1.2. Log-to-canonical stability
In contrast to the log canonical case, imposing the adjoint canonical condition further restricts the admissible configurations. The boundary configuration does not occur if $\epsilon<1/4$ and the singularity is $\epsilon$-adjoint canonical. Thus $\epsilon=1/4$ marks the first wall in the adjoint canonical setting. Below this value, adjoint canonicity forces foliated log canonicity, and the admissible configurations form a strictly smaller class, which we classify completely (see Theorem 5.7).
Analogous to Proposition 1.2, we also have the following log-to-canonical stability proposition.
Proposition 1.4.
[Corollary 5.9+Example 6.3] Fix $\epsilon\in(0,1/4)$. Suppose $p$ is an $\epsilon$-adjoint canonical singularity of $(X,\mathcal{F})$. Then $p$ is a log canonical singularity of $\mathcal{F}$ and a klt singularity of $X$.
The threshold $\epsilon=1/4$ is sharp: at this value, the $\epsilon$-adjoint canonical foliated singularity may fail to be log canonical as a foliated singularity.
1.3. Applications to the adjoint minimal model program
As a direct consequence of Proposition 1.4, we obtain an adjoint minimal model program in the range $\epsilon\in(0,1/4)$. In particular, the threshold $1/4$ determines precisely the maximal range in which the adjoint MMP preserves the expected singularity behavior.
Proposition 1.5.
[See Section 5.1] Let $X$ be a smooth projective surface and $\mathcal{F}$ a rank one foliation with canonical singularities. Then for any $\epsilon\in(0,1/4)\cap\mathbb{Q}$, there exists a birational morphism $\varphi:X\to Y$ (which is a $(K_{\mathcal{F}}+\epsilon K_{X})$-non-positive contraction) such that either
  1. (1)
    $K_{\mathcal{G}}+\epsilon K_{Y}$ is nef, where $\mathcal{G}=\varphi_{*}\mathcal{F}$; or
  2. (2)
    there exists a morphism $f:Y\to Z$ such that $\rho(Y/Z)=1$ and $(K_{\mathcal{G}}+\epsilon K_{Y})$ is $f$-ample,
where $Y$ has klt singularities and $\mathcal{G}$ has log canonical singularities. Moreover, if $K_{\mathcal{F}}+\epsilon K_{X}$ is big, then $K_{\mathcal{G}}+\epsilon K_{Y}$ is ample.
Remark 1.6.
Proposition 1.5 extends [SS23, Theorems 1.1 and 1.2], where the case $\epsilon\in(0,1/5)$ was treated. It also complements [LWX25, Theorem 1.1].
1.4. Interpolated lc thresholds and gap phenomena
The threshold $1/5$ admits a natural interpretation in terms of the interpolated lc threshold, introduced by McKernan [McK22] (see also [CHL+24, Appendix A]).
For a foliated surface $(X,\mathcal{F})$, define
$t_{0}:=\sup\left\{\,t\in[0,1]\mid tK_{\mathcal{F}}+(1-t)K_{X}\text{ is lc}\,\right\}.$
Noting that $K_{\mathcal{F}}+\epsilon K_{X}$ corresponds to the above expression via $\epsilon=\frac{1-t}{t}$ for $t>0$, Proposition 1.2 immediately implies:
Proposition 1.7.
The interpolated lc threshold $t_{0}$ satisfies
$t_{0}=1\quad\text{or}\quad t_{0}\leq\frac{5}{6}.$
More generally, the following 1-gap conjecture for interpolated lc thresholds was formulated in [CHL+24, Conjecture A.6]:
Conjecture 1.8 (1-gap for interpolated lc thresholds).
Let $d$ be a positive integer. Then there exists a positive real number $\tau=\tau(d)$ satisfying the following property:
Let $(X,\mathcal{F},t)$ be an adjoint foliated structure of dimension $d$. Assume that $tK_{\mathcal{F}}+(1-t)K_{X}$ is log canonical for some $t>1-\tau$ and that $K_{\mathcal{F}}$ is $\mathbb{Q}$-Cartier. Then $\mathcal{F}$ is log canonical.
In dimension 2, our results show that the constant is
$\tau(2)=\frac{1}{6},$
and that this value is sharp.
Outlook. Our results reveal a sharp first threshold for adjoint singularities in dimension two. It is natural to ask whether analogous threshold behavior occurs in higher dimensions, and how such gap principles may interact with boundedness, moduli problems, or the birational geometry of foliated pairs.
Organization of the paper. Section 2 reviews background on foliations and singularities. Section 3 establishes structural properties of negative definite configurations. Section 4 classifies $\epsilon$-adjoint log canonical singularities for $\epsilon\in(0,1/3)$ and determines the threshold $1/5$. Section 5 studies $\epsilon$-adjoint canonical singularities, determines the threshold $1/4$, and proves the adjoint minimal model program (MMP). Section 6 presents examples illustrating the sharpness of both thresholds.
2. Preliminaries
2.1. Foliations and singularities
A foliation $\mathcal{F}$ on a normal surface $X$ is a rank 1 saturated subsheaf $T_{\mathcal{F}}$ of the tangent sheaf $T_{X}$. Note that $T_{X}$ is reflexive as it is isomorphic to ${\rm Hom}_{\mathcal{O}_{X}}(\Omega_{X},\mathcal{O}_{X})$. Thus $T_{\mathcal{F}}$ is also reflexive and we can define the canonical divisor $K_{\mathcal{F}}$ of the foliation as a Weil divisor on $X$ satisfying $\mathcal{O}_{X}(-K_{\mathcal{F}})\cong T_{\mathcal{F}}$. A foliated surface $(X,\mathcal{F})$ consists of a normal surface $X$ together with a foliation $\mathcal{F}$ on $X$.
A point $p\in X$ is called a singularity of the foliation $\mathcal{F}$ if either a singular point of $X$ or a point at which the quotient $T_{X}/T_{\mathcal{F}}$ is not locally free. We say $p$ is a regular point of $\mathcal{F}$ if $p$ is a smooth point of $X$ and $T_{X}/T_{\mathcal{F}}$ is locally free at $p$.
Let $p$ be a singularity of $\mathcal{F}$, at which $X$ is smooth. Locally, $\mathcal{F}$ is defined by a vector field
(2.1) $\nu=A(x,y)\frac{\partial}{\partial x}+B(x,y)\frac{\partial}{\partial y},$
where $p=(0,0)$. The two eigenvalues $\lambda_{1},\lambda_{2}$ of the linear part $(D\nu)(p)$ of $\nu$ at $p$ are well defined.
Definition 2.1.
Let $p$ be a singularity of $\mathcal{F}$ (at which $X$ is smooth) and let $\lambda_{1},\lambda_{2}$ be as above.
  1. (1)
    The singularity $p$ is called non-degenerated if the two eigenvalues $\lambda_{1},\lambda_{2}$ are both nonzero.
  2. (2)
    The singularity $p$ is called reduced if one of the two eigenvalues, say $\lambda_{2}$, is nonzero and the quotient $\lambda=\lambda_{1}/\lambda_{2}$ is not a positive rational number. In particular, if $\lambda=0$, we call $p$ a saddle-node.
A foliation $\mathcal{F}$ is said to be reduced if any singularity of $\mathcal{F}$ is reduced.
Remark 2.2.
If $p$ is a saddle-node, by a suitable transition of coordinates, the foliation $\mathcal{F}$ at $p$ can be represented by
$\nu$ $=\left(x+axy^{k}+yF(x,y)\right)\frac{\partial}{\partial x}+y^{k+1}\frac{\partial}{\partial y},$
$\text{or}\quad\omega$ $=\left(x+axy^{k}+yF(x,y)\right){\rm d}y-y^{k+1}{\rm d}x,$
where $a\in\mathbb{C}$, $k\in\mathbb{Z}^{+}$, and $F$ is a holomorphic function which vanishes at $p=(0,0)$ up to order $k$. The curve $(y=0)$ is a separatrix, called strong separatrix. If $F=0$, then the separatrix $(x=0)$ is called the weak separatrix.
Theorem 2.3 (Seidenberg).
Given any foliated surface $(X,\mathcal{F})$, there exists a sequence of blowing-ups $\sigma:X^{\prime}\to X$, such that the induced foliation $\mathcal{F}^{\prime}$ on $X^{\prime}$ is reduced.
Proof.
See [Sei68] or [Bru15, Theorem 1.1]. ∎
2.2. Minimal resolution of foliated surfaces
Definition 2.4.
Let $(X,\mathcal{F})$ be a foliated surface. A curve $C\subset X$ is said to be $\mathcal{F}$-exceptional if the following conditions are satisfied:
  1. (1)
    $C$ is a smooth rational curve with self-intersection 1-1;
  2. (2)
    the contraction of $C$ to a point $p$ yields a foliated surface $(X^{\prime},\mathcal{F}^{\prime})$, where $p$ is either a regular point or a reduced singularity of $\mathcal{F}^{\prime}$.
Definition 2.5.
Let $(X,\mathcal{F})$ be a foliated surface and let $p$ be a singularity of $(X,\mathcal{F})$. A resolution of $(X,\mathcal{F},p)$ is a birational morphism
$\pi:(Y,\mathcal{G})\longrightarrow(X,\mathcal{F}),$
such that
  • (i)
    $Y-\pi^{-1}(p)\cong X-p$,
  • (ii)
    for any $\in\pi^{-1}(p), qq$ is a smooth point of $Y$ and $q$ is a reduced (or regular) point of the induced foliation $\mathcal{G}:=\pi^{*}\mathcal{F}$.
We say the resolution $\pi$ is minimal if moreover
  • (iii)
    there exists no $\mathcal{G}$-exceptional curve contained in $\pi^{-1}(p)$.
2.3. Index Theorems
Let $(X,\mathcal{F})$ be a foliated surface with $X$ smooth.
A reduced curve $C\subseteq X$ is said to be $\mathcal{F}$-invariant if the inclusion $T_{\mathcal{F}}|_{C}\rightarrow T_{X}|_{C}$ factors through $T_{C}$, where $T_{C}$ is the tangent bundle of $C$.
2.3.1. Non-invariant curves
For a non-$\mathcal{F}$-invariant curve $C$ and $p\in C$, we can define the tangency order of $\mathcal{F}$ along $C$ at $p$ to be
${\rm tang}(\mathcal{F},C,p):=\dim_{\mathbb{C}}\frac{\mathcal{O}_{p}}{\langle f,\nu(f)\rangle},$
where $\nu$ is the local generator of $\mathcal{F}$ at $p$ and $f=0$ is the local equation of $C$. We define
${\rm tang}(\mathcal{F},C):=\sum_{p\in C}{\rm tang}(\mathcal{F},C,p).$
Proposition 2.6.
We have
(2.2) ${\rm tang}(\mathcal{F},C)=K_{\mathcal{F}}\cdot C+C^{2}\,(\geq 0).$
Proof.
See [Bru15, Proposition 2.2]. ∎
Remark 2.7.
${\rm tang}(\mathcal{F},C)=0$ implies that $C$ is smooth.
2.3.2. Invariant curves
Let $C$ be an $\mathcal{F}$-invariant curve on $X$, and let $p\in C$ be a point. In a neighborhood of $p$, the foliation $\mathcal{F}$ is defined by a 1-form $\omega$, and $C$ is locally given by the equation $f=0$. Since $C$ is $\mathcal{F}$-invariant, we may write
$g\omega=hdf+f\eta,$
where $\eta$ is a holomorphic 1-form and $g,h$ are holomorphic functions defined around $p$, with $h$ and $f$ are coprime. We define
(2.3) ${\rm Z}(\mathcal{F},C,p)$ $:=\text{vanishing order of $\left.\frac{h}{g}\right|_{C}$ at $p$}$
(2.4) ${\rm CS}(\mathcal{F},C,p)$ $:={\rm Res}_{p}\left\{\left.-\frac{\eta}{h}\right|_{C}\right\}.$
By definition, both ${\rm Z}(\mathcal{F},C,p)$ and ${\rm CS}(\mathcal{F},C,p)$ vanish if $p$ is not a singularity of $\mathcal{F}$. If $p$ is a reduced singularity of $\mathcal{F}$, then ${\rm Z}(\mathcal{F},C,p)\geq 0$. Let
${\rm Z}(\mathcal{F},C):=\sum_{p\in C}{\rm Z}(\mathcal{F},C,p),\quad{\rm CS}(\mathcal{F},C):=\sum_{p\in C}{\rm CS}(\mathcal{F},C,p).$
Proposition 2.8.
Let $C$ be an $\mathcal{F}$-invariant curve on $X$. Then
(2.5) ${\rm Z}(\mathcal{F},C)$ $=K_{\mathcal{F}}\cdot C+2-2p_{a}(C),$
(2.6) ${\rm CS}(\mathcal{F},C)$ $=C^{2},$
where $p_{a}(C)$ denotes the arithmetic genus of $C$. The second equality above is called the Camacho-Sad formula.
Proof.
See [Bru15, Proposition 2.3 and Theorem 3.2]. ∎
Lemma 2.9.
Let $p$ is a reduced singularity of the foliated surface $(X,\mathcal{F})$.
  1. (1)
    If $p$ is nondegerate and $\mathcal{F}$ locally around $p$ defined by $\omega=\lambda y(1+o(1)){\rm d}x-x(1+o(1)){\rm d}y$, then
    $\begin{cases}{\rm CS}(\mathcal{F},x=0,p)=1/\lambda,\qquad{\rm CS}(\mathcal{F},y=0,p)=\lambda,\\ {\rm Z}(\mathcal{F},x=0,p)={\rm Z}(\mathcal{F},y=0,p)=1\end{cases}$
  2. (2)
    If $p$ is a saddle-node and $\mathcal{F}$ locally around $p$ defined by $\omega=y^{k+1}{\rm d}x-(x(1+ay^{k})+yo(k)){\rm d}y$, where $k\in\mathbb{Z}^{>0}$ and $a\in\mathbb{C}$, then
    ${\rm CS}(\mathcal{F},y=0,p)=0,\quad\text{and}\quad{\rm Z}(\mathcal{F},y=0,p)=1.$
    If furthermore there exists a weak separatrix $(x=0)$, then
    ${\rm CS}(\mathcal{F},x=0,p)=a,\quad\text{and}\quad{\rm Z}(\mathcal{F},x=0,p)=k+1\,(\geq 2).$
Proof.
See [Bru15, pp. 30-31]. ∎
Next, we recall the separatrix theorem.
Theorem 2.10 (Separatrix theorem).
Let $\mathcal{F}$ be a foliation on a smooth projective surface $X$ and let $C\subset X$ be a connected compact $\mathcal{F}$-invariant curve such that:
  • (i)
    All the singularities of $\mathcal{F}$ on $C$ are reduced (in particular, $C$ has only normal crossing singularities);
  • (ii)
    If $C_{1},\cdots,C_{n}$ are the irreducible components of $C$, then the intersection matrix $(C_{i}C_{j})_{1\leq i,j\leq n}$ is negative definite and the dual graph $\Gamma$ is a tree.
Then there exists at least one point $p\in C\cap{\rm Sing}(\mathcal{F})$ and a separatrix through $p$ not contained in $C$.
Proof.
See [Bru15, Theorem 3.4]. ∎
2.4. Pairs and Triples
Let $X$ be a normal surface and let $\mathcal{F}$ be a foliation on $X$. Let $D$ be an $\mathbb{R}$-divisor on $X$. Then $D$ admits a unique decomposition
$D=D^{\mathrm{inv}}_{\mathcal{F}}+D^{\mathrm{n\!-\!inv}}_{\mathcal{F}},$
where $D^{\mathrm{inv}}_{\mathcal{F}}$ (resp. $D^{\mathrm{n\!-\!inv}}_{\mathcal{F}}$) denotes the $\mathcal{F}$-invariant part (resp. the non-$\mathcal{F}$-invariant part) of $D$, namely the sum of the components of $D$ whose support is $\mathcal{F}$-invariant (resp. non-$\mathcal{F}$-invariant). If there is no risk of confusion, we will omit the subscript $\mathcal{F}$ from the notation.
We say that $(X,\Delta)$ is a pair if $\Delta$ is an effective $\mathbb{R}$-divisor and $K_{X}+\Delta$ is $\mathbb{R}$-Cartier. A foliated pair $(\mathcal{F},\Delta)$ on $X$ consists of a foliation $\mathcal{F}$ on $X$ and an effective $\mathbb{R}$-divisor $\Delta$ such that $K_{\mathcal{F}}+\Delta$ is $\mathbb{R}$-Cartier.
Definition 2.11.
A foliated triple $(X,\mathcal{F},\Delta)$ consists of a foliation $\mathcal{F}$ on a normal surface $X$ and an effective $\mathbb{R}$-divisor $\Delta\geq 0$ such that both $K_{\mathcal{F}}+\Delta^{\mathrm{n\!-\!inv}}$ and $K_{X}+\Delta$ are $\mathbb{R}$-Cartier.
2.5. Log canonical foliated singularities
Let $\pi\colon X^{\prime}\to X$ be a birational morphism and let $(\mathcal{F},\Delta)$ be a foliated pair on $X$. Denote by $\mathcal{F}^{\prime}$ the induced foliation on $X^{\prime}$. Then we may write
$K_{\mathcal{F}^{\prime}}=\pi^{*}(K_{\mathcal{F}}+\Delta)+\sum_{E}a(E,\mathcal{F},\Delta)\,E,$
where the sum runs over all prime divisors $E$ on $X^{\prime}$ and
$\pi_{*}\!\left(\sum_{E}a(E,\mathcal{F},\Delta)\,E\right)=-\Delta.$
The rational number $a(E,\mathcal{F},\Delta)$ is called the discrepancy of the foliated pair $(\mathcal{F},\Delta)$ with respect to $E$.
Definition 2.12.
Let $X$ be a normal surface and let $(\mathcal{F},\Delta)$ be a foliated pair on $X$. We say that $(\mathcal{F},\Delta)$ is terminal (resp. canonical, klt, log canonical) if
$a(E,\mathcal{F},\Delta)>0\quad(\text{resp.\ }\geq 0,\ >-\iota(E)\ \text{and }\lfloor\Delta\rfloor=0,\ \geq-\iota(E))$
for any birational morphism $\pi\colon X^{\prime}\to X$ and for any prime divisor $E$ on $X^{\prime}$, where
$\iota(E):=\begin{cases}1,&\text{if $E$ is not $\mathcal{F}^{\prime}$-invariant},\\ 0,&\text{if $E$ is $\mathcal{F}^{\prime}$-invariant}.\end{cases}$
Moreover, we say that $(\mathcal{F},\Delta)$ is log terminal if
$a(E,\mathcal{F},\Delta)>-\iota(E)$
for any birational morphism $\pi\colon X^{\prime}\to X$ and for any $\pi$-exceptional prime divisor $E$ on $X^{\prime}$.
2.6. $(\epsilon,\delta)$-adjoint log canonical foliated singularities
Definition 2.13.
Let $(X,\mathcal{F},\Delta)$ be a foliated triple on a normal surface $X$, and fix $\epsilon\geq 0$. The $\epsilon$-adjoint divisor of $(X,\mathcal{F},\Delta)$ is defined by
$K_{(X,\mathcal{F},\Delta),\epsilon}:=(K_{\mathcal{F}}+\Delta^{\mathrm{n\!-\!inv}})+\epsilon(K_{X}+\Delta).$
Definition 2.14.
Let $(X,\mathcal{F},\Delta)$ be a foliated triple. Fix $\epsilon>0$ and $\delta\geq 0$. We say that $(X,\mathcal{F},\Delta)$ is $(\epsilon,\delta)$-adjoint log canonical (resp. $(\epsilon,\delta)$-adjoint klt) if for any birational morphism $\pi\colon X^{\prime}\to X$, writing
$K_{(X^{\prime},\mathcal{F}^{\prime},\Delta^{\prime}),\epsilon}=\pi^{*}K_{(X,\mathcal{F},\Delta),\epsilon}+\sum_{i}a_{i}E_{i},$
where the $E_{i}$ are $\pi$-exceptional divisors and $\Delta^{\prime}:=\pi_{*}^{-1}\Delta$ is the strict transform of $\Delta$, we have, for all $i$,
$a_{i}\geq(\iota(E_{i})+\epsilon)(-1+\delta)$
$(\text{resp.}\quad a_{i}>(\iota(E_{i})+\epsilon)(-1+\delta)\quad\text{and}\quad\lfloor\Delta\rfloor=0).$
When $\delta=1$, we call $(\epsilon,\delta)$-adjoint log canonical $\epsilon$-adjoint canonical. When $\delta=0$, we call $(\epsilon,\delta)$-adjoint log canonical simply $\epsilon$-adjoint log canonical.
Moreover, we say that $(\mathcal{F},\Delta)$ is $\epsilon$-adjoint terminal if $a_{i}>0$ for all $i$.
2.7. Notation and basic configurations
In this subsection, let $(X,\mathcal{F})$ be a foliated surface with $X$ smooth. We introduce notation and basic configurations that will be used throughout the paper.
Definition 2.15 ($\mathcal{F}$-chain).
A divisor $\Theta=\Gamma_{1}+\cdots+\Gamma_{r}$ is called an $\mathcal{F}$-chain if
  1. (1)
    $\Theta$ is a Hirzebruch–Jung string (cf. [BHPV04, Ch.III; Sec.5]);
  2. (2)
    each irreducible component $\Gamma_{j}$ is $\mathcal{F}$-invariant;
  3. (3)
    ${\rm Sing}(\mathcal{F})\cap\Theta$ consists of reduced, non-degenerate singularities;
  4. (4)
    ${\rm Z}(\mathcal{F},\Gamma_{1})=1$ and ${\rm Z}(\mathcal{F},\Gamma_{i})=2$ for all $i\geq 2$.
The curve $\Gamma_{1}$ is called the first curve of the $\mathcal{F}$-chain (see Figure 1).
$\cdots\cdots$
$\Gamma_{2}$
$\Gamma_{1}$
$\Gamma_{r}$
Figure 1. $\mathcal{F}$-chain with the first curve $\Gamma_{1}$
Definition 2.16.
An irreducible curve $C$ is called a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve) if
  1. (1)
    $C$ is a smooth rational $\mathcal{F}$-invariant curve, and
  2. (2)
    ${\rm Z}(\mathcal{F},C)=1$ (resp. ${\rm Z}(\mathcal{F},C)=2$).
Lemma 2.17.
If $C$ is a (1)(-1)-$\mathcal{F}$-curve with $C^{2}=-1$, then $C$ is $\mathcal{F}$-exceptional. Moreover, there exists exactly one singularity of $\mathcal{F}$ on $C$, which is non-degenerate and reduced.
Lemma 2.18.
Suppose $C$ is a (2)(-2)-$\mathcal{F}$-curve with $C^{2}=-1$. Assume that every singularity of $\mathcal{F}$ on $C$ is reduced. If $C$ is not $\mathcal{F}$-exceptional, then at the contraction point $p$ of $C$, the induced foliation is locally defined by
$\omega=(x+y)\,{\rm d}x-x\,{\rm d}y.$
Moreover, there exists exactly one singularity of $\mathcal{F}$ on $C$, which is a saddle-node $q$. In a neighborhood of $q$, the foliation $\mathcal{F}$ is defined by
$\omega^{\prime}=y(s+1)\,{\rm d}s+s^{2}\,{\rm d}y,$
where $C=(y=0)$ is the weak separatrix of $q$.
Proof.
See [Bru15, pp. 7–8]. ∎
Definition 2.19.
A curve $C$ is called a bad tail if
  1. (1)
    $C$ is a smooth rational irreducible $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$ and $C^{2}\leq-2$;
  2. (2)
    $C$ intersects two (1)(-1)-$\mathcal{F}$-curves whose self-intersections are both 2-2.
(See Figure 2.)
$C$
2-2
2-2
Figure 2. $C$ is a bad tail
Definition 2.20.
An $\mathcal{F}$-dihedral fork is a divisor
$F=\Gamma_{1}+\Gamma_{2}+\cdots+\Gamma_{r},$
where
  1. (1)
    $\Gamma_{1}$ and $\Gamma_{2}$ are (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2;
  2. (2)
    $\Gamma_{3}$ is a bad tail, which is attached to a chain of (2)(-2)-$\mathcal{F}$-curves $\Gamma_{4}+\cdots+\Gamma_{r}$.
(See Figure 3.)
$\Gamma_{1}(-2)$
$\cdots\cdots$
$\Gamma_{2}(-2)$
$\Gamma_{4}$
$\Gamma_{3}$
$\Gamma_{r}$
Figure 3. $\mathcal{F}$-dihedral fork
Definition 2.21 (egl).
An elliptic Gorenstein leaf is either a rational $\mathcal{F}$-invariant curve with a single node, or a cycle of (2)(-2)-$\mathcal{F}$-curves. (See Figure 4.)
or
$\vdots$
Figure 4. An elliptic Gorenstein leaf
Definition 2.22 ($\mathcal{F}$-star graph).
We say that
$F=C+\sum_{i=1}^{s}\Theta_{i}\qquad(s\geq 0)$
is an $\mathcal{F}$-star graph (centered at $C$) if
  1. (1)
    $C$ is a smooth irreducible non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$;
  2. (2)
    for each $i=1,\dots,s$, $\Theta_{i}=\sum_{j=1}^{r(i)}\Gamma_{ij}$ is an $\mathcal{F}$-chain whose first curve $\Gamma_{i1}$ satisfies
    $C\cdot\Theta_{i}=C\cdot\Gamma_{i1}=1.$
We say that an $\mathcal{F}$-star graph is of type $(s;m_{1},\dots,m_{s})$, where
$m_{i}=\det\!\bigl(-\Gamma_{ij}\cdot\Gamma_{ik}\bigr)_{1\leq j,k\leq r(i)}.$
The curve $C$ is called the center of the $\mathcal{F}$-star graph (see Figure 5).
Moreover, if $s\leq 2$, we call $F$ an $\mathcal{F}$-star chain; if $s\geq 3$, we call $F$ is an $\mathcal{F}$-star fork.
$C$
$\Theta_{1}$
$\cdots\cdots$
$\Theta_{2}$
$\cdots\cdots$
$\Theta_{s}$
$\cdots\cdots$
$\cdots\cdots\cdots\cdots$
Figure 5. $\mathcal{F}$-star graph centered at $C$
$C$
$\Theta_{1}$
$\Theta_{2}$
$\cdots\cdots$
$\cdots\cdots$
Figure 6. $\mathcal{F}$-star chain centered at $C$, where $\Theta_{1}$ and $\Theta_{2}$ may be empty.
Remark 2.23.
The configurations introduced above arise naturally in the birational geometry of foliated surfaces. They will serve as the basic configurations throughout this paper.
2.8. Classification of lc surfaces singularities and lc foliated surface singularities
In this subsection, we recall the classification of log canonical surface singularities and log canonical foliated surface singularities, serving as a reference point for comparison with the $\epsilon$-adjoint setting.
Theorem 2.24.
Let $p$ be a log canonical singularity of a normal surface $Y$. Let
$\pi:X\to Y$
be the minimal resolution of $(Y,p)$, with the exceptional divisor $E=\cup_{i}E_{i}$. Then $E$ is one of the following:
  1. (1)
    $E$ is a Hirzebruch-Jung chain.
  2. (2)
    $E=C$ is an irreducible curve, where $C$ is a smooth elliptic curve or a rational curve with a single node.
  3. (3)
    $E$ is a cycle of smooth rational curves.
  4. (4)
    $E$ has two forks with the dual graph as follows:
[…]
    where each component is a smooth rational curve with self-intersection at most 2-2.
  5. (5)
    $E$ is a single fork centered at a smooth rational curve $C$, and $E\setminus C$ are pairwise disjoint Hirzebruch-Jung chains $\Theta_{1},\cdots,\Theta_{l}$ of type $(m_{1},\cdots,m_{l})$, which belongs to one of
    1. (a)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, or $(2,3,4)$, or $(2,3,5)$;
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, $(2,4,4)$, $(3,3,3)$.
    3. (c)
      $l=4$ and $(m_{1},m_{2},m_{3},m_{4})=(2,2,2,2)$.
Here, types (2), (3), (4), (5-b), and (5-c) are not klt, while types (1) and (5-1) are klt.
Proof.
See [Alex92] or [KM98, Sec. 4.1]. ∎
Theorem 2.25.
Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface with log canonical foliated singularity at $p$, and let $\pi:(X,\mathcal{F})\to(Y,\mathcal{G})$ be the minimal resolution with exceptional divisors $E=\cup_{i}E_{i}$. Then $E$ is one of the following:
  1. (1)
    An $\mathcal{F}$-chain. (See Figure 1.)
  2. (2)
    A chain of (2)(-2)-$\mathcal{F}$-curves. (See Definition 2.16.)
  3. (3)
    A chain of three $\mathcal{F}$-invariant curves $E_{1}+E_{2}+E_{3}$, where $E_{1},E_{3}$ are (1)(-1)-$\mathcal{F}$-curves and $E_{2}$ is a bad tail. (See Figure 2.)
  4. (4)
    An $\mathcal{F}$-dihedral fork: two (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2 joined by a bad tail connecting to a chain of (2)(-2)-$\mathcal{F}$-curves. (See Figure 3.)
  5. (5)
    An elliptic Gorenstein leaf: either a rational $\mathcal{F}$-invariant curve with a single node or a cycle of (2)(-2)-$\mathcal{F}$-curves. (See Figure 4.)
  6. (6)
    An $\mathcal{F}$-star graph centered at a smooth non-$\mathcal{F}$-invariant curve $C$ with ${\rm tang}(\mathcal{F},C)=0$. (See Definition 2.22 and Figure 5.)
Here, type (1) is terminal and types (1)–(5) are canonical (as foliated singularities).
Proof.
See [Chen23, Theorem 1.1]. We note that cases (6) and (7) in [Chen23, Theorem 1.1] are combined into a single case, namely case (6) in Theorem 2.25. ∎
2.9. A negativity criterion on surfaces
Lemma 2.26.
Let $D=\sum_{i=1}^{n}a_{i}C_{i}$ be a $\mathbb{Q}$-divisor such that the intersection matrix
$(C_{i}\cdot C_{j})_{1\leq i,j\leq n}$
is negative definite.
  1. (1)
    If $D\cdot C_{i}\leq 0$ for all $i=1,\dots,n$, then $D\geq 0$.
  2. (2)
    If $E$ is an effective $\mathbb{Q}$-divisor and $(E-D)\cdot C_{j}\leq 0$ for all $j=1,\dots,n$, then $E-D\geq 0$.
Proof.
Both statements follow from [Luc01, Lemma 14.9 and Lemma 14.15]. ∎
3. Negative definite configurations of curves
In this section, we introduce useful notations and properties for negative definite configurations of curves on a smooth surface, which will be applied in the next section.
Throughout this section, let $\mathfrak{E}=\bigcup_{i=1}^{n}E_{i}$ be a finite collection of irreducible curves on a smooth surface $X$ whose intersection matrix $(E_{i}\cdot E_{j})$ is negative definite. Set $E:=E_{1}+\cdots+E_{n}$.
3.1. Dual graphs and intersection equations
Associated to $\mathfrak{E}$ is its weighted dual graph $G_{E}$, whose vertices correspond to the curves $E_{i}$, with an edge of weight $E_{i}\cdot E_{j}$ between distinct vertices and with vertex weight $E_{i}^{2}$. A subgraph of $G_{E}$ is the dual graph of a divisor $F\subset E$.
Let $D$ be a divisor on $X$, and let $F=C_{1}+\cdots+C_{r}\subset E$ be a divisor whose intersection matrix is negative definite. Since the intersection matrix is invertible, the system of linear equations
$\sum_{i=1}^{r}a_{i}\,C_{i}\cdot C_{j}=D\cdot C_{j},\qquad j=1,\dots,r,$
admits a unique solution $(a_{1},\dots,a_{r})$.
Definition 3.1.
The divisor $M(D,F)$ is the unique divisor supported on $F$ satisfying
$M(D,F)\cdot C_{i}=D\cdot C_{i}\quad\text{for all }i=1,\dots,r.$
The divisor $M(D,F)$ can be viewed as the intersection-theoretic projection of $D$ onto the negative definite subconfiguration $F$.
We say that $F$ (or $G_{F}$) is a $D^{\geq 0}$-graph (resp. $D^{>0}$-graph) if the coefficients of $M(D,F)$ are all non-negative (resp. positive). If $F=C$ is irreducible, then $C$ is a $D^{>0}$-curve if and only if $D\cdot C<0$.
Lemma 3.2.
Assume the negative definite curves $F=\sum_{i=1}^{r}E_{i}$ is connected. If $DE_{i}\leq 0$ for all $i$, then $F$ is a $D^{\geq 0}$-graph. Moreover, $F$ is a $D^{>0}$-graph if and only if $DE_{i}<0$ for some $i$.
Proof.
See [KM98, Lemma 3.41]. ∎
3.2. $D^{>0}$-chains
A divisor $F=C_{1}+\cdots+C_{r}\subset E$ is called a chain if $C_{i}\cdot C_{i+1}=1$ and $C_{i}\cdot C_{j}=0$ for $|i-j|\geq 2$. If moreover $F$ is a $D^{\geq 0}$- (resp. $D^{>0}$-)graph, we call it a $D^{\geq 0}$- (resp. $D^{>0}$-)chain.
Lemma 3.3.
Let $F=C_{1}+\cdots+C_{r}$ be a $D^{>0}$-chain and let $C\subset E$ be a curve not contained in $F$ such that $F\cdot C=C_{r}\cdot C=1$. Then $F+C$ is a $D^{>0}$-chain if and only if
$D\cdot C<M(D,F)\cdot C.$
Proof.
Write
$M(D,F)=\sum_{i=1}^{r}\gamma_{i}C_{i},\qquad M(D,F+C)=\sum_{i=1}^{r+1}\gamma^{\prime}_{i}C_{i},$
where $C_{r+1}:=C$. Then
$\bigl(M(D,F+C)-M(D,F)\bigr)\cdot C_{i}=\begin{cases}0,&i=1,\dots,r,\\ D\cdot C-M(D,F)\cdot C,&i=r+1.\end{cases}$
If $D\cdot C<M(D,F)\cdot C$, then
$\bigl(M(D,F+C)-M(D,F)\bigr)\cdot C_{i}\leq 0\quad\text{for }i=1,\cdots,r+1,$
and hence Lemma 2.26 implies $M(D,F+C)>M(D,F)$; in particular, $\gamma^{\prime}_{r+1}>0$. Thus $\gamma^{\prime}_{i}>0$ for all $i=1,\dots,r+1$, and $F+C$ is a $D^{>0}$-chain.
Conversely, if $D\cdot C\geq M(D,F)\cdot C$, then $\gamma^{\prime}_{r+1}\leq 0$, so $F+C$ is not a $D^{>0}$-chain. ∎
Remark 3.4.
An analogous statement holds for $D^{\geq 0}$-chains with the inequality $D\cdot C\leq M(D,F)\cdot C$.
For later numerical control along chains, we introduce the determinants
(3.1) $\begin{cases}\mu_{0}=\lambda_{r+1}:=1,\qquad\mu_{1}=\lambda_{r}:=1,\\ \mu_{k}:=\det(-C_{i}\cdot C_{j})_{1\leq i,j\leq k-1},&k=2,\dots,r+1,\\ \lambda_{l}:=\det(-C_{i}\cdot C_{j})_{l+1\leq i,j\leq r},&l=0,\dots,r-1,\\ n:=\lambda_{0}=\mu_{r+1}.\end{cases}$
3.3. Special $D^{>0}$-chains
We now isolate the configurations that play a decisive role in the later classification.
Definition 3.5.
A $D^{>0}$- (resp. $D^{\geq 0}$-)chain $F=C_{1}+\cdots+C_{r}$ is called special if
$D\cdot C_{1}<0\quad\text{and}\quad D\cdot C_{i}\geq 0\ \text{for }i\geq 2.$
Here $C_{1}$ is called the first curve of $F$.
The following proposition provides explicit and uniform bounds for the coefficients of $M(D,F)$ along special $D^{>0}$-chains. More precisely, each coefficient is controlled solely by the initial negativity $D\cdot C_{1}$ and the combinatorial structure of the chain.
Proposition 3.6.
Let $F=C_{1}+\cdots+C_{r}$ be a special $D^{>0}$-chain. Write
$M(D,F)=\sum_{i=1}^{r}\gamma_{i}C_{i}.$
Then:
  • (1)
    For $k=1,\dots,r$, we have
    $\gamma_{k}\leq\frac{(-D\cdot C_{1})\lambda_{k}}{n}.$
  • (2)
    Let $F_{k}:=C_{1}+\cdots+C_{k}$ for $k=1,\dots,r$. Then for $k\geq 2$,
    $D\cdot C_{k}<M(D,F_{k-1})\cdot C_{k}\leq\frac{-D\cdot C_{1}}{\mu_{k}}.$
    Moreover, each $F_{k}$ is a special $D^{>0}$-chain, satisfying
    $0<M(D,F_{1})<M(D,F_{2})<\cdots<M(D,F_{r})=M(D,F).$
Proof.
(1) Let $M_{1}(F)$ be the unique effective $\mathbb{Q}$-divisor supported on $F$ such that
$M_{1}(F)\cdot C_{1}=-1,\qquad M_{1}(F)\cdot C_{i}=0\ \text{ for }i\geq 2.$
A straightforward computation shows that
$M_{1}(F)=\sum_{k=1}^{r}\frac{\lambda_{k}}{n}C_{k}.$
On the other hand, we have
$\big[(-D\cdot C_{1})M_{1}(F)-M(D,F)\big]\cdot C_{i}\leq 0,\qquad i=1,\dots,r.$
By Lemma 2.26, this implies
$M(D,F)\leq(-D\cdot C_{1})M_{1}(F)=\sum_{k=1}^{r}\frac{(-D\cdot C_{1})\lambda_{k}}{n}C_{k}.$
This proves (1).
(2) For $k\leq r-1$, we compute
$\big[M(D,F)-M(D,F_{k})\big]\cdot C_{i}=\begin{cases}0,&\text{if }i\leq k,\\[4.0pt] D\cdot C_{k+1}-M(D,F_{k})\cdot C_{k+1},&\text{if }i=k+1,\\[4.0pt] D\cdot C_{i}\ (\geq 0),&\text{if }i\geq k+2.\end{cases}$
If
$D\cdot C_{k+1}-M(D,F_{k})\cdot C_{k+1}\geq 0,$
then Lemma 2.26 implies $M(D,F)\leq M(D,F_{k})$, hence $\gamma_{i}\leq 0$ for all $i\geq k+1$, contradicting the assumption that $F$ is a $D^{>0}$-chain. Therefore,
(3.2) $D\cdot C_{k}<M(D,F_{k-1})\cdot C_{k}\leq\frac{-D\cdot C_{1}}{\mu_{k}},\qquad k\geq 2,$
where the second inequality follows from the same argument as in (1).
Note that $F_{1}$ is a special $D^{>0}$-chain since $D\cdot C_{1}<0$. By (3.2) and Lemma 3.3, an induction on $k$ shows that each $F_{k}$ is a special $D^{>0}$-chain and that $M(D,F_{k})$ is strictly increasing. ∎
4. Classification of $\epsilon$-adjoint log canonical singularities for $\epsilon\in(0,1/3)$
In the previous section, we analyzed negative definite configurations of curves on smooth surfaces and established several structural properties of their intersection matrices. We now apply these results to the study of $\epsilon$-adjoint log canonical surface singularities.
Let $(Y,\mathcal{G})$ be a foliated pair on a normal surface, and let $p$ be a singular point of $(Y,\mathcal{G})$. Fix $\epsilon\in(0,1/3)$. Assume that $(Y,\mathcal{G},p)$ is $\epsilon$-adjoint log canonical. Our goal is to determine all possible configurations arising in its minimal resolution.
Let $\pi:X\to Y$ be the minimal resolution of the germ $(Y,\mathcal{G},p)$ (cf. Definition 2.5), and write $\mathfrak{E}=\bigcup_{i\in I}E_{i}$ for the exceptional divisor. Set
$E=\sum_{i\in I}E_{i}.$
Note that the intersection matrix $(E_{i}\cdot E_{j})_{i,j\in I}$ is negative definite. Consequently, the adjoint log canonical condition can be expressed as a system of linear inequalities for the discrepancies along the components $E_{i}$.
We can write
$K_{(X,\mathcal{F},E),\epsilon}=\pi^{*}K_{(Y,\mathcal{G}),\epsilon}+\sum_{i\in I}a_{i}E_{i},$
where $K_{(X,\mathcal{F},E),\epsilon}$ and $K_{(Y,\mathcal{G}),\epsilon}$ are as in Definition 2.13. Intersecting both sides with $E_{j}$, we obtain
(4.1) $\sum_{i\in I}a_{i}\,E_{i}\cdot E_{j}=K_{(X,\mathcal{F},E),\epsilon}\cdot E_{j},\qquad\text{for all $j\in I$}.$
Since the intersection matrix is negative definite, the system (4.1) admits a unique solution $(a_{i})$. Hence $(Y,\mathcal{G},p)$ is $\epsilon$-adjoint log canonical (resp. $\epsilon$-adjoint klt) if and only if
$a_{i}\geq 0\quad(\text{resp.\ }a_{i}>0)\quad\text{for all }i\in I.$
Thus the classification problem reduces to determining all negative definite configurations for which the right-hand side of (4.1) produces a nonnegative (resp. positive) solution. Equivalently, in the terminology of Section 3, we must determine all possible dual graphs of $E$ which are $K^{\geq 0}$-graphs (resp. $K^{>0}$-graphs), where
$K:=K_{(X,\mathcal{F},E),\epsilon},\qquad\epsilon\in(0,\tfrac{1}{3}).$
The remainder of this section is devoted to a systematic analysis of these possibilities.
4.1. Outline of the classification
The classification is based on the analysis of the intersection numbers $K\cdot C$ for irreducible components $C\subset E$.
Conceptually, the classification can be viewed as a finite assembly problem. Once all irreducible components with prescribed intersection behaviour are classified, the remaining configurations arise from combining a fixed finite collection of admissible building blocks, subject to the numerical constraints imposed by negative definiteness and the adjoint inequalities.
We first classify all components with $K\cdot C\leq 0$, which are shown to be highly restricted and are completely described in Lemmas 4.2 and 4.3. As a consequence, two elementary cases can be separated and treated independently, namely the irreducible case and the $K^{=0}$-graph case (Propositions 4.9 and 4.10). After excluding these cases, we may assume that there exists a component $C\subset E$ with $K\cdot C<0$.
Starting from such components, we introduce special linear configurations, called special $K^{>0}$-chains, whose possible numerical types are finite and classified in Proposition 4.6. Subtracting their contribution from $K$, we obtain a modified divisor $K_{1}$, and—under a disjointness assumption ()(*)—reduce the analysis to the sign of $K_{1}\cdot C$ for the remaining components. The case where ()(*) fails is treated separately in Proposition 4.11.
Using Lemmas 4.12 and 4.13, the case $K_{1}\cdot C<0$ is reduced to explicitly described configurations. Hence, it suffices to analyze the case where $E$ is a $(K_{1})^{=0}$-graph, which reduces to a finite combinatorial problem. This yields the complete list of possible configurations (Proposition 4.14), and hence the classification of $\epsilon$-adjoint log canonical singularities (Theorem 4.15).
4.2. $K$-non-positive curves
Definition 4.1.
Let $C$ be an irreducible curve contained in $E$. We define the following intersection-theoretic quantities:
(4.2) $d_{1}(C):=\left(E^{\rm n\!-\!inv}-\iota(C)C\right)\cdot C,\qquad d_{2}(C):=\left(E^{\rm inv}-(1-\iota(C))C\right)\cdot C,$
and set
$d(C):=d_{1}(C)+d_{2}(C)=(E-C)\cdot C.$
We then analyze complements satisfying $K\cdot C\leq 0$. We first consider the non-$\mathcal{F}$-invariant case.
Lemma 4.2.
Let $C\subset E$ be an irreducible non-$\mathcal{F}$-invariant curve. Then
(4.3) $KC=\epsilon\bigl(2p_{a}(C)-2+d(C)\bigr)+{\rm tang}(\mathcal{F},C)+d_{1}(C).$
Moreover, the inequality $KC\leq 0$ forces ${\rm tang}(\mathcal{F},C)=0$, and under this condition $C$ is necessarily one of the following explicitly listed types:
  1. (1)
    $d(C)=0$.
    1. (a)
      $C$ is a smooth elliptic curve. In this case, $KC=0$.
    2. (b)
      $C$ is a smooth rational curve. In this case, $KC=-2\epsilon$.
  2. (2)
    $d(C)=1$. $C$ is a smooth rational curve with $d_{1}(C)=0$ and $d_{2}(C)=1$. In this case, $KC=-\epsilon$.
  3. (3)
    $d(C)=2$. $C$ is a smooth rational curve with $d_{1}(C)=0$ and $d_{2}(C)=2$. In this case, $KC=0$.
Proof.
We compute
$KC$ $=(K_{\mathcal{F}}+E^{\rm n\!-\!inv})\cdot C+\epsilon(K_{X}+E)\cdot C$
$=(K_{\mathcal{F}}+C)\cdot C+\epsilon(K_{X}+C)\cdot C+(E^{\rm n\!-\!inv}-C)\cdot C+\epsilon(E-C)\cdot C.$
By Proposition 2.6, the genus formula, and the definitions of $d_{1}(C)$ and $d(C)$, this can be rewritten as
$KC={\rm tang}(\mathcal{F},C)+\epsilon\bigl(2p_{a}(C)-2+d(C)\bigr)+d_{1}(C).$
If $p_{a}(C)\geq 1$, then $KC\geq 0$. Moreover, $KC=0$ if and only if $p_{a}(C)=1$ and ${\rm tang}(\mathcal{F},C)=d(C)=0$. In this case, $C$ is a smooth elliptic curve. This is case (1-a).
Next we assume $p_{a}(C)=0$. If ${\rm tang}(\mathcal{F},C)+d_{1}(C)\geq 1$, then $KC\geq 1-2\epsilon>0$. Thus, we may assume ${\rm tang}(\mathcal{F},C)=d_{1}(C)=0$, in which case
$KC=(d(C)-2)\epsilon.$
Then $KC\leq 0$ implies case (1-b), (2) or (3). ∎
We now turn to the $\mathcal{F}$-invariant case.
Lemma 4.3.
Let $C\subset E$ be an irreducible $\mathcal{F}$-invariant curve. Then
(4.4) $KC=(1+\epsilon)\bigl(2p_{a}(C)-2+d(C)\bigr)+{\rm Z}(\mathcal{F},C)-d_{2}(C).$
Furthermore, the inequality $KC\leq 0$ holds if and only if $C$ is one of the following curves:
  1. (1)
    $d(C)=0$.
    1. (a)
      $C$ is an irreducible elliptic Gorenstein leaf. In this case, $KC=0$.
    2. (b)
      $C$ is a (2)(-2)-$\mathcal{F}$-curve. In this case, $KC=-2\epsilon$.
    3. (c)
      $C$ is a (1)(-1)-$\mathcal{F}$-curve. In this case, $KC=-(1+2\epsilon)$.
  2. (2)
    $d(C)=1$.
    1. (a)
      $C$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$. In this case, $KC=-(1+\epsilon)$.
    2. (b)
      $C$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=1$ and $d_{2}=0$. In this case, $KC=-\epsilon$.
    3. (c)
      $C$ is a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$. In this case, $KC=-\epsilon$.
  3. (3)
    $d(C)=2$.
    1. (a)
      $C$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$. In this case, $KC=0$.
    2. (b)
      $C$ is a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=2$. In this case, $KC=0$.
Proof.
We compute
$KC$ $=(K_{\mathcal{F}}+E^{\rm n\!-\!inv})\cdot C+\epsilon(K_{X}+E)\cdot C$
$=K_{\mathcal{F}}\cdot C+\epsilon(K_{X}+C)\cdot C+E^{\rm n\!-\!inv}\cdot C+\epsilon(E-C)\cdot C$
$=K_{\mathcal{F}}\cdot C+\epsilon(2p_{a}(C)-2)+d_{1}(C)+\epsilon d(C)$
$=K_{\mathcal{F}}\cdot C+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d(C)-d_{2}(C).$
By Proposition 2.8, we have
$K_{\mathcal{F}}\cdot C=2p_{a}(C)-2+{\rm Z}(\mathcal{F},C),$
so that
$KC=(1+\epsilon)\bigl(2p_{a}(C)-2+d(C)\bigr)+{\rm Z}(\mathcal{F},C)-d_{2}(C),$
with ${\rm Z}(\mathcal{F},C)-d_{2}(C)\geq 0$.
First, if $p_{a}(C)\geq 1$, then clearly $KC\geq 0$. Moreover, $KC=0$ if and only if $p_{a}(C)=1$ and ${\rm Z}(\mathcal{F},C)=d(C)=0$, in which case $C$ is an irreducible elliptic Gorenstein leaf, corresponding to case (1-a).
Now assume $p_{a}(C)=0$. Then
$KC=(1+\epsilon)(d(C)-2)+{\rm Z}(\mathcal{F},C)-d_{2}(C),$
so that $KC\leq 0$ implies $d(C)\leq 2$. Since $C^{2}<0$ and $p_{a}(C)=0$, Proposition 2.8 together with Lemma 2.9 ensures ${\rm Z}(\mathcal{F},C)\geq 1$.
Case $d(C)=0$. Here $d_{2}(C)=0$ and
$KC=-2(1+\epsilon)+{\rm Z}(\mathcal{F},C).$
Given $\epsilon\in(0,1/3)$, $KC\leq 0$ forces ${\rm Z}(\mathcal{F},C)\in\{1,2\}$, corresponding to cases (1-b) and (1-c).
Case $d(C)=1$. In this case, $d_{2}(C)\in\{0,1\}$, and
$KC=-(1+\epsilon)+{\rm Z}(\mathcal{F},C)-d_{2}(C).$
Since $\epsilon\in(0,1/3)$, $KC\leq 0$ implies
$({\rm Z}(\mathcal{F},C),d_{2}(C))=(1,1),\ (1,0),\ (2,1),$
which correspond to cases (2-b), (2-b) and (2-c), respectively.
Case $d(C)=2$. Then
$KC={\rm Z}(\mathcal{F},C)-d_{2}(C)\geq 0,$
and $KC=0$ occurs precisely when ${\rm Z}(\mathcal{F},C)=d_{2}(C)\in\{1,2\}$, corresponding to cases (3-a) and (3-b). ∎
Corollary 4.4.
Let $C\subset E$ be an irreducible curve with $d(C)\geq 1$.
  1. (1)
    If $0\leq KC<\epsilon$, then $KC=0$.
  2. (2)
    If $C$ is $\mathcal{F}$-invariant and $0\leq KC<(1+\epsilon)/2$, then $KC=0$.
Proof.
Write
$KC=\begin{cases}m\epsilon+n,&\text{if $C$ is non-$\mathcal{F}$-invariant},\\[2.84526pt] m(1+\epsilon)+n,&\text{if $C$ is $\mathcal{F}$-invariant},\end{cases}$
with $m=2p_{a}(C)-2+d(C)\geq-1$ and $n\geq 0$, as in (4.3) and (4.4).
For non-$\mathcal{F}$-invariant $C$, we have m=1m=-1 implies $KC=-\epsilon+n\in\{-\epsilon,1-\epsilon,2-\epsilon,\dots\}$, $m=0$ implies $KC=n\in\{0,1,2,\dots\}$, and $m\geq 1$ implies $KC\geq\epsilon$.
For $\mathcal{F}$-invariant $C$, we have m=1m=-1 implies $KC=-(1+\epsilon)+n\in\{-(1+\epsilon),-\epsilon,1-\epsilon,2-\epsilon,\dots\}$, $m=0$ implies $KC=n\in\{0,1,2,\dots\}$, and $m\geq 1$ implies $KC\geq 1+\epsilon$.
Since $\epsilon\in(0,1/3)$, $0\leq KC<\epsilon$ forces $KC=0$ for non-$\mathcal{F}$-invariant $C$, and $0\leq KC<(1+\epsilon)/2$ forces $KC=0$ for $\mathcal{F}$-invariant $C$. ∎
4.3. Special $K^{>0}$-chains in $E$
Let $C$ be an irreducible component of $E$. By Lemmas 4.2 and 4.3, any irreducible component $C\subset E$ with $KC<0$ satisfies $d(C)\leq 1$. If $d(C)=0$, then necessarily $E=C$. Otherwise, when $d(C)=1$, such a curve intersects ECE-C in a unique point, and hence naturally gives rise to a linear configuration inside $E$. This motivates the introduction of the following notion. These chains capture the configurations where a single curve with $KC<0$ is followed by curves with $KC\geq 0$, forming a linear segment of $E$.
Definition 4.5.
Let $F=C_{1}+\cdots+C_{r}$ be a chain of irreducible curves in $E$. We say that $F$ is a special $K^{>0}$-chain (resp. a special $K^{\geq 0}$-chain) if $F$ is a $K^{>0}$-graph (resp. $K^{\geq 0}$-graph) satisfying
$KC_{1}<0,\qquad KC_{i}\geq 0,\quad i\geq 2.$
We call $C_{1}$ the first curve of $F$.
Proposition 4.6.
Suppose $E$ has at least two components. Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{>0}$-chain with first curve $C_{1}$. Then $F$ belongs to one of the following configurations:
  • (A)
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve, and $C_{i}$ is (2)(-2)-$\mathcal{F}$-curve for all $i\geq 2$.
  • (B)
    Each $C_{i}$ is a (2)(-2)-$\mathcal{F}$-curve.
  • (C)
    $C_{i}$ is a (2)(-2)-$\mathcal{F}$-curve for all $i<r$, and $C_{r}$ is a (1)(-1)-$\mathcal{F}$-curve.
  • (D)
    There exists some $k\in[1,r]$ such that $C_{k}$ is a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{k})=0$. In this case, $C_{k-1}$ (if $k\geq 2$) and $C_{k+1}$ (if $k\leq r-1$) are (1)(-1)-$\mathcal{F}$-curves, and all remaining $C_{i}$ are (2)(-2)-$\mathcal{F}$-curves.
Moreover, in all cases above, the first curve $C_{1}$ satisfies $d(C_{1})=1$, while every other component $C_{i}$ ($i\geq 2$) satisfies
$d(C_{i})=2\quad\text{and}\quad KC_{i}=0.$
Furthermore,
[…]
Proof.
Since $KC_{1}<0$ and $d(C_{1})\geq 1$ (as $E$ has at least two components), Lemmas 4.2 and 4.3 imply that $C_{1}$ must fall into one of the following cases:
  • (i)
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$, in which case $KC_{1}=-(1+\epsilon)$;
  • (ii)
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=1$ and $d_{2}=0$, in which case $KC_{1}=-\epsilon$;
  • (iii)
    $C_{1}$ is a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$, in which case $KC_{1}=-\epsilon$;
  • (iv)
    $C_{1}$ is a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{1})=0$, $d_{1}=0$, and $d_{2}=1$, in which case $KC_{1}=-\epsilon$.
If $r=1$, then $F=C_{1}$, and the statement follows directly from the above list. Hence we may assume that $r\geq 2$.
Since $F$ is a special $K^{>0}$-chain, Proposition 3.6 yields
$0\leq KC_{k}<\frac{-KC_{1}}{\mu_{k}},\qquad k\geq 2,$
where $\mu_{k}=\det\bigl(-C_{i}C_{j}\bigr)_{1\leq i,j\leq k-1}$.
Cases (ii)–(iv). In these cases $KC_{1}=-\epsilon$, hence $0\leq KC_{i}<\epsilon$ for all $i\geq 2$. By Corollary 4.4, this implies $KC_{i}=0$. Consequently, for each $i\geq 2$, the curve $C_{i}$ must be one of the following:
  • a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$;
  • a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=2$;
  • a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{i})=0$, $d_{1}=0$ and $d_{2}=2$.
A straightforward combinatorial analysis of the possible adjacencies then yields the types (B), (C), and (D), and any additional case is excluded by the separatrix theorem (cf. Theorem 2.10).
Case (i). In this case, $KC_{1}=-(1+\epsilon)$ and $d(C_{1})=d_{2}(C_{1})=1$. Since $C_{1}$ is not $\mathcal{F}$-exceptional, we have $\mu_{2}=-C_{1}^{2}\geq 2$. It follows that $C_{2}$ is $\mathcal{F}$-invariant, as $d(C_{1})=d_{2}(C_{1})=1$, and
$0\leq KC_{2}<\frac{1+\epsilon}{\mu_{2}}\leq\frac{1+\epsilon}{2}.$
By Corollary 4.4, this implies $KC_{2}=0$. Consequently, $C_{2}$ is either a (2)(-2)-$\mathcal{F}$-curve or a (1)(-1)-$\mathcal{F}$-curve, by Lemma 4.3. The latter possibility is excluded by the separatrix theorem. Thus, $d(C_{2})=d_{2}(C_{2})=2$. Moreover, $C_{2}^{2}\leq-2$, since otherwise $C_{2}$ would be $\mathcal{F}$-exceptional, contradicting our assumptions. Therefore, $C_{1}+C_{2}$ is an $\mathcal{F}$-chain, and in particular $\mu_{3}\geq 2$.
Applying the same argument to $C_{3}$, we see that $C_{3}$ is $\mathcal{F}$-invariant, $d(C_{3})=d_{2}(C_{3})=2$, and $KC_{3}=0$, so that $C_{1}+C_{2}+C_{3}$ is again an $\mathcal{F}$-chain. Iterating this procedure, we conclude that $F$ is an $\mathcal{F}$-chain with the first curve $C_{1}$, hence of type (A). ∎
Remark 4.7.
Types (A) and (C) are both $\mathcal{F}$-chains. However, when viewed as special $K^{>0}$-chains, they are distinct.
The following result will be used repeatedly to control how a special $K^{\geq 0}$-chain can meet the rest of the exceptional divisor.
Corollary 4.8.
Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{\geq 0}$-chain in $E$ with the first curve $C$. If $C$ is an irreducible curve in $E\setminus F$ such that $F\cdot C>0$, then $F\cdot C=C_{r}\cdot C=1$ and
[…]
where $n=\det(-C_{i}C_{j})_{1\leq i,j\leq r}$. Moreover, such a curve $C$ is non-$\mathcal{F}$-invariant if and only if $F$ is of type (C).
Proof.
The statement is a direct consequence of Proposition 3.6 and Proposition 4.6, including the arguments in its proof. ∎
4.4. Classification
Let $G_{E}$ denote the dual graph of $E$, which is connected. We classify all possible configurations of $E$ according to the structure of $G_{E}$ and the distribution of the intersection numbers $KC$.
4.4.1. Case I: $E$ irreducible
We first treat the case where $E$ is irreducible.
Proposition 4.9.
If $E=C$ is irreducible, then $C$ belongs to one of the following cases:
  • (1)
    $C$ is an irreducible elliptic Gorenstein leaf. In this case, $KC=0$.
  • (2)
    $C$ is a smooth elliptic non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$. In this case, $KC=0$.
  • (3)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve. In this case, $KC=-2\epsilon$.
  • (4)
    $C$ is a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$. In this case, $KC=-2\epsilon$.
  • (5)
    $C$ is a (1)(-1)-$\mathcal{F}$-curve. In this case, $KC=-(1+2\epsilon)$.
Proof.
This corresponds to the case $d(C)=0$. The classification follows directly from Lemmas 4.2 and 4.3. ∎
4.4.2. Case II: $E$ is a $K^{=0}$-graph
We next consider the case where all components of $E$ have zero intersection with the adjoint divisor.
Proposition 4.10.
If $E$ is a $K^{=0}$-graph, that is, $KC=0$ for any component $C$ of $E$, then $E$ belongs to one of the following cases:
  1. (1)
    $E=C$ is an irreducible curve, where $C$ is an irreducible elliptic Gorenstein leaf, or a smooth elliptic non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$.
  2. (2)
    $E$ is a cycle of (2)(-2)-$\mathcal{F}$-curves.
Proof.
If $E$ consists of a single component $C$, then $KC=0$ and $d(C)=0$. Thus $C$ must be either an irreducible elliptic Gorenstein leaf, or a smooth elliptic non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$, as described in Lemmas 4.2 and 4.3. This is case (1).
Assume now that $E$ has at least two components. Then $d(C)\geq 1$ for any component $C$ of $E$. Since $KC=0$, Lemma 4.2 and Lemma 4.3 imply that each $C$ is a smooth rational curve with $d(C)=2$. Hence $E=C_{1}+\cdots+C_{r}$ is a cycle of smooth rational curves.
By Lemma 4.2 and 4.3, each $C_{i}$ must belong to one of the following types:
  • (i)
    a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$;
  • (ii)
    a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=2$;
  • (iii)
    a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$, $d_{1}=0$, and $d_{2}=2$.
A direct analysis using the separatrix theorem (cf. Theorem 2.10) shows that cases (i) and (iii) cannot occur. Therefore, each $C_{i}$ is a (2)(-2)-$\mathcal{F}$-curve, and $E$ is a cycle of such curves. This is case (2). ∎
4.4.3. Case III: the general case
We now treat the remaining configurations, namely those which are neither irreducible nor $K^{=0}$-graphs.
The analysis proceeds by a successive reduction process. Starting from components with $K\cdot C<0$, we isolate certain linear subconfigurations whose numerical behaviour is completely determined. These are the special $K^{>0}$-chains classified in Proposition 4.6.
Step 1: Interaction of special chains.
If two distinct special $K^{>0}$-chains intersect, their intersection forces strong structural constraints. This situation is analyzed in Proposition 4.11, which yields an explicit finite list of admissible configurations.
Proposition 4.11.
Suppose that there exist two special $K^{>0}$-chains $F_{1}$ and $F_{2}$, whose first curves are distinct, such that they share a common component or have nonzero intersection $F_{1}\cdot F_{2}>0$. Then $E=C_{1}+\cdots+C_{r}$ is a chain belonging to one of the following cases:
  • (1)
    $E$ is an $\mathcal{F}$-chain.
  • (2)
    $E$ is a chain of (2)(-2)-$\mathcal{F}$-curves.
  • (3)
    There exists some $k\in[1,r]$ such that $C_{k}$ is a non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{k})=0$, and both $C_{1}+\cdots+C_{k-1}$ and $C_{k+1}+\cdots+C_{r}$ are $\mathcal{F}$-chains whose first curves are $C_{k-1}$ and $C_{k+1}$, respectively. (See Figure 6.)
In particular, the corresponding point $p=\pi(E)$ is $\epsilon$-adjoint klt in all cases above.
Proof.
By Proposition 4.6, a finite analysis of the admissible building blocks yields the desired configurations, and the separatrix theorem (cf. Theorem 2.10) rules out any remaining possibilities.
For all components $C_{i}\subset E$, we have $KC_{i}\leq 0$, and $KC_{1}<0$. By Lemma 3.2, $E$ is a $K^{>0}$-graph, hence $p=\pi(E)$ is $\epsilon$-adjoint klt. ∎
From now on, we may therefore assume the following disjointness condition ()(*).
    $E$ is neither irreducible nor a $K^{=0}$-graph, and any two special $K^{>0}$-chains that start with different first curves are disjoint.
Step 2: Numerical reduction.
Under assumption ()(*), we can define the maximal special $K^{>0}$-chains in $E$. We adopt the following notation:
  • Let $\{\Theta_{1},\dots,\Theta_{s}\}$ denotes the set of all maximal special $K^{>0}$-chains of type (A) contained in $E$.
  • Let $\{\Xi_{1},\dots,\Xi_{t}\}$ denotes the set of all maximal special $K^{>0}$-chains of types (B), (C), or (D) contained in $E$.
We subtract the contribution of all maximal special $K^{>0}$-chains from $K$ and obtain a modified divisor
(4.5) $K_{1}:=K-\sum_{i=1}^{s}M(K,\Theta_{i})-\sum_{j=1}^{t}M(K,\Xi_{j}).$
It is clear that:
  • $E$ is a $(K_{1})^{\geq 0}$-graph (by Lemma 2.26(2));
  • for any component $C$ of $\Theta_{i}$ (resp. $\Xi_{j}$), one has $K_{1}C=0$.
We then analyze the components $C$ of $E$ not contained in any special $K^{>0}$-chains that satisfy $K_{1}\cdot C\leq 0$. We begin with the case where $C$ is not $\mathcal{F}$-invariant.
Lemma 4.12.
Under assumption ()(*), let $C\subset E$ be an irreducible non-$\mathcal{F}$-invariant curve not contained in any special $K^{>0}$-chain. If $K_{1}C\leq 0$, then one of the following holds:
  1. (1)
    $C$ is a smooth rational curve with ${\rm tang}(\mathcal{F},C)=0$, and $E$ is an $\mathcal{F}$-star fork (cf. Definition 2.22) of type $(l;m_{1},\dots,m_{l})$ centered at $C$, where $(l;m_{1},\dots,m_{l})$ is one of the following:
    1. (a)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, $(2,3,4)$, or $(2,3,5)$;
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, or $(2,4,4)$, or $(3,3,3)$;
    3. (c)
      $l=4$ and $(m_{1},\dots,m_{4})=(2,2,2,2)$.
    In particular,
    $d(C)=d_{2}(C)=l,\qquad K_{1}C\begin{cases}<0,&\text{for case (a)},\\[2.84526pt] =0,&\text{for cases (b) and (c)}.\end{cases}$
  2. (2)
    $C$ is a smooth rational curve with ${\rm tang}(\mathcal{F},C)=0$ meeting transversely two (1)(-1)-$\mathcal{F}$-curves $\Gamma_{1}$ and $\Gamma_{2}$ whose self-intersection are 2-2, where
    $d(\Gamma_{1})=d(\Gamma_{2})=1,\qquad d(C)=d_{2}(C)=3.$
    In this case, $K_{1}C=0$ and $C$ is disjoint from all special $K^{>0}$-chains except $\Gamma_{1}$ and $\Gamma_{2}$.
  3. (3)
    $C$ is a smooth rational curve with
    ${\rm tang}(\mathcal{F},C)=0,\qquad d_{1}(C)=0,\qquad d_{2}(C)=2,$
    and is disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
Proof.
We first observe that $d(C)\geq 1$. If $FC>0$ for some special $K^{>0}$-chain $F$, then $F$ must be of type (C) by Proposition 4.6 and Corollary 4.8.
Without loss of generality, assume that $C$ meets transversely the chains $\Xi_{1},\dots,\Xi_{l}$. Let $m_{j}$ denote the absolute value of the determinant of the intersection matrix of $\Xi_{j}$. Since each $\Xi_{j}$ is of type (C), for $j\leq l$ we have $m_{j}\geq 2$. By (4.3), we have
$KC={\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d_{1}(C)+\epsilon d_{2}(C).$
Thus,
[…]
$={\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d_{1}+\epsilon(d_{2}-l)+\sum_{j=1}^{l}\epsilon\left(1-\frac{1}{m_{j}}\right)$
$\geq{\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d_{1}+\epsilon(d_{2}-l)+\frac{\epsilon l}{2},$
where $d_{2}\geq l$ and $m_{j}\geq 2$.
Observe that if ${\rm tang}(\mathcal{F},C)\geq 1$, $p_{a}(C)\geq 1$, or $d_{1}\geq 1$, then $K_{1}C>0$ (since $d=d_{1}+d_{2}\geq 1$ and $\epsilon\in(0,1/3)$). Hence we may assume
${\rm tang}(\mathcal{F},C)=p_{a}(C)=d_{1}(C)=0,$
so that
$K_{1}C=\epsilon(d_{2}-l-2)+\sum_{j=1}^{l}\epsilon\left(1-\frac{1}{m_{j}}\right)\geq\frac{\epsilon}{2}(2(d_{2}-l)+l-4).$
Set
$x=d_{2}-l\,(\geq 0).$
Then $K_{1}C\leq 0$ gives
(4.6) $\epsilon(2x+l-4)\leq 0.$
On the other hand,
(4.7) $KC=\epsilon(x+l-2)\geq 0,$
because otherwise $C$ would start a special $K^{>0}$-chain, contradicting our assumption. Moreover, if equality holds in (4.7) and $l\geq 1$, then adjoining $C$ to a neighbouring $K^{>0}$-chain would produce a new special $K^{>0}$-chain, contradicting maximality. Hence,
(4.8) $\epsilon(x+l-2)=0\quad\Longrightarrow\quad l=0.$
We are thus reduced to a finite numerical problem for
$x,l\in\mathbb{Z}_{\geq 0},\qquad\epsilon\in(0,1/3),$
subject to (4.6)–(4.8). A straightforward inspection gives the only possible pairs:
$(x,l)=(2,0),\quad(1,2),\quad(0,3),\quad(0,4).$
Case $(2,0)$: Here $l=0$, $d=d_{2}=2$ and $K_{1}C=KC=0$. Moreover, $C$ is disjoint from all special $K^{>0}$-chains, giving case (3).
Case $(1,2)$: Here $l=2$, $d=d_{2}=3$, and
$K_{1}C=\epsilon\left(1-\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\geq 0,$
and equality holds if and only if $(m_{1},m_{2})=(2,2)$. This corresponds to case (2).
Case $(0,3)$ and $(0,4)$: Here $l\in\{3,4\}$ and
$K_{1}C=-2\epsilon+\sum_{j=1}^{l}\epsilon\left(1-\frac{1}{m_{j}}\right)=\epsilon\left(-2+l-\sum_{j=1}^{l}\frac{1}{m_{j}}\right).$
Hence, the condition $K_{1}C\leq 0$ is equivalent to
$\frac{1}{m_{1}}+\cdots+\frac{1}{m_{l}}\geq l-2.$
In this situation, $E=C+\Xi_{1}+\cdots+\Xi_{l}$ is an $\mathcal{F}$-star fork, giving case (1). ∎
We now consider the $\mathcal{F}$-invariant case.
Lemma 4.13.
Under the assumption ()(*), let $C$ be an irreducible $\mathcal{F}$-invariant curve contained in $E$, and assume that $C$ is not contained in any special $K^{>0}$-chain. If $K_{1}C\leq 0$, then one of the following holds:
  1. (1)
    $E=\Theta_{1}+\Theta_{2}+C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, meeting transversely two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of the following:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2;
    2. (b)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
    3. (c)
      $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
  2. (2)
    $E=C+\Theta_{1}+\Theta_{2}+\Xi_{1}$ is a fork centered at $C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, $\Theta_{1}$ and $\Theta_{2}$ are $\mathcal{F}$-chains, and $\Xi_{1}$ is a chain of (2)(-2)-$\mathcal{F}$-curves. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of the following:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2;
    2. (b)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
    3. (c)
      $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
  3. (3)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve that meets transversely two (1)(-1)-$\mathcal{F}$-curves $\Gamma_{1}$ and $\Gamma_{2}$, where
    $d(\Gamma_{1})=d(\Gamma_{2})=1,\qquad d(C)=d_{2}(C)=3.$
    In this case, $K_{1}C=0$ and $C$ is disjoint from all special $K^{>0}$-chains except $\Gamma_{1}$ and $\Gamma_{2}$.
  4. (4)
    $C$ is a (1)(-1)-$\mathcal{F}$-curve with
    $d_{1}(C)=d_{2}(C)=1,$
    and is disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
  5. (5)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve with
    $d(C)=d_{2}(C)=2,$
    and is disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
Moreover, we have:
  • Cases (1-b), (1-c), (2-b) and (2-c) cannot occur if $\epsilon<\tfrac{1}{5}$.
  • $K_{1}C<0$ for cases (1-a) and (2-a).
Proof.
We first observe that $d(C)\geq 1$. Without loss of generality, assume that $C$ meets the chains $\Theta_{1},\dots,\Theta_{k}$ transversely at points $p_{1},\dots,p_{k}$, and the chains $\Xi_{1},\dots,\Xi_{l}$ at points $q_{1},\dots,q_{l}$. Set
$V:=\sum_{i=1}^{k}M(K,\Theta_{i})+\sum_{j=1}^{l}M(K,\Xi_{j}).$
By Corollary 4.8, we have
$VC=\sum_{i=1}^{k}\frac{1+\epsilon}{n_{i}}+\sum_{j=1}^{l}\frac{\epsilon}{m_{j}},$
where $n_{i}$ (resp. $m_{j}$) denotes the absolute value of the determinant of the intersection matrix of $\Theta_{i}$ (resp. $\Xi_{j}$). In particular, $n_{i}\geq 2$.
By (4.4), we have
$KC=(1+\epsilon)(2p_{a}(C)-2+d(C))+{\rm Z}(\mathcal{F},C)-d_{2}(C).$
Hence
$K_{1}C$ $=KC-VC$
$=(1+\epsilon)(2p_{a}(C)-2+d-k-l)+l+{\rm Z}(\mathcal{F},C)-d_{2}$
$\quad+\sum_{i=1}^{k}(1+\epsilon)\Bigl(1-\frac{1}{n_{i}}\Bigr)+\sum_{j=1}^{l}\epsilon\Bigl(1-\frac{1}{m_{j}}\Bigr)$
$\geq(1+\epsilon)(2p_{a}(C)-2+d-k-l)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\frac{(1+\epsilon)k}{2}$
where $d-k-l\geq 0$ and ${\rm Z}(\mathcal{F},C)-d_{2}\geq 0$.
If $p_{a}(C)\geq 1$, then $K_{1}C\leq 0$ forces $k=l=d=0$, which contradicts our assumption. Hence we may assume $p_{a}(C)=0$. In this case, the dual graph of $C+\sum_{i=1}^{k}\Theta_{i}$ forms a tree, which implies by the separatrix theorem that
(4.9) ${\rm Z}(\mathcal{F},C)\geq k+1.$
Set
$x=d-k-l\,(\geq 0),\qquad y={\rm Z}(\mathcal{F},C)-d_{2}\,(\geq 0).$
Then we have
$K_{1}C\geq(1+\epsilon)(x-2)+l+y+\frac{(1+\epsilon)k}{2}.$
Since $K_{1}C\leq 0$, this yields
(4.10) $(1+\epsilon)\Bigl(x+\frac{k}{2}-2\Bigr)+l+y\leq 0.$
On the other hand,
(4.11) $KC=(1+\epsilon)(x+k+l-2)+y\geq 0,$
because otherwise $C$ would start a special $K^{>0}$-chain, contradicting our assumption. Moreover, if equality holds in (4.11) and $k+l\geq 1$, then adjoining $C$ to a neighbouring special $K^{>0}$-chain would produce a new special $K^{>0}$-chain, contradicting maximality. Hence
(4.12) $(1+\epsilon)(x+k+l-2)+y=0\quad\Longrightarrow\quad k=l=0.$
Finally, we claim
(4.13) $x+l+y\geq 1.$
Indeed, otherwise $x=l=y=0$, which implies
$k\leq{\rm Z}(\mathcal{F},C)=d_{2}\leq d=k,$
so that ${\rm Z}(\mathcal{F},C)=k$, contradicting (4.9).
We are thus reduced to a finite numerical problem for
$x,k,l,y\in\mathbb{Z}_{\geq 0},\qquad\epsilon\in(0,1/3),$
subject to (4.10)–(4.13). A straightforward inspection gives the only possible quadruples:
$(x,k,l,y)=(2,0,0,0),\quad(1,2,0,0),\quad(0,2,0,1),\quad(0,2,1,0).$
Case $(2,0,0,0)$. Here $k=l=0$, $d=2$, ${\rm Z}(\mathcal{F},C)=d_{2}\in\{1,2\}$, and $K_{1}C=KC=0$. Hence $C$ is either a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$ or a (2)(-2)-$\mathcal{F}$-curve with $d=d_{2}=2$, disjoint from all special $K^{>0}$-chains. This gives case (4) or (5).
Case $(1,2,0,0)$. Here $k=2$, $l=0$, $d=3$. We have
$K_{1}C=(1+\epsilon)\Bigl(1-\frac{1}{n_{1}}-\frac{1}{n_{2}}\Bigr),$
so that $K_{1}C\leq 0$ forces $(n_{1},n_{2})=(2,2)$, which further implies $K_{1}C=0$. Hence $C$ meets transversely two (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2. By (4.9), ${\rm Z}(\mathcal{F},C)\geq 3$, which leads to $d_{2}=d=3$. This gives case (3).
Cases $(0,2,0,1)$ and $(0,2,1,0)$. Here $k=2$, $d=l+2\in\{2,3\}$, and
$K_{1}C\geq 1-(1+\epsilon)\Bigl(\frac{1}{n_{1}}+\frac{1}{n_{2}}\Bigr),$
so that $K_{1}C\leq 0$ implies $\frac{1}{n_{1}}+\frac{1}{n_{2}}\geq 1/(1+\epsilon)>3/4$, hence $(n_{1},n_{2})=(2,2)$ or $(2,3)$, the latter only if $\epsilon\geq 1/5$. By (4.9), ${\rm Z}(\mathcal{F},C)\geq 3$.
If $l=0$, then $d=d_{2}=2$ and ${\rm Z}(\mathcal{F},C)=3$, giving case (1); if $l=1$, then $d=d_{2}={\rm Z}(\mathcal{F},C)=3$, and $\Xi_{1}$ is necessarily a special $K^{>0}$-chain of type (B), giving case (2). Moreover, $K_{1}C=-\epsilon$ for case (1-a) and $K_{1}C=-\epsilon/m_{1}$ for case (2-a). ∎
Step 3: Residual configurations. By Lemmas 4.12 and 4.13, components satisfying $K_{1}\cdot C\leq 0$ are again severely restricted. If some component satisfies $K_{1}\cdot C<0$, we obtain one of finitely many explicitly described configurations. Otherwise, $E$ is a $(K_{1})^{=0}$-graph, which also reduces to a finite combinatorial classification. More precisely, we have the following proposition.
Proposition 4.14.
Under the assumption ()(*), $E$ belongs to one of the following cases:
  1. (1)
    $E$ is an $\mathcal{F}$-star graph of type $(l;m_{1},\dots,m_{l})$ centered at $C$, where $C$ is a smooth rational curve with ${\rm tang}(\mathcal{F},C)=0$ and $(l;m_{1},\dots,m_{l})$ is one of the following:
    1. (a)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, $(2,3,4)$, or $(2,3,5)$;
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, or $(2,4,4)$, or $(3,3,3)$;
    3. (c)
      $l=4$ and $(m_{1},\dots,m_{4})=(2,2,2,2)$.
  2. (2)
    $E=\Theta_{1}+\Theta_{2}+C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, meeting transversely two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of the following:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2;
    2. (b)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
    3. (c)
      $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
  3. (3)
    $E=C+\Theta_{1}+\Theta_{2}+\Xi_{1}$ is a fork centered at $C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, $\Theta_{1}$ and $\Theta_{2}$ are $\mathcal{F}$-chains, and $\Xi_{1}$ is a chain of (2)(-2)-$\mathcal{F}$-curves. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of the following:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2;
    2. (b)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
    3. (c)
      $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
Moreover, we have:
  • (i)
    cases (2-b), (2-c), (3-b) and (3-c) cannot occur if $\epsilon<\tfrac{1}{5}$.
  • (ii)
    the corresponding point $p=\pi(E)$ is $\epsilon$-adjoint klt for cases (1-a), (2-a) and (3-a).
  • (iii)
    the corresponding point $p=\pi(E)$ is not $\epsilon$-adjoint klt for cases (1-b) and (1-c).
Proof.
If there exists an irreducible curve $C\subset E$ as in Lemma 4.12(1) or Lemma 4.13(1)(2), then we are in cases (1)–(3).
Otherwise, no such curve exists. Hence $E$ is a $(K_{1})^{=0}$-graph. By the preceding lemmas, $E$ is obtained by combining configurations of the following types:
  • special $K^{>0}$-chains;
  • configurations as in Lemma 4.12(2)(3);
  • configurations as in Lemma 4.13(3)(4)(5).
Nevertheless, any such $E$ is ruled out by the separatrix theorem (cf. Theorem 2.10) and assumption ()(*).
Next, we prove the last part.
(i) follows from Lemma 4.13.
(ii) For cases (1-a), (2-a) and (3-a), we see $K_{1}E_{i}\leq 0$ for any component $E_{i}\subset E$ and $K_{1}C<0$ by Lemma 4.12 and 4.13. On the other hand, $E$ is connected. It follows from Lemma 3.2 that $E$ is a $(K_{1})^{>0}$-graph, so $K^{>0}$-graph. This is equivalent to saying that $p=\pi(E)$ is $\epsilon$-adjoint klt.
(iii) For cases (1-b) and (1-c), $K_{1}C=0$ by Lemma 4.12. Thus, $M(K_{1},E)=0$ and so
$M(K,E)=M(K,\Xi_{1})+M(K,\Xi_{2}).$
This implies the coefficient of $C$ in $M(K,E)$ is 0, which means $p$ is not $\epsilon$-adjoint klt. ∎
4.4.4. Summary of the classification
Combining Propositions 4.9, 4.10, 4.11, and 4.14, we obtain a complete classification of admissible configurations.
Theorem 4.15.
Fix $\epsilon\in(0,1/3)$. Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface such that $p$ is $\epsilon$-adjoint log canonical. Let
$\pi\colon(X,\mathcal{F})\to(Y,\mathcal{G})$
be the minimal resolution, and write $E=\bigcup_{i=1}^{r}E_{i}$ for the exceptional divisor.
Then the exceptional divisor $E$ is one of the following configurations.
  1. (I)
    Foliated canonical configurations.
    • (1)
      A chain of smooth rational $\mathcal{F}$-invariant curves, consisting of:
      • (a)
        A $\mathcal{F}$-chain (Figure 1).
      • (b)
        Two (1)(-1)-$\mathcal{F}$-curves of self-intersection 2-2 joined by a bad tail (Figure 2).
      • (c)
        A chain of (2)(-2)-$\mathcal{F}$-curves.(See Definition 2.16)
    • (2)
      An $\mathcal{F}$-dihedral fork (Figure 3).
    • (3)
      An elliptic Gorenstein leaf, which is either a rational $\mathcal{F}$-invariant curve with a single node, or a cycle of (2)(-2)-$\mathcal{F}$-curves. (See Figure 4)
  2. (II)
    Foliated lc (non-canonical) configurations.
    • (1)
      An $\mathcal{F}$-star chain centered at a smooth rational curve: a chain $E=E_{1}+\cdots+E_{r}$ of smooth rational curves such that for some index $k$, the curve $E_{k}$ is non-$\mathcal{F}$-invariant with ${\rm tang}(\mathcal{F},E_{k})=0$, and the two subchains $E_{1}+\cdots+E_{k-1}$ and $E_{k+1}+\cdots+E_{r}$ (if nonempty) are $\mathcal{F}$-chains whose first curves are $E_{k-1}$ and $E_{k+1}$ respectively.
      (See Figure 6.)
    • (2)
      An $\mathcal{F}$-star fork of type $(l;m_{1},\dots,m_{l})$ centered at a smooth rational curve $C$ (cf. Definition 2.22), where $C$ is non-$\mathcal{F}$-invariant and ${\rm tang}(\mathcal{F},C)=0$, and where $(l;m_{1},\dots,m_{l})$ is one of the following:
      1. (a)
        $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, $(2,3,4)$, $(2,3,5)$;
      2. (b)
        $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, or $(2,4,4)$, or $(3,3,3)$;
      3. (c)
        $l=4$ and $(m_{1},\dots,m_{4})=(2,2,2,2)$.
      (See Figure 5.)
    • (3)
      A smooth elliptic non-$\mathcal{F}$-invariant curve $E_{1}$ with ${\rm tang}(\mathcal{F},E_{1})=0$.
  3. (III)
    Boundary configurations.
    • (1)
      $E=\Theta_{1}+\Theta_{2}+C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, meeting transversely two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$ of type $(2,3)$. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of:
      1. (a)
        $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
      2. (b)
        $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
    • (2)
      $E=C+\Theta_{1}+\Theta_{2}+\Xi_{1}$ is a single fork centered at $C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, $\Theta_{1}$ and $\Theta_{2}$ are $\mathcal{F}$-chains of type $(2,3)$, and $\Xi_{1}$ is a chain of (2)(-2)-$\mathcal{F}$-curves. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of:
      1. (a)
        $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 3-3;
      2. (b)
        $\Theta_{2}=\Gamma_{21}+\Gamma_{22}$, where $\Gamma_{21}$ (resp. $\Gamma_{22}$) is a (1)(-1)-$\mathcal{F}$-curve (resp. a (2)(-2)-$\mathcal{F}$-curve), both with self-intersection 2-2.
In addition, the following properties hold.
  • (i)
    Type (I) is foliated canonical, Type (II) is foliated lc (non-canonical), and Type (III) is not foliated lc. In particular, Type (III) cannot occur if $\epsilon<1/5$.
  • (ii)
    Among Type (I) and (II),
    • (I-1), (I-2), (II-1) and (II-2-a) are $\epsilon$-adjoint klt;
    • (I-3), (II-2-b),(II-2-c) and (II-3) are not $\epsilon$-adjoint klt.
Proof.
All possible configurations of the exceptional divisor $E$ are covered by Proposition 4.9, Proposition 4.10, Proposition 4.11, and Proposition 4.14. Hence, by collecting the conclusions of these propositions, we obtain both the classification of $E$ and the additional properties listed in the theorem. ∎
Corollary 4.16.
Fix $\epsilon\in(0,1/5)$. Suppose $p$ is an $\epsilon$-adjoint log canonical singularity of $(X,\mathcal{F})$. Then $p$ is log canonical both as a foliated singularity and as a surface singularity. Moreover, if $p$ is $\epsilon$-adjoint klt, then $p$ is klt as a singularity of $X$.
Proof.
It follows directly from Theorem 4.15, together with the classification of lc surface singularities (Theorem 2.24). ∎
Remark 4.17.
The classification carried out in this section should be understood as a finite numerical reduction procedure on the negative definite exceptional divisor $E$, rather than as a mere enumeration of combinatorial possibilities.
From this viewpoint, the same reduction mechanism, with minor adjustments to the numerical parameters, recovers the classical classifications of log canonical surface singularities (cf. Theorem 2.24) and log canonical foliated surface singularities (cf. Theorem 2.25), which fit naturally into the same numerical reduction framework.
5. $\epsilon$-adjoint canonical singularities for $\epsilon\in(0,1/4)$
Throughout this section, we fix $\epsilon\in(0,1/4)$. Under the $\epsilon$-adjoint canonical condition, we determine which of the configurations listed in Theorem 4.15 are ruled out and which can still occur as $\epsilon$-adjoint canonical singularities.
Lemma 5.1.
Types (I-3), (II-2-b), (II-2-c), and (II-3) are not $\epsilon$-adjoint canonical.
Proof.
This follows directly from Theorem 4.15. ∎
Lemma 5.2.
Types (III) is not $\epsilon$-adjoint canonical.
Proof.
Let $K^{\prime}:=K_{\mathcal{F}}+\epsilon K_{X}$. For each $\Theta_{i}$, let $M(K^{\prime},\Theta_{i})$ denote the effective divisor supported on $\Theta_{i}$ such that
$M(K^{\prime},\Theta_{i})\cdot\Gamma=K^{\prime}\cdot\Gamma$
for any irreducible component $\Gamma$ of $\Theta_{i}$. Set
$W:=K^{\prime}-M(K^{\prime},\Theta_{1})-M(K^{\prime},\Theta_{2}).$
Suppose that $p$ is an $\epsilon$-adjoint canonical singularity. Then $E$ is a $(K^{\prime})^{\geq 0}$-graph and hence also a $W^{\geq 0}$-graph. Clearly, $W\cdot\Gamma\geq 0$ for any component $\Gamma\subset E$ with $\Gamma\neq C$. We claim that $W\cdot C>0$, contradicting the assumption that $E$ is a $W^{\geq 0}$-graph.
We have
$M(K^{\prime},\Theta_{1})=\frac{1}{2}\Gamma_{11}.$
If $\Theta_{2}$ is of type (III-1-a) or (III-2-a), then
$M(K^{\prime},\Theta_{2})=\frac{1-\epsilon}{3}\Gamma_{21}.$
Note that
$K^{\prime}\cdot C=K_{\mathcal{F}}\cdot C+\epsilon K_{X}\cdot C\geq 1-\epsilon.$
Hence,
$W\cdot C$ $=K^{\prime}\cdot C-M(K^{\prime},\Theta_{1})\cdot C-M(K^{\prime},\Theta_{2})\cdot C$
$\geq 1-\epsilon-\frac{1}{2}-\frac{1-\epsilon}{3}=\frac{1-4\epsilon}{6}>0.$
If $\Theta_{2}$ is of type (III-1-b) or (III-2-b), then
$M(K^{\prime},\Theta_{2})=\frac{2}{3}\Gamma_{21}+\frac{1}{3}\Gamma_{22}.$
In this case, we must have $C^{2}\leq-2$; otherwise, the intersection matrix of $\Gamma_{11}+C+\Gamma_{21}$ is not negative definite. Hence $K^{\prime}\cdot C\geq K_{\mathcal{F}}C=-C^{2}\geq 2$, and therefore
$W\cdot C\geq 2-\frac{1}{2}-\frac{1}{3}=\frac{7}{6}>0.$
Lemma 5.3.
Type (II-2-a) is not $\epsilon$-adjoint canonical.
Proof.
We follow the notation of Lemma 5.2. Let
$W:=K^{\prime}-\sum_{i=1}^{3}M(K^{\prime},\Theta_{i}),$
where $\Theta_{1},\Theta_{2},\Theta_{3}$ are the $\mathcal{F}$-chains attached to $C$ as in type (II-2-a) of Theorem 4.15. Similar to the proof of Lemma 5.2, it suffices to show that $W\cdot C>0$.
In case (II-2-a), we have $e_{C}:=-C^{2}\geq 2$, since the intersection matrix of $C+\Theta_{1}+\Theta_{2}+\Theta_{3}$ is negative definite. Hence
$K^{\prime}\cdot C=K_{\mathcal{F}}C+\epsilon K_{X}C=(1+\epsilon)e_{C}-2\epsilon\geq 2.$
Case $(2,2,n)$, $n\geq 2$.
We have
$M(K^{\prime},\Theta_{1})=\tfrac{1}{2}\Gamma_{11},\qquad M(K^{\prime},\Theta_{2})=\tfrac{1}{2}\Gamma_{21},$
so
$M(K^{\prime},\Theta_{1})\cdot C=M(K^{\prime},\Theta_{2})\cdot C=\tfrac{1}{2}.$
Since the coefficients of $M(K^{\prime},\Theta_{3})$ are at most (n1)/n(n-1)/n, we obtain
$M(K^{\prime},\Theta_{3})\cdot C\leq\frac{n-1}{n}.$
Therefore
$W\cdot C\geq 2-\frac{1}{2}-\frac{1}{2}-\frac{n-1}{n}=\frac{1}{n}>0.$
Case $(2,3,n)$, $n=3,4,5$.
Here $\Theta_{1}=\Gamma_{11}$ with $\Gamma_{11}^{2}=-2$, and $\Theta_{2}$ is either
  • (i)
    $\Gamma_{21}$ with $\Gamma_{21}^{2}=-3$, or
  • (ii)
    $\Gamma_{21}+\Gamma_{22}$ with $\Gamma_{21}^{2}=\Gamma_{22}^{2}=-2$.
In case (i),
$M(K^{\prime},\Theta_{1})\cdot C=\tfrac{1}{2},\quad M(K^{\prime},\Theta_{2})\cdot C=\frac{1-\epsilon}{3},\quad M(K^{\prime},\Theta_{3})\cdot C\leq\frac{n-1}{n}\leq\frac{4}{5},$
hence
$W\cdot C\geq 2-\frac{1}{2}-\frac{1-\epsilon}{3}-\frac{4}{5}=\frac{11}{30}+\frac{\epsilon}{3}>0.$
In case (ii),
$M(K^{\prime},\Theta_{1})\cdot C=\tfrac{1}{2},\quad M(K^{\prime},\Theta_{2})\cdot C=\tfrac{2}{3},\quad M(K^{\prime},\Theta_{3})\cdot C\leq\frac{n-1}{n}\leq\frac{4}{5},$
hence
$W\cdot C\geq 2-\frac{1}{2}-\frac{2}{3}-\frac{4}{5}=\frac{1}{30}>0.$
Lemma 5.4.
Let $E=C_{1}+\cdots+C_{r}$ be as in type (I-1-c) of Theorem 4.15. If $p=\pi(E)$ is $\epsilon$-adjoint canonical, then
$(-C_{1}^{2},\ldots,-C_{r}^{2})=(2,\ldots,2)\quad\text{or}\quad(2,\ldots,2,3,2,\ldots,2,1).$
If moreover $p=\pi(E)$ is $\epsilon$-adjoint terminal, then $p$ is a smooth point of $Y$, and around $p$ the foliation $\mathcal{G}$ is locally defined (cf. [Bru15, pp. 8]) by
$\omega=(ry+x^{r})\,{\rm d}x-x\,{\rm d}y.$
Proof.
Note that each $C_{i}$ is not $\mathcal{F}$-exceptional, and $E$ is a $(K^{\prime})^{\geq 0}$-graph for $K^{\prime}=K_{\mathcal{F}}+\epsilon K_{X}$.
If $C_{i}^{2}\leq-2$ for all $i$, then
$K^{\prime}C=K_{\mathcal{F}}C-\epsilon K_{X}C=\epsilon K_{X}C\geq 0$
and equality holds if and only if $C_{i}^{2}=-2$. Thus, the fact that $E$ is a $(K^{\prime})^{\geq 0}$-graph implies $C_{i}^{2}=-2$ for all $i$. In this case, $E$ is a $(K^{\prime})^{=0}$-graph.
Suppose that $C_{i}^{2}=-1$ for some $i$. Then necessarily $i=1$ or $i=r$; otherwise, $C_{i}$ would be $\mathcal{F}$-exceptional, a contradiction. By the separatrix theorem (cf. Theorem 2.10) and Lemma 2.18, there exists at most one such curve $C_{i}$. Without loss of generality, assume $C_{r}^{2}=-1$. By Lemma 2.18, there exists exactly one singularity of $\mathcal{F}$ on $C_{r}$, which is a saddle-node whose weak separatrix is contained in $C_{r}$.
For $i\leq r-1$, we have $C_{i}^{2}\leq-2$. Let $l$ be the integer such that $C_{i}^{2}=-2$ for $l\leq i\leq r-1$ and $C_{l-1}^{2}\leq-3$. Contracting $C_{l}+C_{l+1}+\cdots+C_{r}$, we obtain a birational morphism
$\sigma\colon(X,\mathcal{F})\longrightarrow(X^{\prime},\mathcal{F}^{\prime}).$
Let
$E^{\prime}=C^{\prime}_{1}+\cdots+C^{\prime}_{l-1},\qquad C^{\prime}_{i}:=\sigma_{*}(C_{i}).$
Then $E^{\prime}$ is a $(K_{\mathcal{F}^{\prime}}+\epsilon K_{X^{\prime}})^{\geq 0}$-graph. Moreover, $(C^{\prime}_{i})^{2}\leq-2$ and $K_{\mathcal{F}^{\prime}}\cdot C^{\prime}_{i}=0$ for all $i\leq l-1$. It follows that $(C^{\prime}_{i})^{2}=-2$ for all $i\leq l-1$. Therefore,
$(-C_{1}^{2},\ldots,-C_{r}^{2})=(2,\ldots,2,3,2,\ldots,2,1).$
If $l\geq 2$, then $p=\pi(E)$ is not $\epsilon$-adjoint terminal. If $l=1$, then $p$ is a smooth point of $Y$. In this case, around $p$ the foliation $\mathcal{G}$ is locally defined (cf. [Bru15, pp. 8]) by
$\omega=(ry+x^{r})\,{\rm d}x-x\,{\rm d}y.$
Lemma 5.5.
If type (I-1-b) (resp. type (I-2)) is $\epsilon$-adjoint canonical, then every component of $E$ has self-intersection 2-2.
Proof.
Let $C$ denote the tail in type (I-1-b) (resp. the fork center in type (I-2)). Let $\Theta_{1},\Theta_{2}$ be the $\mathcal{F}$-chains attached to $C$ as in type (I-1-b) (resp. (I-2)). Then $\Theta_{i}=\Gamma_{i}$, where $\Gamma_{i}$ is a (1)(-1)-$\mathcal{F}$-curve with self-intersection 2-2 for $i=1,2$. Moreover, every component of $E$ has self-intersection $\leq-2$ in both types (I-1-b) and (I-2).
We follow the notation of Lemma 5.2. Set
$W:=K^{\prime}-\sum_{i=1}^{2}M(K^{\prime},\Theta_{i})=K^{\prime}-\frac{1}{2}\Gamma_{1}-\frac{1}{2}\Gamma_{2}.$
Since $E$ is a $(K^{\prime})^{\geq 0}$-graph, it is also a $W^{\geq 0}$-graph.
We compute
$W\Gamma_{1}=W\Gamma_{2}=0$
and
$WC=K^{\prime}C-1=K_{\mathcal{F}}C+\epsilon K_{X}C-1=\epsilon(-2-C^{2})\geq 0.$
For any remaining curve $C^{\prime}$ in type (I-2), we have
$WC^{\prime}=K^{\prime}C^{\prime}=K_{\mathcal{F}}C^{\prime}+\epsilon K_{X}C^{\prime}=\epsilon(-2-(C^{\prime})^{2})\geq 0.$
Thus, by Lemma 2.26, the fact that $E$ is a $W^{\geq 0}$-graph implies $C^{2}=-2$ and $(C^{\prime})^{2}=-2$. Therefore, every component of $E$ has self-intersection 2-2, as claimed. ∎
Lemma 5.6.
Let $E=E_{1}+\cdots+E_{k}+\cdots+E_{r}$ be as in type (II-1) of Theorem 4.15, where $E_{k}$ is the unique non-$\mathcal{F}$-invariant curve. If $p=\pi(E)$ is $\epsilon$-adjoint canonical, then $E_{k}^{2}=-1$ and $r\geq 2$.
Proof.
We use the notation as in Lemma 5.2. Let $C:=E_{k}$. In this setting, the curve $C$ meets $l\leq 2$ $\mathcal{F}$-chains, denoted by $\Theta_{1},\dots,\Theta_{l}$. For each $i$, let $\Theta^{\prime}_{i}$ be a maximal subchain of $\Theta_{i}$ such that $\Theta^{\prime}_{i}$ is a $(K^{\prime})^{\geq 0}$-graph. Set
$W:=K^{\prime}-\sum_{i=1}^{l}M(K^{\prime},\Theta^{\prime}_{i}).$
Then $E$ is a $W^{\geq 0}$-graph.
By the definition of $M(K^{\prime},\Theta^{\prime}_{i})$ (cf. Definition 3.1) and the maximality of the $\Theta^{\prime}_{i}$’s, we have:
  • $W\cdot\Gamma=0$ for any irreducible curve $\Gamma\subset\bigcup_{i}\Theta^{\prime}_{i}$, and
  • $W\cdot\Gamma\geq 0$ for any irreducible curve $\Gamma\subset\bigl(\bigcup_{i}\Theta_{i}\bigr)\setminus\bigl(\bigcup_{i}\Theta^{\prime}_{i}\bigr)$.
Assume that $C^{2}\leq-2$. Then
$K^{\prime}\cdot C=K_{\mathcal{F}}\cdot C+\epsilon K_{X}\cdot C\geq 2.$
Since $M(K^{\prime},\Theta^{\prime}_{i})\cdot C<1$ for all $i$, it follows that
$W\cdot C>2-l\geq 0,$
which contradicts the fact that $E$ is a $W^{\geq 0}$-graph. Thus $C^{2}=-1$.
Next, we show that $l\geq 1$, or equivalently $r\geq 2$. If $l=0$, then $W=K^{\prime}$. In this case,
$W\cdot C=K^{\prime}\cdot C=1-\epsilon>0,$
which again contradicts the fact that $E$ is a $W^{\geq 0}$-graph. Therefore, $l\geq 1$, and hence $r\geq 2$. ∎
The above lemmas exclude all configurations in Theorem 4.15 except the following five types. We therefore obtain the complete classification of $\epsilon$-adjoint canonical singularities for $\epsilon\in(0,1/4)$.
Theorem 5.7.
Fix $\epsilon\in(0,\tfrac{1}{4})$. Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface, and assume that $p$ is an $\epsilon$-adjoint canonical singularity of $(Y,\mathcal{G})$. Let
$\pi\colon(X,\mathcal{F})\to(Y,\mathcal{G})$
be the minimal resolution of $(Y,\mathcal{G},p)$, with exceptional divisor $E=\bigcup_{i=1}^{r}E_{i}$. Then $E$ is one of the following:
  1. (1)
    An $\mathcal{F}$-chain. (See Figure 1.)
  2. (2)
    A chain of (2)(-2)-$\mathcal{F}$-curves. (See Definition 2.16.)
  3. (3)
    A chain $E_{1}+E_{2}+E_{3}$, where $E_{1}$ and $E_{3}$ are (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2 and $E_{2}$ is a bad tail. Moreover, in this case, $E_{2}^{2}=-2$. (See Figure 2.)
  4. (4)
    An $\mathcal{F}$-dihedral fork, all of whose components have self-intersection 2-2. (See Figure 3.)
  5. (5)
    An $\mathcal{F}$-star chain centered at a smooth rational curve, that is, a chain with exactly one smooth rational non-$\mathcal{F}$-invariant curve $E_{k}$ with ${\rm tang}(\mathcal{F},E_{k})=0$, such that the two sides are $\mathcal{F}$-chains. Moreover, in this case, $E_{k}^{2}=-1$ and $r\geq 2$. (See Figure 6.)
Moreover, the following holds:
  • (i)
    Types (1)(4) are foliated canonical and type (5) is foliated log canonical.
  • (ii)
    In type (2), the point $p\in Y$ is either a Du Val singularity or a smooth point of $Y$. In the former case, $p$ is not $\epsilon$-adjoint terminal.
  • (iii)
    Types (3) and (4) are not $\epsilon$-adjoint terminal; moreover, they are Du Val singularities of $Y$.
Remark 5.8.
By Theorem 5.7, the instability of the canonical condition under adjoint perturbations is entirely due to configuration (5).
There is no uniform constant $\epsilon_{0}>0$ that excludes configuration (5) for all $\epsilon<\epsilon_{0}$ (see Example 6.1). However, for each fixed index $i(\mathcal{G})$, there exists a constant $\epsilon_{0}(i(\mathcal{G}))>0$ such that configuration (5) does not occur whenever $\epsilon<\epsilon_{0}(i(\mathcal{G}))$. For instance, [LWX25, Proposition 1.9] shows that if $K_{\mathcal{G}}$ is big, one may take $\epsilon_{0}(i(\mathcal{G}))=\frac{1}{3i(\mathcal{G})}$.
Corollary 5.9.
Fix $\epsilon\in(0,1/4)$. Suppose $p$ is an $\epsilon$-adjoint canonical singularity of $(X,\mathcal{F})$. Then $p$ is log canonical as a foliated singularity and klt as a surface singularity.
Proof.
It follows directly from Theorem 5.7, together with the classification of lc surface singularities (Theorem 2.24). ∎
5.1. Application to adjoint MMP
We now explain how the $\epsilon$-adjoint canonical classification yields a natural threshold for the adjoint minimal model program.
Let $(X,\mathcal{F})$ be a foliated surface, where $X$ is smooth and $\mathcal{F}$ is a rank one foliation with canonical singularities. Fix $\epsilon\in(0,1/4)\cap\mathbb{Q}$. We now run an $\epsilon$-adjoint MMP for $(X,\mathcal{F})$.
Choose a maximal adjoint collection of $(K_{\mathcal{F}}+\epsilon K_{X})^{\geq 0}$-graphs:
$\{\,\mathfrak{E}_{1},\cdots,\mathfrak{E}_{s}\,\}.$
Such a collection may not be unique if $K_{\mathcal{F}}+\epsilon K_{X}$ is not big.
By contracting $\mathfrak{E}_{1},\cdots,\mathfrak{E}_{s}$, we obtain a $(K_{\mathcal{F}}+\epsilon K_{X})$-non-positive birational contraction
$\varphi:(X,\mathcal{F})\longrightarrow(Y,\mathcal{G}).$
The pair $(Y,\mathcal{G})$ satisfies the following properties:
  • (a)
    For any irreducible curve $C\subset Y$, if
    $(K_{\mathcal{G}}+\epsilon K_{Y})\cdot C\leq 0,$
    then $C^{2}\geq 0$.
  • (b)
    $(Y,\mathcal{G})$ has at most $\epsilon$-adjoint canonical singularities. By Corollary 5.9, it follows that $Y$ has klt (hence rational) singularities and that $\mathcal{G}$ has lc singularities.
Consequently, we distinguish two cases:
  • (i)
    If $K_{\mathcal{F}}+\epsilon K_{X}$ is pseudo-effective, then $K_{\mathcal{G}}+\epsilon K_{Y}$ is nef. In particular, if $K_{\mathcal{F}}+\epsilon K_{X}$ is big, then $K_{\mathcal{G}}+\epsilon K_{Y}$ is ample. Since $Y$ has rational singularities, this follows from [Tan04, Theorem 4.8]. Moreover, [Tan04] provides an explicit number $\alpha=\alpha(K_{\mathcal{G}}+\epsilon K_{Y})$ such that $\alpha\cdot(K_{\mathcal{G}}+\epsilon K_{Y})$ is very ample (see also [LWX25, Corollary 1.2]).
  • (ii)
    If $K_{\mathcal{F}}+\epsilon K_{X}$ is not pseudo-effective, then neither is $K_{\mathcal{G}}+\epsilon K_{Y}$. Hence there exists an irreducible curve $C\subset Y$ with $C^{2}\geq 0$ and
    $(K_{\mathcal{G}}+\epsilon K_{Y})\cdot C<0.$
    In particular,
    $K_{Y}\cdot C<0\quad\text{or}\quad K_{\mathcal{G}}\cdot C<0.$
    By [KM98, Corollary 3.17] and [SS23, Theorem 2.8], there exists a morphism $f:Y\to Z$ such that $\rho(Y/Z)=1$ and $(K_{\mathcal{G}}+\epsilon K_{Y})$ is $f$-ample.
6. Key Examples
In this section, we present two illustrative examples.
Example 6.1.
Let $p$ be a singularity of a foliated surface $(X,\mathcal{F})$. Assume that $X$ is smooth at $p$, and that $\mathcal{F}$ is locally defined near $p=(0,0)$ by the $1$-form
$\omega=ny\,\mathrm{d}x-x\,\mathrm{d}y.$
Consider the minimal resolution $\sigma\colon(X^{\prime},\mathcal{F}^{\prime})\to(X,\mathcal{F})$ over $p$ (see Figure 7).
$\cdots\cdots$
$\bar{E}_{2}(-2)$
$\bar{E}_{1}(-2)$
$\bar{E}_{n-1}(-2)$
$E_{n}(-1)$
Figure 7.
Denote by $q_{i}$ (resp. $E_{i}$), $i=1,\dots,n$, the successive blow-up points (resp. exceptional curves). In this case, we have
$l(q_{i})=a(q_{i})=1,\quad i\leq n-1,\qquad\text{and}\qquad l(q_{n})=2>a(q_{n})=1.$
In particular, each $E_{i}$ is $\mathcal{F}^{\prime}$-invariant for $i\leq n-1$, while $E_{n}$ is not $\mathcal{F}^{\prime}$-invariant. A straightforward computation yields
$K_{\mathcal{F}^{\prime}}$ $=\sigma^{*}(K_{\mathcal{F}})-E_{n},$
$K_{X^{\prime}}$ $=\sigma^{*}(K_{X})+\bar{E}_{1}+2\bar{E}_{2}+\cdots+(n-1)\bar{E}_{n-1}+nE_{n}.$
Hence,
$K_{\mathcal{F}^{\prime}}+\epsilon K_{X^{\prime}}=\sigma^{*}(K_{\mathcal{F}}+\epsilon K_{X})+\epsilon\sum_{i=1}^{n-1}i\bar{E}_{i}+(n\epsilon-1)E_{n}.$
For any $\epsilon\in(0,1)$, one can choose $n$ sufficiently large so that $n\epsilon-1\geq 0$. In this situation, the point $p$ is $\epsilon$-adjoint canonical but not foliated canonical.
Remark 6.2.
Example 6.1 corresponds to type (5) in Theorem 5.7, which is the only configuration in which an $\epsilon$-adjoint canonical singularity fails to be foliated canonical for sufficiently small $\epsilon$.
Example 6.3.
Let $p$ be a singularity of a foliated surface $(X,\mathcal{F})$. Assume that $X$ is smooth at $p$, and that $\mathcal{F}$ is locally defined near $p=(0,0)$ by the $1$-form
$\omega=x\,\mathrm{d}x+y^{2}\,\mathrm{d}y.$
Consider the minimal resolution $\sigma\colon(X^{\prime},\mathcal{F}^{\prime})\to(X,\mathcal{F})$ over $p$ (see Figure 8).
$\bar{E}_{1}(-3)$
$\bar{E}_{2}(-2)$
$E_{3}(-1)$
Figure 8.
Denote by $q_{i}$ (resp. $E_{i}$) the successive blow-up points (resp. exceptional curves). In this case,
$l(q_{1})=l(q_{2})=1,\qquad l(q_{3})=a(q_{3})=2.$
In particular, all $E_{i}$ are $\mathcal{F}^{\prime}$-invariant. A direct computation gives
$K_{\mathcal{F}^{\prime}}$ $=\sigma^{*}(K_{\mathcal{F}})-E_{3},$
$K_{X^{\prime}}$ $=\sigma^{*}(K_{X})+\bar{E}_{1}+2\bar{E}_{2}+4E_{3}.$
Hence,
$K_{\mathcal{F}^{\prime}}+\frac{1}{4}K_{X^{\prime}}$ $=\sigma^{*}\!\left(K_{\mathcal{F}}+\frac{1}{4}K_{X}\right)+\frac{1}{4}\bar{E}_{1}+\frac{1}{2}\bar{E}_{2},$
$K_{\mathcal{F}^{\prime}}+\frac{1}{5}K_{X^{\prime}}$ $=\sigma^{*}\!\left(K_{\mathcal{F}}+\frac{1}{5}K_{X}\right)+\frac{1}{5}\bar{E}_{1}+\frac{2}{5}\bar{E}_{2}-\frac{1}{5}E_{3}.$
It follows that $p$ is $\tfrac{1}{4}$-adjoint canonical and $\tfrac{1}{5}$-adjoint log canonical, but it is not a log canonical singularity of $\mathcal{F}$.
Remark 6.4.
Example 6.3 corresponds to type (III-1-a) in Theorem 4.15, i.e., the boundary configuration that is not foliated log canonical.
Acknowledgements
The author sincerely thanks Professors Shengli Tan, Jun Lu, and Xin Lü for their continuous encouragement and support, and Professor Xiaohang Wu for many helpful discussions. He is also grateful to Professor Jihao Liu for raising the question concerning the sharp stability threshold in the adjoint log canonical setting, which inspired the development of this work.